### v2

# Einstein metrics on homogeneous superspaces

Abstract
This paper initiates the study of the Einstein equation on homogeneous supermanifolds. First, we produce explicit curvature formulas for graded Riemannian metrics on these spaces. Next, we present a construction of homogeneous supermanifolds by means of Dynkin diagrams, resembling the construction of generalised flag manifolds in classical (non-super) theory. We describe the Einstein metrics on several classes of spaces obtained through this approach. Our results provide examples of compact homogeneous supermanifolds on which the Einstein equation has no solutions, discrete families of solutions, and continuous families of Ricci-flat solutions among invariant metrics. These examples demonstrate that the finiteness conjecture from classical homogeneous geometry fails on supermanifolds, and challenge the intuition furnished by Bochner’s vanishing theorem.
1 Introduction
Over the past few decades, considerable efforts have been made to extend Riemannian geometry to the realm of supermanifolds; see, for example, the classic book [DeW92] and the more recent texts [Rog07, CCF11, Kes19]. Notions such as Riemannian metric, Levi-Civita connection and Riemann curvature have thus been generalised to this setting. The purpose of the present paper is to address some of the fundamental mathematical aspects of these notions, building on explicit curvature formulas that we derive for certain types of supermanifolds.
[…]
One of the major goals in Riemannian geometry is to determine which manifolds admit ‘special’ geometric structures. Once the foundational concepts are established and their basic properties understood, one may search for metrics with distinguished features. For example, there is a large body of literature on metrics with positive or non-negative sectional curvature; see the collection [HHL14]. Of particular interest in Riemannian geometry are metrics whose curvature is, in some sense, constant (think of round metrics on spheres as model examples). A natural way to capture the idea of constant curvature mathematically is through the Einstein equation
$\operatorname{Ric}(g)=cg,$ (1.1)
where $\operatorname{Ric}$ denotes the Ricci curvature, $g$ the metric, and $c$ the Einstein constant. Solutions to this equation are known as Einstein metrics, and many of the classical results on the associated Einstein manifolds are collected in the celebrated monograph [Bes87]. The analysis of (1.1) also underpins the classification of three-manifolds via Thurston’s geometrisation conjecture [MT06], and plays a pivotal role in a raft of other breakthrough insights and constructions in geometry, analysis and theoretical physics.
Kähler geometry provides an important class of Einstein metrics. These include Ricci-flat metrics on Calabi–Yau manifolds appearing in string theory; see [CHSW85, Gre97, Pol98] and references therein. Outside the Kähler setting, the majority of known Einstein metrics have been found through symmetry reduction. One typically assumes that the manifold admits an action of a Lie group that either is transitive or has principal orbits of co-dimension one. This helps reduce (1.1) from a system of PDEs to a system of algebraic equations or ODEs. The theory of Einstein metrics on homogeneous spaces is particularly fruitful, and includes a plethora of infinite families and isolated examples [WZ86, Boh04, BWZ04, Arv15]. The constructions in the present paper are inspired by that theory.
Another goal in Riemannian geometry is the establishment and exploration of links between the geometry and the topology of a manifold. For instance, the Bonnet–Myers theorem states that the positivity of the Ricci curvature implies the finiteness of the fundamental group. In the presence of symmetry, the links are particularly strong. A compact homogeneous space, for example, typically does not admit a metric with non-positive Ricci curvature. According to the Bochner theorem (see [Bes87, Theorem 1.84]), if $G_{0}/K_{0}$ is a compact homogeneous space with non-positive Ricci curvature, then the connected component of the identity of $G_{0}$ must be abelian.
While supergeometric generalisations of Ricci-flat Kähler metrics on Calabi–Yau manifolds have been considered in the literature (see, e.g., [RW05, Zho05, ARS16]), there seem to be no known results regarding metrics on supermanifolds with distinguished curvature properties but without an immediate connection to a Kähler-type structure. In this paper, we initiate the study of the Einstein equation (1.1) on homogeneous superspaces. Our work is motivated by the theory of Einstein metrics on homogeneous spaces $G_{0}/K_{0}$, which is a particularly rich theory for $G_{0}$ compact. For example, it was shown in [BL23] that every non-compact manifold with a homogeneous Einstein metric of negative scalar curvature must be diffeomorphic to $\mathbb{R}^{n}$.
Our key initial step is to compute the curvature of a metric on a homogeneous superspace $G/K$. Building on this, we solve (1.1) on a large class of such spaces, and find examples of homogeneous supermanifolds on which (1.1) has no solutions, (up to four) discrete families of solutions, and continuous families of solutions among invariant metrics. The solutions we produce appear to be the first nontrivial Einstein metrics obtained in the framework of supergeometry without exploiting methods or ideas from Kähler geometry. The existence of the discrete families is to be expected in light of the results available in the non-super setting. However, the presence of continuous families (constructed on $G/K$ with $G=\operatorname{SU}(m|n)$) shows that the well-known finiteness conjecture from homogeneous geometry (see [BWZ04]) fails on supermanifolds. Remarkably, all metrics in the continuous families are Ricci-flat, while Bochner’s theorem tells us that no Ricci-flat metrics exist on $G_{0}/K_{0}$ if $G_{0}$ is a compact Lie group whose identity component is not a torus. Thus, the exploration of links between the geometry of a supermanifold and the topology of its associated reduced manifold is an intriguing research direction that is likely to challenge intuition based on classical results only.
Often motivated by physical theories, homogeneous superspaces, including special types such as flag supermanifolds, have been studied before; see [BV91, Cor00, Goe08, San10, CCF11, Vis15, Vis22], and references therein. In fact, one may view our work as a continuation of these papers, offering a further but distinct study of the nature of homogeneous superspaces with distinguished structures.
The paper is organised as follows. In Section 2, we recall some definitions and properties of Lie superalgebras and their real forms, Lie supergroups, Harish-Chandra pairs, and homogeneous superspaces. In Section 3.1 and Section 3.2, we collect definitions and basic results from Riemannian supergeometry needed in subsequent sections. We also show that every left-invariant metric on $\operatorname{SL}(1|1)$ satisfies (1.1). This gives us, arguably, the simplest nontrivial examples of Einstein metrics on supermanifolds and a toy model for more complicated cases. In Section 3.3, we present our formulas for the curvature tensors associated with a metric on a homogeneous superspace, with the proof of the Ricci curvature formula deferred until Section 3.4. While the expression for the Riemann curvature is considerably bulkier than its non-super counterpart, we find that the expression for the Ricci curvature is comparatively simple. In both expressions, a key difference from the non-super setting stems from the use of dual bases. Specialisations and simplifications for naturally reductive and diagonal metrics are presented in Section 3.5 and Section 3.6, respectively. The possible non-positivity of superdimensions and structure constants presents additional challenges in the diagonal case.
Section 4 is devoted to Einstein metrics on homogeneous superspaces $G/K$, where $G$ is a connected Lie supergroup with Lie superalgebra $\mathfrak{g}$. We assume that $\mathfrak{g}$ is basic classical to ensure that $\mathfrak{g}^{\mathbb{C}}$ admits a non-degenerate $\operatorname{ad}_{\mathfrak{g}^{\mathbb{C}}}$-invariant supersymmetric even bilinear form, which we use to obtain our background metric on $G/K$. Furthermore, we demand that $\mathfrak{g}$ be the compact real form of $\mathfrak{g}^{\mathbb{C}}$ induced by a so-called star operation, as is the case in the non-super setting when $G_{0}$ is compact. This leaves $\mathfrak{g}^{\mathbb{C}}$ to be of type $A$ or $C$, whence $G=\operatorname{SU}(m|n)$ or $G=\operatorname{SOSp}(2|2n)$; see [SNR77, AN15, FG23]. Imitating the structure of so-called generalised flag manifolds in classical Riemannian geometry (see [Arv03]), we will assume that the Dynkin diagram of $K$ is obtained from that of $G$ by removing one or more nodes. Here, we restrict ourselves to removing one or two nodes only but intend to discuss elsewhere the removal of an arbitrary number of nodes. Section 4.1 thus outlines the construction of homogeneous superspaces by means of Dynkin diagrams, while Section 4.2 concerns the corresponding flag supermanifolds and their compact real forms. In Section 4.3 and Section 4.4, we obtain our classification results for Einstein metrics on $G/K$ with $G=\operatorname{SU}(m|n)$. Notably, some of the metrics we find are Ricci-flat, and our construction in Section 4.3 produces supergeometric analogues of Fubini–Study metrics on projective spaces. Section 4.5 concerns Einstein metrics on $G/K$ with $G=\operatorname{SOSp}(2|2n)$.
Notation
Throughout, the ground field is the field $\mathbb{R}$ of real numbers, unless otherwise stated. We let $\mathbb{C}$ denote the field of complex numbers, $\mathbb{H}$ the quaternions (appearing in Section 4.2.2), $\mathbb{Z}$ the ring of integers, $\mathbb{Z}_{2}=\{\bar{0},\bar{1}\}$ the ring of integers modulo 2, $\mathbb{R}^{+}$ (respectively $\mathbb{R}^{-}$) the set of positive (respectively negative) real numbers, and $\imath$ the imaginary unit, and set $\mathbb{R}^{\times}=\mathbb{R}-\{0\}$. Let $\langle-,$-$\rangle$ denote a scalar superproduct on a vector superspace $V$. If $\{v_{1},\ldots,v_{d}\}$ is a basis for $V$, we write $\{\bar{v}_{1},\ldots,\bar{v}_{d}\}$ for the corresponding right dual basis, i.e., $\langle v_{i},\bar{v}_{j}\rangle=\delta_{ij}$ for all $i,j\in\{1,\ldots,d\}$. We denote by $g$ a graded Riemannian metric on a Riemannian supermanifold $M$. We write $\mathfrak{g}^{\mathbb{C}}$ for a complex Lie superalgebra, with $\mathfrak{g}$ its real form over $\mathbb{R}$. If $\mathfrak{g}^{\mathbb{C}}$ is a basic classic complex Lie superalgebra, we denote by $Q$ a fixed, non-degenerate, supersymmetric, invariant, even bilinear form on $\mathfrak{g}^{\mathbb{C}}$.
2 Supermanifolds
Here, we recall some basic notions from supergeometry. We refer to [Kos77, Bat79, Lei80, BBH91, DeW92, DM99, Var04, Goe08, CCF11] for further background, noting (as in [BCF10]) that these references employ different approaches to supergeometry. Focus here is on notions and results needed in subsequent sections.
2.1 Lie superalgebras and real forms
A vector superspace of dimension $m|n$ is a $\mathbb{Z}_{2}$-graded vector space $V=V_{\bar{0}}\oplus V_{\bar{1}}$, where $m=\dim(V_{\bar{0}})$ and $n=\dim(V_{\bar{1}})$. The superdimension of $V$ is defined as $\operatorname{sdim}(V)=m-n$. A nonzero element of $V_{\bar{0}}\cup V_{\bar{1}}$ is said to be homogeneous, with elements of $V_{\bar{0}}$ called even and those of $V_{\bar{1}}$ odd, and a basis $\{v_{1},\dots,v_{m+n}\}$ for $V$ is said to be homogeneous if every $v_{i}$ is homogeneous. A homogeneous vector $v\in V_{a}$ is said to have parity $[v]=a$. For the tensor product of vector superspaces $V$ and $W$, we have
$(V\otimes W)_{\bar{0}}=(V_{\bar{0}}\otimes W_{\bar{0}})\oplus(V_{\bar{1}}\otimes W_{\bar{1}}),\qquad(V\otimes W)_{\bar{1}}=(V_{\bar{0}}\otimes W_{\bar{1}})\oplus(V_{\bar{1}}\otimes W_{\bar{0}}).$ (2.1)
Throughout this paper, definitions and calculations for vector superspaces are given in terms of homogeneous elements and then extended linearly. Moreover, the ground field $\mathbb{R}$ is considered an even vector space over $\mathbb{R}$, and a linear map $f\in\mathrm{Hom}_{\mathbb{R}}(V,W)$ is said to be even if $f(V_{a})\subseteq W_{a}$ for $a\in\mathbb{Z}_{2}$. This notion extends readily to bilinear maps.
A Lie superalgebra is a $\mathbb{Z}_{2}$-graded vector space (over $\mathbb{R}$ or $\mathbb{C}$) equipped with a Lie superbracket [,][-,-] which is super-skew-symmetric and satisfies the super-Jacobi identity. The (even bilinear) Killing form is defined for every finite-dimensional Lie superalgebra $\mathfrak{g}$, by
$B(X,Y):=\operatorname{Str}_{\mathfrak{g}}\!\big(\!\operatorname{ad}(X)\operatorname{ad}(Y)\big),\qquad X,Y\in\mathfrak{g},$
where $\operatorname{Str}_{\mathfrak{g}}$ is the supertrace over $\mathrm{End}(\mathfrak{g})$. The Killing form may be identically zero.
A finite-dimensional complex Lie superalgebra $\mathfrak{g}^{\mathbb{C}}=\mathfrak{g}_{\bar{0}}^{\mathbb{C}}\oplus\mathfrak{g}_{\bar{1}}^{\mathbb{C}}$ is called basic classical if $\mathfrak{g}^{\mathbb{C}}$ is simple, $\mathfrak{g}_{\bar{0}}^{\mathbb{C}}$ is reductive, and $\mathfrak{g}^{\mathbb{C}}$ admits a non-degenerate $\operatorname{ad}_{\mathfrak{g}^{\mathbb{C}}}$-invariant supersymmetric even bilinear form. Following Kac [Kac77, Kac78], there are two types of basic classical Lie superalgebras, characterised by whether $\mathfrak{g}_{\bar{1}}^{\mathbb{C}}$ is reducible (type I) or irreducible (type II) as a $\mathfrak{g}_{\bar{0}}^{\mathbb{C}}$-module. Thus, $A(m,n)$ and $C(n)$ comprise the basic classical Lie superalgebras of type I, while $B(m,n)$, $D(m,n)$, $F(4)$, $G(3)$ and $D(2,1,\alpha)$ comprise the ones of type II. Concretely, a basic classical Lie superalgebra of type I is isomorphic to one of the following Lie superalgebras: $A(m,n)=\mathfrak{sl}(m+1|n+1)^{\mathbb{C}}$ with $m>n\geq 0$; $A(n|n)=\mathfrak{psl}(n+1|n+1)^{\mathbb{C}}$ with $n\geq 1$; or $C(n)=\mathfrak{osp}(2|2n-2)^{\mathbb{C}}$ with $n\geq 3$ (noting that $C(2)\cong A(1|0)$).
A real Lie superalgebra $\mathfrak{g}$ is called a real form of a complex Lie superalgebra $\mathfrak{g}^{\mathbb{C}}$ if $\mathfrak{g}^{\mathbb{C}}\cong\mathfrak{g}\otimes_{\mathbb{R}}\mathbb{C}$, while a star operation $\mathfrak{g}^{\mathbb{C}}\to\mathfrak{g}^{\mathbb{C}}$ is an even anti-linear map such that
$[X,Y]^{*}=[Y^{*},X^{*}],\qquad(X^{*})^{*}=X,$
for any $X,Y\in\mathfrak{g}^{\mathbb{C}}$; see [SNR77] or [DM99, §4.1–4.3]. Here, anti-linearity means that $(cX)^{*}=\bar{c}X^{*}$, where $\bar{c}$ is the complex conjugate of $c\in\mathbb{C}$. With $\sqrt{\imath}=e^{\frac{\pi}{4}\imath}$, a real form $\mathfrak{g}=\mathfrak{g}_{\bar{0}}\oplus\mathfrak{g}_{\bar{1}}$ can now be constructed as
$\begin{gathered}\mathfrak{g}_{\bar{0}}=\operatorname{span}_{\mathbb{R}}\{X\in\mathfrak{g}_{\bar{0}}^{\mathbb{C}}\,|\,X^{*}=-X\},\\[2.84544pt] \mathfrak{g}_{\bar{1}}=\operatorname{span}_{\mathbb{R}}\{\sqrt{\imath}X\in\mathfrak{g}_{\bar{1}}^{\mathbb{C}}\,|\,X^{*}=-X\}=\operatorname{span}_{\mathbb{R}}\{X\in\mathfrak{g}_{\bar{1}}^{\mathbb{C}}\,|\,X^{*}=\imath X\}.\end{gathered}$ (2.4)
Note that $\mathfrak{g}_{\bar{0}}$ is a real form of the complex Lie algebra $\mathfrak{g}_{\bar{0}}^{\mathbb{C}}$. We say that a real form of $\mathfrak{g}^{\mathbb{C}}$ is compact if $\mathfrak{g}_{\bar{0}}$ is a compact Lie algebra; cf. [AN15, FG23]. It was shown in [SNR77] that among basic classical Lie superalgebras, only type I Lie superalgebras admit star operations. In that case, there are actually two star operations, and they are related by duality. Throughout this paper, we are only concerned with the so-called type (1) star operation; see, e.g., [GZ90a, GZ90b] and Section 4.2.
2.2 Lie supergroups
The approach taken when introducing the notion of a supermanifold is typically context-dependent. In the following, we largely follow [Kos77, Lei80].
An archetypal supermanifold is the superspace
$\mathbb{R}^{m|n}=(\mathbb{R}^{m},\mathcal{O}_{\mathbb{R}^{m|n}}),$
the topological space $\mathbb{R}^{m}$ equipped with the sheaf of supercommutative superalgebras
$\mathcal{O}_{\mathbb{R}^{m|n}}(U)=\mathcal{C}_{\mathbb{R}^{m}}^{\infty}(U)\otimes\Lambda_{\mathbb{R}}[\xi_{1},\dots,\xi_{n}]$
for any open set $U\subseteq\mathbb{R}^{m}$. Here, $\mathcal{C}_{\mathbb{R}^{m}}^{\infty}$ denotes the sheaf of smooth (even) functions on $\mathbb{R}^{m}$, and $\Lambda_{\mathbb{R}}[\xi_{1},\dots,\xi_{n}]$ is the Grassman algebra in the odd variables $\xi_{1},\dots,\xi_{n}$.
A (real smooth) supermanifold of dimension $m|n$ is a ringed space $M=(|M|,\mathcal{O}_{M})$, where the underlying topological space $|M|$ is Hausdorff and second-countable, and $\mathcal{O}_{M}=(\mathcal{O}_{M})_{\bar{0}}\oplus(\mathcal{O}_{M})_{\bar{1}}$ is a sheaf of superfunctions, i.e., for each $x\in|M|$ there exists an open neighbourhood $V_{x}\subseteq|M|$ of $x$ and an open set $U\subseteq\mathbb{R}^{m|n}$ such that $\mathcal{O}_{M}(V_{x})\cong\mathcal{O}_{\mathbb{R}^{m|n}}(U)$. Let $\mathcal{I}_{M}=\langle(\mathcal{O}_{M})_{\bar{1}}\rangle$ be the two-sided ideal of $\mathcal{O}_{M}$ generated by all nilpotent sections in $(\mathcal{O}_{M})_{\bar{1}}$. The corresponding subspace $M_{0}=(|M|,\mathcal{O}_{M}/\mathcal{I}_{M})$ is called the reduced manifold associated with $M$, where $\mathcal{O}_{M}/\mathcal{I}_{M}\cong\mathcal{C}^{\infty}_{|M|}$ is regarded as the sheaf of smooth functions on $|M|$. Hence, $M_{0}$ lies in the category of ordinary smooth manifolds. The canonical projection $\mathrm{ev}:\mathcal{O}_{M}\to\mathcal{O}_{M}/\mathcal{I}_{M}$ is the corresponding evaluation map. For any point $p\in|M|$, let $\mathcal{O}_{M,p}$ denote the stalk at $p$. The evaluation at $p$ is then defined as
$\mathrm{ev}_{p}:\mathcal{O}_{M,p}\to\mathbb{R},\qquad f\mapsto f(p):=\mathrm{ev}(f)(p).$
For each supermanifold $M=(|M|,\mathcal{O}_{M})$, the tangent sheaf $\mathcal{T}_{M}$ is defined by
$\mathcal{T}_{M}(U):=\operatorname{Der}(\mathcal{O}_{M}(U))$
for any open subset $U\subseteq|M|$. Here, $\operatorname{Der}(\mathcal{O}_{M}(U))$ denotes the $\mathcal{O}_{M}(U)$-module of (left) super-derivations $\varphi$, the ones satisfying the super-Leibniz rule
$\varphi(fg)=\varphi(f)g+(-1)^{[\varphi][f]}f\varphi(g),\qquad f,g\in\mathcal{O}_{M}(U).$
The sections of $\mathcal{T}_{M}$ are called vector fields. For every point $p\in|M|$, the tangent superspace $T_{p}M$ at $p$ is the vector superspace of dimension $m|n$, defined by
$T_{p}M:=\{\psi\in\mathrm{Hom}_{\mathbb{R}}(\mathcal{O}_{M,p},\mathbb{R})\,|\,\psi(fg)=\psi(f)g(p)+(-1)^{[\psi][f]}f(p)\psi(g)\}.$
Elements of $T_{p}M$ are called tangent vectors. For any $p\in U$ and vector field $X\in\mathcal{T}_{M}(U)$, the value of $X$ at $p$ is the tangent vector $X_{p}:\mathcal{O}_{M,p}\to\mathbb{R}$ defined by $X_{p}(f):=\mathrm{ev}_{p}(X(f))$ for all $f\in\mathcal{O}_{M,p}$. A vector field is only determined by its values at all points if the odd dimension of $M$ is zero. We also introduce the dual space
$\mathcal{T}_{M}^{*}:=\mathrm{Hom}_{\mathcal{O}_{M}}(\mathcal{T}_{M},\mathcal{O}_{M}).$
A Lie supergroup $G$ is a group object in the category of supermanifolds, that is, it is a real smooth supermanifold together with three morphisms
$m:G\times G\to G,\qquad i:G\to G,\qquad e:\mathbb{R}^{0|0}\to G,$
called multiplication, inverse and unit, respectively, satisfying the usual group axioms that encode associativity and the existence of identity and of inverse (see [CCF11, Definition 7.1]). The reduced manifold $G_{0}$ is an ordinary Lie group. A Lie supergroup $K$ is a Lie subsupergroup of $G$ if the reduced group $K_{0}$ is a Lie subgroup of $G_{0}$ and the inclusion morphism $K\to G$ is an even immersion. Moreover, $K$ is said to be closed, respectively connected, if $K_{0}$ is closed, respectively connected, in $G_{0}$.
A vector field $X$ on $G$ is left-invariant if and only if
$(1\otimes X)\circ m^{*}=m^{*}\circ X,$
where $m^{*}:\mathcal{O}_{G}\to\mathcal{O}_{G\times G}$ is a morphism on sheaves. The Lie superalgebra $\mathfrak{g}$ of $G$ consists of all left-invariant vector fields on $G$. We can identify $\mathfrak{g}$ with the tangent superspace of $G$ at the identity $e$, i.e., the linear map $\mathfrak{g}\to T_{e}G$, $X\mapsto X_{e}$, is an isomorphism of vector superspaces, with inverse map given by
$v\mapsto X_{v}=(1\otimes v)\circ m^{*},$
where $X_{v}$ is the left invariant vector field with $(X_{v})_{e}=v$. We refer to [CCF11] for more details.
2.3 Harish-Chandra pairs
Complementary to the introduction in Section 2.2, the theory of Lie supergroups can be approached equivalently through Harish-Chandra pairs [Kos77, DM99, CCF11].
Definition 2.1.
A (real) Harish-Chandra pair $(G_{0},\mathfrak{g})$ consists of an ordinary Lie group $G_{0}$; a Lie superalgebra $\mathfrak{g}=\mathfrak{g}_{\bar{0}}\oplus\mathfrak{g}_{\bar{1}}$, where $\mathfrak{g}_{\bar{0}}$ is the Lie algebra of $G_{0}$; and a representation $\sigma$ of $G_{0}$ on $\mathfrak{g}$ such that (i) $\sigma$ induces the adjoint representation of $G_{0}$ on $\mathfrak{g}_{\bar{0}}$, and (ii) the differential $(d\sigma)_{e}$ at the identity $e\in G_{0}$ acts on $\mathfrak{g}$ as the adjoint representation of $\mathfrak{g}_{\bar{0}}$.
Given a Lie supergroup $G$, we can define a Harish-Chandra pair $(G_{0},\mathfrak{g})$ by taking $G_{0}$ to be the underlying reduced Lie group of $G$, and $\mathfrak{g}$ to be the Lie superalgebra of $G$. It is well known that the category of (real) Harish-Chandra pairs is equivalent to the category of (real) Lie supergroups; see, e.g., [CCF11, Theorem 7.4.5]. Thus, a Lie supergroup only depends on the underlying ordinary Lie group and Lie superalgebra, together with the compatibility conditions given in Definition 2.1.
As an example, the general linear supergroup $\operatorname{GL}(m|n)$ is the Lie supergroup associated with the Harish-Chandra pair $(\operatorname{GL}(m)\times\operatorname{GL}(n),\mathfrak{gl}(m|n))$, where $\mathfrak{gl}(m|n)$ is the real general linear superalgebra comprising the endomorphisms on $\mathbb{R}^{m|n}$. The special linear supergroup $\operatorname{SL}(m|n)=(\operatorname{SL}(m|n)_{0},\mathfrak{sl}(m|n))$ is the Lie subsupergroup of $\operatorname{GL}(m|n)$ with even subsupergroup
$\operatorname{SL}(m|n)_{0}=\{(A,B)\in\operatorname{GL}(m)\times\operatorname{GL}(n)\,|\,\mathrm{det}(A)=\mathrm{det}(B)>0\}$
and Lie superalgebra
$\mathfrak{sl}(m|n)=\{X\in\mathfrak{gl}(m|n)\,|\,\operatorname{Str}(X)=0\}.$
A representation of a Harish-Chandra pair $G=(G_{0},\mathfrak{g})$ on a vector superspace $V$ consists of a representation of $G_{0}$ on $V$ and one of $\mathfrak{g}$ on $V$, such that the differential of the representation of $G_{0}$ agrees with the restriction to $\mathfrak{g}_{0}$ of the $\mathfrak{g}$-representation.
2.4 Homogeneous superspaces
Let $G$ be a Lie supergroup in the category of real smooth supermanifolds and $K$ a closed Lie subsupergroup of $G$. Classically, the set $G_{0}/K_{0}$ is a topological space with the quotient topology induced by the natural projection $\pi_{0}:G_{0}\to G_{0}/K_{0}$. Moreover, it has a unique classical manifold structure compatible with the action of $G_{0}$. As discussed in the following, this classical construction can be extended to the super-setting, thereby giving rise to homogeneous superspaces.
Following [Kos77, Theorem 3.9], the super-structure sheaf $\mathcal{O}_{G/K}$ on $G_{0}/K_{0}$ is defined as follows. Let $\mathrm{pr}_{1}:G\times K\to G$ denote the projection onto the first factor, and $\psi:G\times K\to G$ the morphism induced by the multiplication $m$. For any open subset $U\subseteq G_{0}/K_{0}$, define the set of $K$-invariant sections by
$\mathcal{O}_{G/K}(U):=\{f\in\mathcal{O}_{G}(\pi_{0}^{-1}(U))\,|\,\psi^{*}f=\mathrm{pr}_{1}^{*}f\}.$
This gives rise to a supermanifold,
$G/K:=(G_{0}/K_{0},\mathcal{O}_{G/K}),$
referred to as a homogeneous superspace. Let $\pi:G\to G/K$ be the corresponding canonical surjective homomorphism of supermanifolds, where the sheaf morphism is given by inclusion. The differential of $\pi$ at the identity $e\in G$ yields a surjective linear map $\mathfrak{g}\cong T_{e}G\to T_{K}(G/K)$ with kernel $\mathfrak{k}$, the Lie superalgebra of $K$. This induces a canonical identification: $T_{K}(G/K)\cong\mathfrak{g}/\mathfrak{k}$.
A Harish-Chandra pair $G=(G_{0},\mathfrak{g})$ acts on a supermanifold $M$ if the reduced Lie group $G_{0}$ acts on $|M|$, and there exists a Lie superalgebra anti-homomorphism from $\mathfrak{g}$ to the Lie superalgebra $\mathcal{T}_{M}(M)$ of vector fields on $M$ such that its restriction to $\mathfrak{g}_{\bar{0}}$ agrees with the differential of the $G_{0}$-action at the identity [DM99, §3.8]. Indeed, with $\alpha:G\times M\to M$ denoting the Lie supergroup action, the anti-homomorphism is given by
$\mathfrak{g}\to\mathcal{T}_{M}(M),\qquad X\mapsto(X_{e}\otimes 1)\circ\alpha^{*}.$ (2.5)
A $G$-action is said to be transitive if (i) the reduced $G_{0}$-action is transitive, and (ii) for every $p\in|M|$, the even linear map
$\mathfrak{g}\to T_{p}M,\qquad X\mapsto\mathrm{ev}_{p}\circ(X_{e}\otimes 1)\circ\alpha^{*},$ (2.6)
is surjective [Goe08, Section 4.11].
A supermanifold $M$ is said to be $G$-homogeneous if the Lie supergroup $G$ acts on it transitively. Every $G$-homogeneous supermanifold $M$ is isomorphic to $G/G_{p}$ with $G_{p}$ being the isotropy subsupergroup of $G$ at $p\in|M|$ [Kos77, Proposition 3.10.3]. By definition, the isotropy subsupergroup $G_{p}$ is the Harish-Chandra pair $((G_{0})_{p},\mathfrak{g}_{p})$, where $(G_{0})_{p}$ denotes the ordinary isotropy subgroup of $G_{0}$ at $p$, and $\mathfrak{g}_{p}$ consists of all left-invariant vector fields $X$ on $G$ such that the infinitesimal action of $X$ at $p$ is trivial; see [Goe08, Section 4.11] and [CCF11, §8.4]. The isotropy representation of $G_{p}$ on $T_{p}M$ is equivalent to the adjoint representation on $\mathfrak{g}/\mathfrak{g}_{p}$ via the natural isomorphism $\mathfrak{g}/\mathfrak{g}_{p}\cong T_{p}M$ [Goe08, Proposition 12].
3 Curvature
3.1 Graded Riemannian metrics and curvatures
Let $V=V_{\bar{0}}\oplus V_{\bar{1}}$ be a vector superspace over $\mathbb{R}$. A scalar superproduct on $V$ is a non-degenerate, supersymmetric, even bilinear form $\langle-,-\rangle:V\times V\rightarrow\mathbb{R}$. By the non-degeneracy, $\dim(V_{\bar{1}})$ must be even. The supersymmetry implies that $\langle-,$-$\rangle$ is symmetric on $V_{\bar{0}}\times V_{\bar{0}}$, and skew-symmetric on $V_{\bar{1}}\times V_{\bar{1}}$, while the evenness implies that it is zero on $V_{\bar{0}}\times V_{\bar{1}}$ and $V_{\bar{1}}\times V_{\bar{0}}$. Moreover, the restriction of $\langle-,$-$\rangle$ to $V_{\bar{0}}$ may be indefinite.
Definition 3.1.
[Goe08] A graded Riemannian metric on a supermanifold $M=(|M|,\mathcal{O}_{M})$ is an $\mathcal{O}_{M}$-linear morphism of sheaves $g:\mathcal{T}_{M}\otimes\mathcal{T}_{M}\to\mathcal{O}_{M}$ such that for any $X,Y\in\mathcal{T}_{M}(U)$, where $U$ is any open subset of $|M|$,
  1. (1)
    (non-degeneracy) the map $X\mapsto g(X,-)$ is an isomorphism $\mathcal{T}_{M}\rightarrow\mathcal{T}_{M}^{*}$;
  2. (2)
    (supersymmetry) $g(X,Y)=(-1)^{[X][Y]}g(Y,X)$;
  3. (3)
    (evenness) $[g(X,Y)]=[X]+[Y]$.
For each $p\in|M|$, the sheaf morphism $g$ induces a scalar superproduct $\langle-,-\rangle_{p}:=g_{p}$ on the tangent superspace $T_{p}M$,
$\langle-,-\rangle_{p}:T_{p}M\times T_{p}M\to\mathbb{R}.$
If the restriction of $\langle-,-\rangle_{p}$ to $(T_{p}M)_{\bar{0}}$ is indefinite, it induces a pseudo-Riemannian metric on the reduced manifold $M_{0}$. We stress that the scalar superproduct $\langle-,-\rangle_{p}$ does not determine the graded Riemannian metric unless the odd dimension of $M$ is zero. A supermanifold $M$ equipped with a graded Riemannian metric $g$ is called a Riemannian supermanifold, and is denoted by $(M,g)$.
An affine connection on a supermanifold $M$ is an even morphism $\nabla:\mathcal{T}_{M}\otimes_{\mathbb{R}}\mathcal{T}_{M}\rightarrow\mathcal{T}_{M}$ of sheaves of vector superspaces such that, for every open subset $U\subseteq|M|$,
$\nabla_{fX}Y=f\nabla_{X}Y,\qquad\nabla_{X}(fY)=X(f)Y+(-1)^{[X][f]}f\nabla_{X}Z,$
where $f\in\mathcal{O}_{M}(U)$ and $X,Y\in\mathcal{T}_{M}(U)$. If $(M,g)$ is a Riemannian supermanifold, there is a unique affine connection, called the Levi-Civita connection, satisfying
$Xg(Y,Z)$ $=g(\nabla_{X}Y,Z)+(-1)^{[X][Y]}g(Y,\nabla_{X}Z),$ (3.1)
$[X,Y]$ $=\nabla_{X}Y-(-1)^{[X][Y]}\nabla_{Y}X,$ (3.2)
for all vector fields $X,Y,Z$ on $M$, meaning that the connection is metric and torsion-free. As in the non-super case, the Levi-Civita connection is determined by the following super analogue of the Koszul formula.
Proposition 3.2.
[MS96, Theorem 4.2] Let $(M,g)$ be a Riemannian supermanifold. Then, the Levi-Civita connection $\nabla$ is given implicitly by the Koszul formula
$2\,g(\nabla_{X}Y,Z)$ $=Xg(Y,Z)+(-1)^{[X][Y]}Yg(X,Z)-(-1)^{([X]+[Y])[Z]}Zg(X,Y)$
$\quad-g(X,[Y,Z])-(-1)^{[X][Y]}g(Y,[X,Z])+(-1)^{([X]+[Y])[Z]}g(Z,[X,Y]),$
with $X,Y,Z$ any vector fields on $M$.
A graded Killing vector field on a Riemannian supermanifold is a vector field $X$ that preserves the metric, that is,
$Xg(Y,Z)=g([X,Y],Z)+(-1)^{[X][Y]}g(Y,[X,Z]),$ (3.3)
where $Y,Z$ are any vector fields, and we note that the set of graded Killing vector fields forms a Lie superalgebra. Using (3.1) and (3.2), a vector field $X$ is seen to be a graded Killing vector field if and only if
$g(\nabla_{Y}X,Z)+(-1)^{[X][Z]}g(Y,\nabla_{Z}X)=0$ (3.4)
for all vector fields $Y,Z$. For graded Killing vector fields, the Koszul formula in Proposition 3.2 readily simplifies as follows.
Lemma 3.3.
Let $X,Y,Z$ be graded Killing vector fields on a Riemannian supermanifold $(M,g)$. Then,
$2g(\nabla_{X}Y,Z)=g(X,[Y,Z])+Xg(Y,Z)=g(X,[Y,Z])+g([X,Y],Z)+(-1)^{[X][Y]}g(Y,[X,Z]).$
Following [DeW92], the Riemann curvature tensor on a Riemannian supermanifold $(M,g)$ is defined by
$R(X,Y)Z:=\nabla_{[X,Y]}Z-\nabla_{X}\nabla_{Y}Z+(-1)^{[X][Y]}\nabla_{Y}\nabla_{X}Z,$
where $X,Y,Z\in\mathcal{T}_{M}(U)$ are any vector fields and $U$ is any open subset of $|M|$. Furthermore, the Ricci curvature tensor is defined as
$\mathrm{Ric}(Y,Z)=\mathrm{Str}\big(X\mapsto(-1)^{[X][Z]}R(Y,X)Z\big),$
where the supertrace $\mathrm{Str}$ is taken over the $\mathcal{O}_{M}(U)$-linear superspace $\mathcal{T}_{M}(U)$. Similarly, the scalar curvature is defined as the supertrace of the Ricci curvature tensor with respect to the metric $g$.
In the remainder of this paper, we will work with these curvature tensors for homogeneous superspaces. To present the curvature tensors evaluated at a point $p\in|M|$, let $\{X_{1},\dots,X_{d}\}$ be an ordered homogeneous basis for the tangent superspace $T_{p}(M)$, with $\{\widebar{X}_{1},\ldots,\widebar{X}_{d}\}$ the corresponding right dual basis. For a vector $X\in T_{p}(M)$, we thus have
$X=\sum_{i=1}^{d}\,\langle X,\widebar{X}_{i}\rangle_{p}X_{i}=\sum_{i=1}^{d}\,\langle X_{i},X\rangle_{p}\widebar{X}_{i}.$ (3.5)
At $p\in|M|$, the Ricci curvature is given by
$\operatorname{Ric}(X_{i},X_{k})_{p}:=\sum_{j=1}^{d}(-1)^{[X_{j}]+[X_{j}][X_{k}]}\langle R(X_{i},X_{j})X_{k},\widebar{X}_{j}\rangle_{p}.$ (3.6)
As $\operatorname{Ric}(-,-)_{p}$ is even, we have $\operatorname{Ric}(X_{i},X_{j})_{p}=0$ if $X_{i}]\neq[X_{j}]$. Finally, the scalar curvature at $p$ is
$S(p):=\sum_{i=1}^{d}(-1)^{[X_{i}]}\operatorname{Ric}(X_{i},\widebar{X}_{i})_{p}.$ (3.7)
3.2 Invariant metrics
Let $G=(G_{0},\mathfrak{g})$ be a Lie supergroup. A graded Riemannian metric $g$ on $G$ is called left-invariant if $g(X,Y)$ is a constant function on $G$ for all left-invariant vector fields $X,Y$. Recall that we identify the Lie superalgebra $\mathfrak{g}$ of left-invariant vector fields with the tangent space $T_{e}G$ at the identity $e\in G$ via the isomorphism $X\mapsto X_{e}$. Similarly to the non-super case, the left-invariant graded Riemannian metrics $g$ on $G$ are in natural bijection with the scalar superproducts $\langle-,-\rangle:=g_{e}$ on $\mathfrak{g}\cong T_{e}G$ [Goe08, Section 4.9].
Lemma 3.4.
Let $g$ be a left-invariant graded Riemannian metric on $(G_{0},\mathfrak{g})$. Then,
$2\,g(\nabla_{X}Y,Z)=g([X,Y],Z)-g(X,[Y,Z])-(-1)^{[X][Y]}g(Y,[X,Z]),\qquad X,Y,Z\in\mathfrak{g}.$
Proof.
Since $g$ is left-invariant, $Xg(Y,Z)=0$ for all $X,Y,Z\in\mathfrak{g}$, so the result follows from Proposition 3.2. ∎
A scalar superproduct $\langle-,$-$\rangle$ on a Lie superalgebra $\mathfrak{g}\cong T_{e}G$ is called $\operatorname{Ad}_{G}$-invariant if it is both $\operatorname{Ad}_{G_{0}}$- and $\operatorname{ad}_{\mathfrak{g}}$-invariant, i.e., for every $X,Y,Z\in\mathfrak{g}$ and every $a\in G_{0}$, we have
$\langle\operatorname{Ad}_{a}(X),\operatorname{Ad}_{a}(Y)\rangle=\langle X,Y\rangle,\qquad\langle[X,Y],Z\rangle=\langle X,[Y,Z]\rangle.$
Note that if $G$ is connected, then $\operatorname{ad}_{\mathfrak{g}}$-invariance implies $\operatorname{Ad}_{G_{0}}$-invariance.
Definition 3.5.
A graded Riemannian metric on a $G$-homogeneous superspace $M$ is $G$-invariant if (i) every element of $G_{0}$ acts on $M$ by isometries, and (ii) the image of the anti-homomorphism (2.5) is contained in the Lie superalgebra of graded Killing vector fields.
Theorem 3.6.
[Goe08, Theorem 3] Let $G=(G_{0},\mathfrak{g})$ be a Lie supergroup and $K=(K_{0},\mathfrak{k})$ a closed Lie subsupergroup of $G$. The map $g\mapsto g_{K}$ is a bijective correspondence between $G$-invariant graded Riemannian metrics on $G/K$ and $\operatorname{Ad}_{K}$-invariant scalar superproducts on the tangent superspace $T_{K}(G/K)$ at the base point $K$.
Theorem 3.6 implies that every $G$-invariant graded Riemannian metric $g$ on $G/K$ is uniquely determined by its evaluation $g_{K}$ at the base point $K$. Using the method in the proof of this result in [Goe08], one shows that $\mathrm{Ric}(g)$ on $G/K$ is uniquely determined by $\mathrm{Ric}(g)_{K}$, even though the Ricci tensor may be degenerate. Here, $\mathrm{Ric}(g)_{K}$ is a supersymmetric, $\mathrm{Ad}_{K}$-invariant even bilinear form on $T_{K}(G/K)$. More generally [San10, Theorem 4.16], there is a bijective correspondence between $G$-invariant tensor fields of type $(r,s)$ on $G/K$ and isotropy-invariant tensors of type $(r,s)$ on $T_{K}(G/K)$, via evaluation at $K$. In our case, both $g$ and $\mathrm{Ric}(g)$ are $G$-invariant tensor fields of type $(0,2)$. Thus, to solve the Einstein equation $\mathrm{Ric}(g)=cg$ on a homogeneous Riemannian supermanifold $(G/K,g)$, it is necessary and sufficient to solve the corresponding equation at the base point $K$, that is, to find $g_{K}$ satisfying $\mathrm{Ric}(g)_{K}=cg_{K}$. The bijective correspondence ensures that a solution at $K$ uniquely extends to a $G$-invariant solution on the entire supermanifold.
Example 3.7.
In the spirit of [Mil76] on left-invariant metrics on Lie groups, we now compute the Ricci curvature of the three-dimensional Lie supergroup $\operatorname{SL}(1|1)$, and demonstrate that every left-invariant graded Riemannian metric on $\operatorname{SL}(1|1)$ is an Einstein metric. Although $\operatorname{SL}(1|1)$ is not part of our $A$-type classification in Section 4.2.1 and Section 4.4, we find that the analysis below serves to illustrate some of the key notions.
The Lie superalgebra $\mathfrak{sl}(1|1)$ of the Lie supergroup $\operatorname{SL}(1|1)$ has basis
$X_{1}=\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\qquad X_{2}=\begin{pmatrix}0&1\\ 0&0\end{pmatrix},\qquad X_{3}=\begin{pmatrix}0&0\\ 1&0\end{pmatrix}$,
with graded commutation relations
$[X_{1},X_{2}]=[X_{1},X_{3}]=0,\qquad[X_{2},X_{3}]=X_{1},$
where $X_{1}$ is even and $X_{2},X_{3}$ odd. Clearly, the centre of $\mathfrak{sl}(1|1)$ is spanned by $X_{1}$.
Let $g$ be a left-invariant graded Riemannian metric on $\operatorname{SL}(1|1)$. Then, $g$ can be identified with a scalar superproduct $\langle-,$-$\rangle$ on $\mathfrak{sl}(1|1)$, which is a non-degenerate supersymmetric even bilinear form. This implies that $\langle X_{1},X_{2}\rangle=\langle X_{1},X_{3}\rangle=\langle X_{2},X_{2}\rangle=\langle X_{3},X_{3}\rangle=0$ and $\langle X_{2},X_{3}\rangle=-\langle X_{3},X_{2}\rangle$, so relative to the ordered basis $\{X_{1},X_{2},X_{3}\}$, we may assume that the metric $g$ has the matrix form
$(g(X_{i},X_{k}))_{3\times 3}=\begin{pmatrix}x_{1}&0&0\\ 0&0&x_{2}\\ 0&-x_{2}&0\end{pmatrix},\qquad x_{1},x_{2}\in\mathbb{R}^{\times}.$
The corresponding right dual basis is thus given by $\{\widebar{X}_{1}=\frac{1}{x_{1}}X_{1},\widebar{X}_{2}=\frac{1}{x_{2}}X_{3},\widebar{X}_{3}=-\frac{1}{x_{2}}X_{2}\}$.
Using Lemma 3.4, we find $\nabla_{X_{1}}X_{1}=\nabla_{X_{2}}X_{2}=\nabla_{X_{3}}X_{3}=0$ and
$\nabla_{X_{1}}X_{2}=\nabla_{X_{2}}X_{1}=-\frac{x_{1}}{2x_{2}}X_{2},\qquad\nabla_{X_{1}}X_{3}=\nabla_{X_{3}}X_{1}=\frac{x_{1}}{2x_{2}}X_{3},\qquad\nabla_{X_{2}}X_{3}=\nabla_{X_{3}}X_{2}=\frac{1}{2}X_{1},$
from which it follows that the Ricci curvature at the identity $e\in\operatorname{SL}(1|1)$ is given by
$\operatorname{Ric}(X_{i},X_{k})=\frac{x_{1}}{2x_{2}^{2}}\,g(X_{i},X_{k}).$
We thus conclude that every left-invariant graded Riemannian metric on $\operatorname{SL}(1|1)$ is an Einstein metric.
3.3 Curvature formulas
Let $G=(G_{0},\mathfrak{g})$ be a connected Lie supergroup, with $\mathfrak{g}$ a compact real form of a basic classical Lie superalgebra $\mathfrak{g}^{\mathbb{C}}$. As $\mathfrak{g}^{\mathbb{C}}$ is basic classical, there exists a non-degenerate supersymmetric even bilinear form $Q$ on $\mathfrak{g}^{\mathbb{C}}$, with $\operatorname{ad}_{\mathfrak{g}^{\mathbb{C}}}$-invariance given by
$Q([X,Y],Z)=Q(X,[Y,Z]),\qquad X,Y,Z\in\mathfrak{g}^{\mathbb{C}}.$
Let $K=(K_{0},\mathfrak{k})$ be a closed connected Lie subsupergroup of $G$, and let the homogeneous superspace $M=G/K$ be equipped with a $G$-invariant graded Riemannian metric $g$. Our goal here is to give explicit formulas for the various curvatures of this metric at the point $K$ of $M$. As discussed following Theorem 3.6, these formulas determine the full curvature tensors.
Since $M$ is equipped with a $G$-invariant metric, the image under the anti-homomorphism (2.5) of the Lie superalgebra $\mathfrak{g}$ of left-invariant vector fields is contained in the Lie superalgebra of graded Killing fields, cf. Definition 3.5. As $\mathfrak{g}$ is simple, the anti-homomorphism, mapping $X\mapsto X^{\dagger}$, is injective, so we can identity the left-invariant vector field $X\in\mathfrak{g}$ with the Killing vector field $X^{\dagger}$ on $(M,g)$. With [,][-,-] denoting the Lie superbracket of vector fields on $M$, and $_{\mathfrak{g}}$ the Lie superbracket of $\mathfrak{g}$, the anti-homomorphicity then implies that
$[X^{\dagger},Y^{\dagger}]=-[X,Y]_{\mathfrak{g}}^{\dagger},\qquad X,Y\in\mathfrak{g}.$ (3.8)
We now identify the Lie subsuperlagebra $\mathfrak{k}$ of $\mathfrak{g}$ with the subsuperalgebra of the Killing vector fields $X^{\dagger}$ which vanish at the base point $K$, and choose once and for all a $Q$-orthogonal complement $\mathfrak{m}$ of $\mathfrak{k}$. We thus have the vector superspace decomposition $\mathfrak{g}=\mathfrak{k}\,\oplus\,\mathfrak{m}$ and ditto isomorphisms $\mathfrak{g}/\mathfrak{k}\cong\mathfrak{m}\cong T_{K}M$, where we identify $\mathfrak{m}$ with $T_{K}M$ by evaluating the corresponding graded Killing vector field at the point $K$. In this way, the isotropy representation of $K$ in $T_{K}M$ is identified with the restriction of the $\operatorname{Ad}_{G}$-representation of $K$ to $\mathfrak{m}$.
We denote by $\langle-,$-$\rangle$ the scalar superproduct on $\mathfrak{m}$ induced by $g_{K}$ through the identification of $\mathfrak{m}$ with $T_{K}M$. By Theorem 3.6, $\langle-,$-$\rangle$ is an $\mathrm{Ad}_{K}$-invariant, supersymmetric, even bilinear form on $\mathfrak{m}$. As $K$ is connected, the $\mathrm{Ad}_{K}$-invariance is equivalent to $\operatorname{ad}_{\mathfrak{k}}$-invariance, that is,
$\langle X,[Y,Z]_{\mathfrak{m}}\rangle=\langle[X,Y]_{\mathfrak{m}},Z\rangle,$
for all $X,Z\in\mathfrak{m}$ and $Y\in\mathfrak{k}$. We extend the superproduct $\langle-,$-$\rangle$ on $\mathfrak{m}$ to an $\operatorname{Ad}_{K}$-invariant superproduct on $\mathfrak{g}$ by choosing on $\mathfrak{k}$ an $\operatorname{Ad}_{K}$-invariant superproduct, such as $Q|_{\mathfrak{k}}$, and setting $\langle\mathfrak{k},\mathfrak{m}\rangle=0$. In the following, we primarily use $X,Y,Z,\ldots$ to denote elements of $\mathfrak{g}$, despite the use of the same symbols to represent vector fields in Section 3.1 and Section 3.2. We hope this practice will not cause any confusion.
The next result evaluates the Levi-Civita connection on $(M,g)$ at the base point $K$. We denote by $Z_{\mathfrak{k}}$, respectively $Z_{\mathfrak{m}}$, the $\mathfrak{k}$- and $\mathfrak{m}$-components of $Z\in\mathfrak{g}$, so
$[-,-]_{\mathfrak{g}}=[-,-]_{\mathfrak{k}}+[-,-]_{\mathfrak{m}}.$ (3.9)
We follow [Bes87, Proposition 7.28] on classical homogeneous spaces and introduce the even bilinear map $U:\mathfrak{m}\times\mathfrak{m}\to\mathfrak{m}$ given implicitly by
$\langle U(X,Y),Z\rangle=-\langle X,[Y,Z]_{\mathfrak{m}}\rangle-(-1)^{[X][Y]}\langle Y,[X,Z]_{\mathfrak{m}}\rangle,\qquad X,Y,Z\in\mathfrak{m}.$ (3.10)
It readily follows that $U(X,Y)=(-1)^{[X][Y]}U(Y,X)$ for any homogeneous elements $X,Y\in\mathfrak{m}$.
Remark 3.8.
To reduce the number of fractions in the ensuing expressions, especially in Proposition 3.12 and Section 3.4, we have opted to scale $U$ by 22 relative to the similar function in [Bes87].
Proposition 3.9.
Let $M=G/K$ be a homogeneous superspace with a $G$-invariant graded Riemannian metric $g$, and let $X^{\dagger},Y^{\dagger}$ be graded Killing vector fields associated with $X,Y\in\mathfrak{m}$. Then,
$2(\nabla_{X^{\dagger}}Y^{\dagger})_{K}=-[X,Y]_{\mathfrak{m}}+U(X,Y).$
Proof.
[…]
$2g(\nabla_{X^{\dagger}}Y^{\dagger},Z^{\dagger})=-g(X^{\dagger},[Y,Z]_{\mathfrak{g}}^{\dagger})-g([X,Y]^{\dagger}_{\mathfrak{g}},Z^{\dagger})-(-1)^{[X^{\dagger}][Y^{\dagger}]}g(Y^{\dagger},[X,Z]^{\dagger}_{\mathfrak{g}}).$
Evaluating at $K$ and using (3.10), having identified $\mathfrak{m}$ with $T_{K}M$, we obtain
$2\,\langle(\nabla_{X^{\dagger}}Y^{\dagger})_{K},Z\rangle=-\langle[X,Y]_{\mathfrak{m}},Z\rangle+\langle U(X,Y),Z\rangle.$
This completes the proof. ∎
We note that the restriction of $\langle-,$-$\rangle$ to $\mathfrak{m}_{\bar{0}}\times\mathfrak{m}_{\bar{0}}$ is symmetric but may be indefinite, that the restriction to $\mathfrak{m}_{\bar{1}}\times\mathfrak{m}_{\bar{1}}$ is skew-symmetric, that $\dim(\mathfrak{m}_{\bar{1}})$ is even, and that $\langle\mathfrak{m}_{\bar{0}},\mathfrak{m}_{\bar{1}}\rangle=\langle\mathfrak{m}_{\bar{1}},\mathfrak{m}_{\bar{0}}\rangle=0$. It follows that $\mathfrak{m}$ admits an ordered $g$-normalised homogeneous basis $\{X_{1},\dots,X_{c},X_{c+1},\dots,X_{d}\}$, where $X_{i}]=\bar{0}$ for $i\leq c$ and $X_{i}]=\bar{1}$ for $i>c$, such that the corresponding matrix representation of $\langle-,$-$\rangle$ is of the form
$(\langle X_{i},X_{j}\rangle)_{d\times d}=\begin{pmatrix}I_{p}&0&0&0\\ 0&-I_{q}&0&0\\ 0&0&0&I_{r}\\ 0&0&-I_{r}&0\end{pmatrix}$, (3.11)
where $p+q=c$ and 2r=dc2r=d-c. The right dual basis elements are given by
$\begin{array}[]{rllll}&\widebar{X}_{i}=X_{i},\quad&1\leq i\leq p;\qquad&\widebar{X}_{i}=X_{i+r},\quad&c<i\leq c+r,\\[5.69046pt] &\widebar{X}_{i}=-X_{i},\quad&p<i\leq c;\qquad&\widebar{X}_{i}=-X_{i-r},\quad&d-r<i\leq d.\end{array}$ (3.14)
We have the following useful observation.
Lemma 3.10.
Let $f$ be an even bilinear map on $\mathfrak{m}\times\mathfrak{m}$. Then,
$\sum_{j=1}^{d}f(\widebar{X}_{j},X_{j})=\sum_{j=1}^{d}(-1)^{[X_{j}]}f(X_{j},\widebar{X}_{j}).$
Proof.
Using (3.14), we have
$\sum_{j=1}^{d}f(\widebar{X}_{j},X_{j})$ $=\sum_{j=1}^{p}f(X_{j},X_{j})+\sum_{j=p+1}^{c}f(-X_{j},X_{j})+\sum_{j=c+1}^{c+r}f(X_{j+r},X_{j})+\sum_{j=c+r+1}^{d}f(-X_{j-r},X_{j})$
$=\sum_{j=1}^{p}f(X_{j},X_{j})+\sum_{j=p+1}^{c}f(X_{j},-X_{j})-\sum_{j=c+r+1}^{d}f(X_{j},-X_{j-r})-\sum_{j=c+1}^{d-r}f(X_{j},X_{j+r}),$
which is recognised as the righthand side of the stated relation. ∎
Remark 3.11.
Using the super-skew-symmetry of $_{\mathfrak{m}}$, it readily follows from Lemma 3.10 that
$\sum_{j=1}^{d}\,[\widebar{X}_{j},X_{j}]_{\mathfrak{m}}=\sum_{j=1}^{d}(-1)^{[X_{j}]}[X_{j},\widebar{X}_{j}]_{\mathfrak{m}}=0.$ (3.15)
Using the $\operatorname{ad}_{\mathfrak{k}}$-invariance of $\langle-,$-$\rangle$, it also follows that, for every $Y\in\mathfrak{k}_{\bar{0}}$,
$\sum_{j=1}^{d}\,\langle[Y,\widebar{X}_{j}]_{\mathfrak{m}},X_{j}\rangle=0.$ (3.16)
We are now in a position to state and prove our main curvature formulas: Proposition 3.12, Theorem 3.13 and Corollary 3.14. The proof of the expression for the Ricci curvature given in Theorem 3.13 is lengthy and deferred until Section 3.4. We emphasise that the formulas below only give the evaluations of the corresponding curvature tensors at the base point $K$. However, by [San10, Theorem 4.16] (cf. the discussion following Theorem 3.6), this data determines the full curvature tensors.
Proposition 3.12.
Let $M=G/K$ be a homogeneous superspace with $G$-invariant graded Riemannian metric $g$, where $G=(G_{0},\mathfrak{g})$ is a connected Lie supergroup, $\mathfrak{g}$ a compact real form of a basic classical Lie superalgebra, and $K=(K_{0},\mathfrak{k})$ a connected closed Lie subsupergroup of $G$. Let $\mathfrak{m}$ be a $Q$-orthogonal complement of $\mathfrak{k}$, where $Q$ is an $\operatorname{ad}_{\mathfrak{g}^{\mathbb{C}}}$-invariant non-degenerate supersymmetric even bilinear form on $\mathfrak{g}^{\mathbb{C}}$, and let $X_{i},X_{j},X_{k},X_{\ell}\in\mathfrak{m}$. Then, at the base point $K$,
$4\,\langle(R$ $(X_{i},X_{j})X_{k},X_{\ell}\rangle=-2\,\langle[X_{i},X_{j}]_{\mathfrak{m}},[X_{k},X_{\ell}]_{\mathfrak{m}}\rangle+(-1)^{[X_{i}]([X_{j}]+[X_{k}])}\langle[X_{j},X_{k}]_{\mathfrak{m}},[X_{i},X_{\ell}]_{\mathfrak{m}}\rangle$
$-(-1)^{[X_{j}][X_{k}]}\langle[X_{i},X_{k}]_{\mathfrak{m}},[X_{j},X_{\ell}]_{\mathfrak{m}}\rangle-\langle X_{i},[[X_{j},X_{k}]_{\mathfrak{m}},X_{\ell}]_{\mathfrak{m}}\rangle+\langle[X_{i},[X_{j},X_{k}]_{\mathfrak{m}}]_{\mathfrak{m}},X_{\ell}\rangle$
$+(-1)^{[X_{i}][X_{j}]}\langle X_{j},[[X_{i},X_{k}]_{\mathfrak{m}},X_{\ell}]_{\mathfrak{m}}\rangle+(-1)^{[X_{j}][X_{k}]}\langle[[X_{i},X_{k}]_{\mathfrak{m}},X_{j}]_{\mathfrak{m}},X_{\ell}\rangle$
$+(-1)^{[X_{i}]([X_{j}]+[X_{k}])}\langle X_{j},[X_{k},[X_{i},X_{\ell}]_{\mathfrak{m}}]_{\mathfrak{m}}\rangle+(-1)^{[X_{i}][X_{j}]+[X_{k}][X_{\ell}]}\langle[X_{j},[X_{i},X_{\ell}]_{\mathfrak{m}}]_{\mathfrak{m}},X_{k}\rangle$
$-(-1)^{[X_{j}][X_{k}]}\langle X_{i},[X_{k},[X_{j},X_{\ell}]_{\mathfrak{m}}]_{\mathfrak{m}}\rangle-(-1)^{[X_{k}][X_{\ell}]}\langle[X_{i},[X_{j},X_{\ell}]_{\mathfrak{m}}]_{\mathfrak{m}},X_{k}\rangle$
$+2\,\langle X_{i},[X_{j},[X_{k},X_{\ell}]_{\mathfrak{g}}]_{\mathfrak{m}}\rangle-2(-1)^{[X_{i}][X_{j}]}\langle X_{j},[X_{i},[X_{k},X_{\ell}]_{\mathfrak{g}}]_{\mathfrak{m}}\rangle$
$+(-1)^{[X_{i}]([X_{j}]+[X_{k}])}\langle U(X_{j},X_{k}),U(X_{i},X_{\ell})\rangle-(-1)^{[X_{j}][X_{k}]}\langle U(X_{i},X_{k}),U(X_{j},X_{\ell})\rangle.$
Proof.
With the shorthand notation $\nabla_{i}=\nabla_{X_{i}}$ and $i]=[X_{i}]$ for each $i$, and applying (3.1) and (3.4) to the defining expression for the Riemannian curvature, we get
$\langle R(X_{i},X_{j})X_{k},X_{\ell}\rangle$ $=\langle\nabla_{[X_{i},X_{j}]}X_{k},X_{\ell}\rangle-\langle\nabla_{i}\nabla_{j}X_{k},X_{\ell}\rangle+(-1)^{[i][j]}\langle\nabla_{j}\nabla_{i}X_{k},X_{\ell}\rangle$
$=-(-1)^{[k][\ell]}\langle[X_{i},X_{j}],\nabla_{\ell}X_{k}\rangle-X_{i}\,\langle\nabla_{j}X_{k},X_{\ell}\rangle+(-1)^{[i]([j]+[k])}\langle\nabla_{j}X_{k},\nabla_{i}X_{\ell}\rangle$
$\quad+(-1)^{[i][j]}X_{j}\,\langle\nabla_{i}X_{k},X_{\ell}\rangle-(-1)^{[j][k]}\langle\nabla_{i}X_{k},\nabla_{j}X_{\ell}\rangle.$
Applying (3.2) and Lemma 3.3 to this expression, we obtain
$\langle R(X_{i},X_{j})X_{k},X_{\ell}\rangle=$ $-(-1)^{[k][\ell]}\langle\nabla_{i}X_{j},\nabla_{\ell}X_{k}\rangle+(-1)^{[i][j]+[k][\ell]}\langle\nabla_{j}X_{i},\nabla_{\ell}X_{k}\rangle$
$+(-1)^{[i]([j]+[k])}\langle\nabla_{j}X_{k},\nabla_{i}X_{\ell}\rangle-(-1)^{[j][k]}\langle\nabla_{i}X_{k},\nabla_{j}X_{\ell}\rangle$
$-\tfrac{1}{2}X_{i}\,\langle X_{j},[X_{k},X_{\ell}]\rangle-\tfrac{1}{2}(-1)^{[j][k]}X_{i}\,\langle X_{k},[X_{j},X_{\ell}]\rangle$
$-\tfrac{1}{2}(-1)^{[\ell]([j]+[k])}X_{i}\,\langle X_{\ell},[X_{j},X_{k}]\rangle+\tfrac{1}{2}(-1)^{[i][j]}X_{j}\,\langle X_{i},[X_{k},X_{\ell}]\rangle$
$+\tfrac{1}{2}(-1)^{[i]([j]+[k])}X_{j}\,\langle X_{k},[X_{i},X_{\ell}]\rangle+\tfrac{1}{2}(-1)^{[\ell]([i]+[k])+[i][j]}X_{j}\,\langle X_{\ell},[X_{i},X_{k}]\rangle.$
Applying Proposition 3.9 to each of the first four terms, and (3.3) to each of the last six terms, yields
$4\,\langle R(X_{i},X_{j})X_{k},X_{\ell}\rangle$ $=2\,\langle[X_{i},X_{j}]_{\mathfrak{m}},U(X_{k},X_{\ell})\rangle-(-1)^{[i]([j]+[k])}\langle[X_{j},X_{k}]_{\mathfrak{m}},U(X_{i},X_{\ell})\rangle$
$\quad-(-1)^{[i]([j]+[k])}\langle U(X_{j},X_{k}),[X_{i},X_{\ell}]_{\mathfrak{m}}\rangle+(-1)^{[j][k]}\langle[X_{i},X_{k}]_{\mathfrak{m}},U(X_{j},X_{\ell})\rangle$
$\quad+(-1)^{[j][k]}\langle U(X_{i},X_{k}),[X_{j},X_{\ell}]_{\mathfrak{m}}\rangle-2\,\langle[X_{i},X_{j}]_{\mathfrak{m}},[X_{k},X_{\ell}]_{\mathfrak{m}}\rangle$
$\quad+(-1)^{[i]([j]+[k])}\langle[X_{j},X_{k}]_{\mathfrak{m}},[X_{i},X_{\ell}]_{\mathfrak{m}}\rangle+(-1)^{[i]([j]+[k])}\langle U(X_{j},X_{k}),U(X_{i},X_{\ell})\rangle$
$\quad-(-1)^{[j][k]}\langle[X_{i},X_{k}]_{\mathfrak{m}},[X_{j},X_{\ell}]_{\mathfrak{m}}\rangle-(-1)^{[j][k]}\langle U(X_{i},X_{k}),U(X_{j},X_{\ell})\rangle$
$\quad-2(-1)^{([i]+[j])[k]}\langle X_{k},[[X_{i},X_{j}]_{\mathfrak{g}},X_{\ell}]_{\mathfrak{m}}\rangle-2(-1)^{[\ell]}\langle X_{\ell},[[X_{i},X_{j}]_{\mathfrak{g}},X_{k}]_{\mathfrak{m}}\rangle$
$\quad-2(-1)^{[i][j]}\langle X_{j},[X_{i},[X_{k},X_{\ell}]_{\mathfrak{g}}]_{\mathfrak{m}}\rangle+2\,\langle X_{i},[X_{j},[X_{k},X_{\ell}]_{\mathfrak{g}}]_{\mathfrak{m}}\rangle.$
The desired formula now follows by applying (3.10) to the first five terms, and using
$(-1)^{[X_{k}][X_{\ell}]}\langle[[X_{i},X_{j}]_{\mathfrak{k}},X_{\ell}]_{\mathfrak{m}},X_{k}\rangle$ $=-\langle[[X_{i},X_{j}]_{\mathfrak{k}},X_{k}]_{\mathfrak{m}},X_{\ell}\rangle,$
which follows from the $\operatorname{ad}_{\mathfrak{k}}$-invariance. ∎
Theorem 3.13.
Let $M=G/K$ be a homogeneous superspace with $G$-invariant graded Riemannian metric $g$, where $G=(G_{0},\mathfrak{g})$ is a connected Lie supergroup, $\mathfrak{g}$ a compact real form of a basic classical Lie superalgebra, and $K$ a connected closed Lie subsupergroup of $G$. Let $\{X_{1},\ldots,X_{d}\}$ be a $g$-normalised homogeneous basis for $\mathfrak{m}=T_{K}(M)$, with right dual basis $\{\widebar{X}_{1},\ldots,\widebar{X}_{d}\}$. Then, at the base point $K$, for every $i\in\{1,\ldots,d\}$,
$\operatorname{Ric}(X_{i},\widebar{X}_{i})=-\frac{1}{2}B(X_{i},\widebar{X}_{i})+\frac{1}{2}\sum_{j=1}^{d}\,\langle[X_{i},\widebar{X}_{j}]_{\mathfrak{m}},[X_{j},\widebar{X}_{i}]_{\mathfrak{m}}\rangle-\frac{1}{4}\sum_{j,k=1}^{d}\,\langle X_{i},[\widebar{X}_{j},\widebar{X}_{k}]_{\mathfrak{m}}\rangle\,\langle[X_{k},X_{j}]_{\mathfrak{m}},\widebar{X}_{i}\rangle,$
where $B$ denotes the Killing form on $\mathfrak{g}$.
Corollary 3.14.
Let the notation be as in Theorem 3.13. Then, at the base point $K$, the scalar curvature of $g$ satisfies
$S(K)=-\frac{1}{2}\sum_{i=1}^{d}B(\widebar{X}_{i},X_{i})+\frac{1}{4}\sum_{i,j=1}^{d}\langle[\widebar{X}_{i},\widebar{X}_{j}]_{\mathfrak{m}},[X_{j},X_{i}]_{\mathfrak{m}}\rangle.$
Proof.
First apply the Ricci curvature formula from Theorem 3.13 to (3.7). Use Lemma 3.10 and apply $\sum_{i=1}^{d}\langle[\widebar{X}_{j},\widebar{X}_{k}]_{\mathfrak{m}},\widebar{X}_{i}\rangle\,X_{i}=[\widebar{X}_{j},\widebar{X}_{k}]_{\mathfrak{m}}$ to the ensuing triple sum. Change summation variables to obtain the desired formula. ∎
3.4 Proof of the Ricci curvature formula
This section is devoted to proving the Ricci curvature formula given in Theorem 3.13. Although our general strategy is similar to that in [Bes87, Chapter 7], the super-setting complicates matters considerably. Throughout, $\{X_{1},\ldots,X_{d}\}$ denotes a $g$-normalised homogeneous basis for $\mathfrak{m}=T_{K}(M)$, with right dual basis $\{\widebar{X}_{1},\ldots,\widebar{X}_{d}\}$, where $\widebar{X}_{i}]=[X_{i}]$ for all $i$. As in the proof of Proposition 3.12, we will use the shorthand notation $\nabla_{i}=\nabla_{X_{i}}$ and $i]=[X_{i}]$. Since the extended superproduct $\langle-,-\rangle:\mathfrak{g}\times\mathfrak{g}\to\mathbb{R}$ is an even bilinear map, an expression like $\langle Y_{1},[Y_{2},Y_{3}]\rangle$, with $Y_{1},Y_{2},Y_{3}\in\mathfrak{g}$, is zero if $Y_{1}]\neq[Y_{2}]+[Y_{3}]$. We first establish a sequence of technical results: Lemma 3.15 to Lemma 3.19.
To set the stage, note that
$\begin{gathered}U(\widebar{X}_{i},X_{j})=\sum_{k=1}^{d}\,\langle U(\widebar{X}_{i},X_{j}),\widebar{X}_{k}\rangle\,X_{k}=-\sum_{k=1}^{d}\Big(\langle\widebar{X_{i}},[X_{j},\widebar{X}_{k}]_{\mathfrak{m}}\rangle+(-1)^{[i][j]}\langle X_{j},[\widebar{X}_{i},\widebar{X}_{k}]_{\mathfrak{m}}\rangle\Big)X_{k},\\ U(X_{i},\widebar{X}_{j})=\sum_{k=1}^{d}\,\langle X_{k},U(X_{i},\widebar{X}_{j})\rangle\,\widebar{X}_{k}=\sum_{k=1}^{d}\Big(\langle[X_{k},X_{i}]_{\mathfrak{m}},\widebar{X}_{j}]\rangle+(-1)^{[i][j]}\langle[X_{k},\widebar{X}_{j}]_{\mathfrak{m}},X_{i}\rangle\Big)\widebar{X}_{k},\end{gathered}$ (3.19)
where the second equality in each line follows from (3.10).
For clarity, we also note that
$\operatorname{Str}_{\mathfrak{m}}(\operatorname{ad}_{\mathfrak{m}}(X))=\sum_{i=1}^{d}\,\langle\widebar{X}_{i},[X,X_{i}]_{\mathfrak{m}}\rangle=\sum_{i=1}^{d}(-1)^{[i]}\langle X_{i},[X,\widebar{X}_{i}]_{\mathfrak{m}}\rangle,\qquad X\in\mathfrak{g},$ (3.20)
with the second equality following from Lemma 3.10, and that this expression is zero if $X\in\mathfrak{g}_{\bar{1}}$.
Lemma 3.15.
It holds that
$\sum_{i=1}^{d}U(\widebar{X}_{i},X_{i})=\sum_{i=1}^{d}(-1)^{[i]}U(X_{i},\widebar{X}_{i})=0.$
Proof.
The first equality follows from Lemma 3.10. For the second equality, let
$Z:=\sum_{i=1}^{d}U(\widebar{X}_{i},X_{i}),$
so $Z\in\mathfrak{m}_{\bar{0}}$. We want to show that $\langle Z,X\rangle=0$ for all $X\in\mathfrak{g}$. Since $\langle\mathfrak{k},\mathfrak{m}\rangle=0$, it suffices to consider $X\in\mathfrak{m}$, and since $Z\in\mathfrak{m}_{\bar{0}}$, the relation trivially holds if $X\in\mathfrak{m}_{\bar{1}}$, so we will assume that $X\in\mathfrak{m}_{\bar{0}}$. It then follows from (3.10) that
$\langle Z,X\rangle=\sum_{i=1}^{d}\,\langle\widebar{X}_{i},[X,X_{i}]_{\mathfrak{m}}\rangle+\sum_{i=1}^{d}(-1)^{[i]}\langle X_{i},[X,\widebar{X}_{i}]_{\mathfrak{m}}\rangle=2\operatorname{Str}_{\mathfrak{m}}(\operatorname{ad}_{\mathfrak{m}}(X))=2\operatorname{Str}_{\mathfrak{g}}(\operatorname{ad}_{\mathfrak{g}}(X)),$
where the last equality uses that $X\in\mathfrak{m}$, $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{m}$ and $\mathfrak{k},\mathfrak{m}]\subseteq\mathfrak{m}$. Since $\mathfrak{g}$ is simple, we have $\mathfrak{g}=[\mathfrak{g},\mathfrak{g}]$, and since $\operatorname{Str}([f_{1},f_{2}])=0$ for all $f_{1},f_{2}\in\mathrm{End}(\mathfrak{g})$, it follows that $\operatorname{Str}_{\mathfrak{g}}(\operatorname{ad}_{\mathfrak{g}}(X))=0$. ∎
Corollary 3.16.
For each $i=1,\ldots,d$, we have
$\sum_{j=1}^{d}\,\langle U(X_{i},\widebar{X}_{i}),U(\widebar{X}_{j},X_{j})\rangle=0.$
Remark 3.17.
The lemma holds as long as $\operatorname{Str}_{\mathfrak{g}}(\operatorname{ad}_{\mathfrak{g}}(X))=0$ for all $X$, even if $\mathfrak{g}$ is not simple. For example, the lemma holds for $\mathfrak{g}=\mathfrak{sl}(n|n)$.
Lemma 3.18.
For each $i=1,\ldots,d$, we have
$\sum_{j=1}^{d}(-1)^{[i]+[j]}\langle U(X_{j},\widebar{X}_{i}),U(X_{i},\widebar{X}_{j})\rangle=-\sum_{j,k=1}^{d}\,\langle X_{i},[\widebar{X}_{k},\widebar{X}_{j}]_{\mathfrak{m}}\rangle\,\langle[X_{j},X_{k}]_{\mathfrak{m}},\widebar{X}_{i}\rangle$
$+\sum_{k=1}^{d}\Big(\langle X_{i},[[\widebar{X}_{i},\widebar{X}_{k}]_{\mathfrak{m}},X_{k}]_{\mathfrak{m}}\rangle+\langle[[X_{i},\widebar{X}_{k}]_{\mathfrak{m}},X_{k}]_{\mathfrak{m}},\widebar{X}_{i}\rangle-\langle[X_{i},\widebar{X}_{k}]_{\mathfrak{m}},[X_{k},\widebar{X}_{i}]_{\mathfrak{m}}\rangle\Big).$
Proof.
Use (3.19) to rewrite the lefthand side as a sum of four double sums. Then, apply (3.5) to $X=[\widebar{X}_{i},\widebar{X}_{k}]_{\mathfrak{m}}$, respectively $X=[X_{i},X_{k}]_{\mathfrak{m}}$ (twice), to reduce three of the double sums to single sums, and apply Lemma 3.10. ∎
Lemma 3.19.
Let $X,Y\in\mathfrak{m}$ be homogeneous. Then,
$B(X,Y)=\sum_{i=1}^{d}\Big(\langle\widebar{X}_{i},[X,[Y,X_{i}]_{\mathfrak{g}}]_{\mathfrak{m}}\rangle+(-1)^{[X]}\langle\widebar{X}_{i},[Y,[X,X_{i}]_{\mathfrak{k}}]_{\mathfrak{m}}\rangle\Big).$
Proof.
Let $\{Y_{1},\dots,Y_{h}\}$ be a homogeneous basis for $\mathfrak{k}$, with right dual basis $\{\widebar{Y}_{1},\dots,\widebar{Y}_{h}\}$, and recall that $\langle\mathfrak{k},\mathfrak{m}\rangle=0$. Expressing $\operatorname{ad}(X)\operatorname{ad}(Y)(\widebar{X}_{i})$ and $\operatorname{ad}(X)\operatorname{ad}(Y)(\widebar{Y}_{k})$ in terms of the dual basis, and using a formula similar to (3.20) but for a supertrace over all of $\mathfrak{g}$, yields
$B(X,Y)=\sum_{i=1}^{d}(-1)^{[X_{i}]}\langle X_{i},[X,[Y,\widebar{X}_{i}]_{\mathfrak{g}}]_{\mathfrak{m}}\rangle+\sum_{k=1}^{h}(-1)^{[Y_{k}]}\langle Y_{k},[X,[Y,\widebar{Y_{k}}]_{\mathfrak{g}}]_{\mathfrak{k}}\rangle.$
Use (3.9) and $X,[Y,\widebar{Y_{k}}]_{\mathfrak{k}}]_{\mathfrak{g}}\in\mathfrak{m}$ to rewrite the sum over $k$ as
$\sum_{k=1}^{h}(-1)^{[Y_{k}]}\langle Y_{k},[X,[Y,\widebar{Y_{k}}]_{\mathfrak{m}}]_{\mathfrak{k}}\rangle$ $=-\sum_{i=1}^{d}\sum_{k=1}^{h}(-1)^{[Y_{k}]+[Y][Y_{k}]}\langle X_{i},[\widebar{Y}_{k},Y]_{\mathfrak{m}}\rangle\,\langle Y_{k},[X,\widebar{X}_{i}]_{\mathfrak{k}}\rangle$
$=\sum_{i=1}^{d}(-1)^{[X_{i}]+[X]}\langle X_{i},[Y,[X,\widebar{X}_{i}]_{\mathfrak{k}}]_{\mathfrak{m}}\rangle,$
where the first equality follows from $Y,\widebar{Y_{k}}]_{\mathfrak{m}}=\sum_{i=1}^{d}\,\langle X_{i},[Y,\widebar{Y_{k}}]_{\mathfrak{m}}\rangle\widebar{X}_{i}$, and the second from $X,\widebar{X}_{i}]_{\mathfrak{k}}=\sum_{k=1}^{h}\,\langle Y_{k},[X,\widebar{X}_{i}]_{\mathfrak{k}}\rangle\,\widebar{Y}_{k}$. By Lemma 3.10, this establishes the desired expression. ∎
Using Lemma 3.10, it follows from Lemma 3.19 that, for each $i=1,\ldots,d$,
$B(X_{i},\widebar{X}_{i})=\sum_{j=1}^{d}\Big(\langle[X_{i},[\widebar{X}_{i},\widebar{X}_{j}]_{\mathfrak{g}}]_{\mathfrak{m}},X_{j}\rangle+\langle[[\widebar{X}_{j},X_{i}]_{\mathfrak{k}},\widebar{X}_{i}]_{\mathfrak{m}},X_{j}\rangle\Big).$ (3.21)
Proof of Theorem 3.13.
Since the $g$-normalised basis and its right-dual companion are related as in (3.14), we can use (3.6) and Proposition 3.12 to obtain an expression for $\operatorname{Ric}(X_{i},\widebar{X}_{i})$. Applying Lemma 3.16 and (3.15) to the expression for $\mathrm{Ric}(X_{i},\widebar{X}_{i})$, the corresponding terms vanish, yielding
$4\mathrm{Ric}(X_{i},\widebar{X_{i}})$ $=-2\sum_{j=1}^{d}(-1)^{[j]+[i][j]}\langle[X_{i},X_{j}]_{\mathfrak{m}},[\widebar{X}_{i},\widebar{X}_{j}]_{\mathfrak{m}}\rangle+\sum_{j=1}^{d}(-1)^{[i]+[j]}\langle[X_{j},\widebar{X}_{i}]_{\mathfrak{m}},[X_{i},\widebar{X}_{j}]_{\mathfrak{m}}\rangle$
$-\sum_{j=1}^{d}(-1)^{[j]+[i][j]}\langle X_{i},[[X_{j},\widebar{X}_{i}]_{\mathfrak{m}},\widebar{X}_{j}]_{\mathfrak{m}}\rangle+\sum_{j=1}^{d}(-1)^{[j]+[i][j]}\langle[X_{i},[X_{j},\widebar{X}_{i}]_{\mathfrak{m}}]_{\mathfrak{m}},\widebar{X}_{j}\rangle$
$+\sum_{j=1}^{d}(-1)^{[j]}\langle X_{j},[[X_{i},\widebar{X}_{i}]_{\mathfrak{m}},\widebar{X_{j}}]_{\mathfrak{m}}\rangle+\sum_{j=1}^{d}(-1)^{[j]}\langle[[X_{i},\widebar{X}_{i}]_{\mathfrak{m}},X_{j}]_{\mathfrak{m}},\widebar{X}_{j}\rangle$
$+\sum_{j=1}^{d}(-1)^{[i]+[j]}\langle X_{j},[\widebar{X}_{i},[X_{i},\widebar{X}_{j}]_{\mathfrak{m}}]_{\mathfrak{m}}\rangle+\sum_{j=1}^{d}(-1)^{[j]+[i][j]}\langle[X_{j},[X_{i},\widebar{X}_{j}]_{\mathfrak{m}}]_{\mathfrak{m}},\widebar{X}_{i}\rangle$
$+2\sum_{j=1}^{d}(-1)^{[j]+[i][j]}\langle X_{i},[X_{j},[\widebar{X}_{i},\widebar{X}_{j}]_{\mathfrak{g}}]_{\mathfrak{m}}\rangle-2\sum_{j=1}^{d}(-1)^{[j]}\langle X_{j},[X_{i},[\widebar{X}_{i},\widebar{X}_{j}]_{\mathfrak{g}}]_{\mathfrak{m}}\rangle$
$+\sum_{j=1}^{d}(-1)^{[i]+[j]}\langle U(X_{j},\widebar{X}_{i}),U(X_{i},\widebar{X}_{j})\rangle.$
Signs may be eliminated by invoking Lemma 3.10 and by applying the skew-supersymmetry of the Lie bracket together with the supersymmetry of the scalar product. Applying Lemma 3.18 and (3.9) to the ensuing expression for $\operatorname{Ric}(X_{i},\widebar{X}_{i})$, and using (3.21), yields
$4\operatorname{Ric}(X_{i},\widebar{X_{i}})+B(X_{i},\widebar{X}_{i})+(-1)^{[i]}B(\widebar{X}_{i},X_{i})$
$=2\sum_{j=1}^{d}\,\langle[X_{i},\widebar{X}_{j}]_{\mathfrak{m}},[X_{j},\widebar{X}_{i}]_{\mathfrak{m}}\rangle-\sum_{j,k=1}^{d}\,\langle X_{i},[\widebar{X}_{j},\widebar{X}_{k}]_{\mathfrak{m}}\rangle\,\langle[X_{k},X_{j}]_{\mathfrak{m}},\widebar{X}_{i}\rangle$
$+2\sum_{j=1}^{d}\Big(\langle[[\widebar{X}_{j},X_{i}]_{\mathfrak{g}},\widebar{X}_{i}]_{\mathfrak{m}},X_{j}\rangle+\langle[[X_{i},\widebar{X}_{i}]_{\mathfrak{m}},\widebar{X}_{j}]_{\mathfrak{m}},X_{j}\rangle$
$\qquad\qquad-\langle[X_{i},[\widebar{X}_{i},\widebar{X}_{j}]_{\mathfrak{m}}]_{\mathfrak{m}},X_{j}\rangle-\langle X_{i},[[\widebar{X}_{i},\widebar{X}_{j}]_{\mathfrak{k}},X_{j}]_{\mathfrak{m}}\rangle\Big).$ (3.22)
We can use the super-Jacobi identity
$[[\widebar{X}_{j},X_{i}]_{\mathfrak{g}},\widebar{X}_{i}]_{\mathfrak{m}}=-[[X_{i},\widebar{X}_{i}]_{\mathfrak{g}},\widebar{X}_{j}]_{\mathfrak{m}}+[X_{i},[\widebar{X}_{i},\widebar{X}_{j}]_{\mathfrak{g}}]_{\mathfrak{m}}$
to rewrite the first term in the last $j$-sum, and use (3.16) to conclude that
$\sum_{j=1}^{d}\,\langle[[X_{i},\widebar{X}_{i}]_{\mathfrak{g}},\widebar{X}_{j}]_{\mathfrak{m}},X_{j}\rangle=\sum_{j=1}^{d}\,\langle[[X_{i},\widebar{X}_{i}]_{\mathfrak{m}},\widebar{X}_{j}]_{\mathfrak{m}},X_{j}\rangle,$
while $\operatorname{ad}_{\mathfrak{k}}$-invariance implies that
$\langle X_{i},[[\widebar{X}_{i},\widebar{X}_{j}]_{\mathfrak{k}},X_{j}]_{\mathfrak{m}}\rangle=\langle[X_{i},[\widebar{X}_{i},\widebar{X}_{j}]_{\mathfrak{k}}]_{\mathfrak{m}},X_{j}\rangle.$
It follows that the last $j$-sum vanishes, and as $B(X_{i},\widebar{X}_{i})=(-1)^{[i]}B(\widebar{X}_{i},X_{i})$, the formula in Theorem 3.13 follows. ∎
3.5 Naturally reductive metrics
We say that the $G$-invariant metric $g$ on the homogeneous superspace $M=G/K$ is naturally reductive if
$\langle[X,Y]_{\mathfrak{m}},Z\rangle=\langle X,[Y,Z]_{\mathfrak{m}}\rangle$ (3.23)
for all $X,Y,Z\in\mathfrak{m}$. This mimics the definition of a naturally reductive metric in classical homogeneous geometry [DZ79]. Such metrics have been used extensively in the study of the Einstein equation, spectral geometry, the prescribed Ricci curvature problem, and other topics; see, e.g., [DZ79, GS10, Lau20, AGP20]. When $g$ is naturally reductive, the curvature formulas obtained above simplify substantially.
Lemma 3.20.
Let $g$ be naturally reductive. Then, $U=0$.
Proof.
Applying (3.23) to (3.10) yields
$\langle U(X,Y),Z\rangle=0,\qquad X,Y,Z\in\mathfrak{m}.$
Since this holds for all $Z$, the result follows from the non-degeneracy of $\langle-,$-$\rangle$. ∎
Proposition 3.21.
Let $M=G/K$ be a homogeneous superspace with $G$-invariant graded naturally reductive Riemannian metric $g$, where $G=(G_{0},\mathfrak{g})$ is a connected Lie supergroup, $\mathfrak{g}$ a compact real form of a basic classical Lie superalgebra, and $K$ a connected closed Lie subsupergroup of $G$. Let $\{X_{1},\ldots,X_{d}\}$ be a $g$-normalised homogeneous basis for $\mathfrak{m}=T_{K}(M)$, with right dual basis $\{\widebar{X}_{1},\ldots,\widebar{X}_{d}\}$. Then, for every $i,j,k,\ell\in\{1,\ldots,d\}$,
$4\,\langle R(X_{i},X_{j})X_{k},X_{\ell}\rangle$ $=2\,\langle X_{i},[X_{j},[X_{k},X_{\ell}]_{\mathfrak{m}}]_{\mathfrak{m}}\rangle+4\,\langle X_{i},[X_{j},[X_{k},X_{\ell}]_{\mathfrak{k}}]_{\mathfrak{m}}\rangle$
$\quad+(-1)^{[X_{j}][X_{k}]}\langle X_{i},[X_{k},[X_{j},X_{\ell}]_{\mathfrak{m}}]_{\mathfrak{m}}\rangle-(-1)^{([X_{j}]+[X_{k}])[X_{\ell}]}\langle X_{i},[X_{\ell},[X_{j},X_{k}]_{\mathfrak{m}}]_{\mathfrak{m}}\rangle$
and
$\operatorname{Ric}(X_{i},\widebar{X}_{i})$ $=-\frac{1}{2}B(X_{i},\widebar{X}_{i})+\frac{1}{4}\sum_{j=1}^{d}\,\langle[X_{i},\widebar{X}_{j}]_{\mathfrak{m}},[X_{j},\widebar{X}_{i}]_{\mathfrak{m}}\rangle.$
Proof.
By applying (3.23) and Lemma 3.20 to the curvature expressions in Proposition 3.12 and Theorem 3.13, the result follows. ∎
3.6 Diagonal metrics
Here, we discuss particular specialisations of the formula for the Ricci curvature that we will use to obtain our main results on Einstein metrics in Section 4.
Recall that $\mathfrak{g}$ admits the $Q$-orthogonal decomposition $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{m}$. As $K$ is connected, this means that $\mathfrak{m}$ is an $\operatorname{ad}_{\mathfrak{k}}$-representation, that is, $\mathfrak{k},\mathfrak{m}]\subseteq\mathfrak{m}$. In the following, we will assume $\mathfrak{m}$ decomposes into $Q$-orthogonal $\operatorname{ad}_{\mathfrak{k}}$-irreducible representations,
$\mathfrak{m}=\mathfrak{m}_{1}\oplus\cdots\oplus\mathfrak{m}_{s},$ (3.24)
where $\mathfrak{m}_{i}\not\cong\mathfrak{m}_{j}$ as $\operatorname{ad}_{\mathfrak{k}}$-representations, for all $i\neq j$. Such a decomposition is unique up to permutation of summands. As $\mathfrak{m}_{i}$ is $\operatorname{ad}_{\mathfrak{k}}$-irreducible, and both $Q$ and the Killing form $B$ are even bilinear forms, Schur’s lemma for Lie superalgebras [Kac77] tells us that there exist $b_{1},\ldots,b_{s}\in\mathbb{R}$ such that
$B|_{\mathfrak{m}_{i}}=-b_{i}Q|_{\mathfrak{m}_{i}},\qquad i=1,\ldots,s,$ (3.25)
where the sign convention is adopted from the non-super setting.
As we will discuss in the following, the curvature formulas obtained in Section 3.3 simplify quite dramatically for diagonal metrics. We will frequently use an ordered basis for $\mathfrak{m}$ that is adapted to the decomposition (3.24). For each $i=1,\ldots,s$, we thus let $\mathfrak{M}_{i}=\{e^{i}_{\alpha}\,|\,\alpha=1,\ldots,\dim(\mathfrak{m}_{i})\}$ be an ordered $Q$-normalised homogeneous basis for $\mathfrak{m}_{i}$ such that the corresponding matrix representation of $Q$ is of the form (3.11), and let $\{\bar{e}_{\alpha}^{i}\,|\,\alpha=1,\ldots,\dim(\mathfrak{m}_{i})\}$ be the right $Q$-dual basis, defined by
$Q(e_{\alpha}^{i},\bar{e}_{\alpha^{\prime}}^{i})=\delta_{\alpha,\alpha^{\prime}},\qquad\alpha,\alpha^{\prime}\in\{1,\ldots,\dim(\mathfrak{m}_{i})\}.$
Ordered as indicated, $\mathfrak{M}_{1}\cup\cdots\cup\mathfrak{M}_{s}$ then constitutes a suitably ordered basis for $\mathfrak{m}$. For each $i=1,\ldots,s$, we let $I^{i}$ denote the index set of the basis $\mathfrak{M}_{i}$ for $\mathfrak{m}_{i}$.
Let $g$ be a $G$-invariant graded Riemannian metric on $M$. Then, by Theorem 3.6, $g$ is identified with an $\operatorname{Ad}_{K}$-invariant scalar superproduct $\langle-,$-$\rangle$ on $\mathfrak{m}$. Since $K$ is connected, the $\operatorname{Ad}_{K}$-invariance is equivalent to the $\operatorname{ad}_{\mathfrak{k}}$-invariance of $\langle-,$-$\rangle$ at the level of Lie superalgebras. By Schur’s lemma, the metric $g$ is therefore of the form
$\langle-,-\rangle=\sum_{i=1}^{s}x_{i}Q|_{\mathfrak{m}_{i}},\qquad x_{1},\ldots,x_{s}\in\mathbb{R}^{\times},$ (3.26)
and we say that it is diagonal with respect to the decomposition (3.24). Unlike in classical Riemannian geometry, the parameters $x_{i}$ need not be positive. If they all are, then we say that the metric is positive. Moreover, the Ricci curvature is an even $\operatorname{ad}_{\mathfrak{k}}$-invariant supersymmetric bilinear form on $\mathfrak{m}$, so it is of the form
$\operatorname{Ric}=\sum_{i=1}^{s}r_{i}Q|_{\mathfrak{m}_{i}},\qquad r_{1},\ldots,r_{s}\in\mathbb{R}.$ (3.27)
With this, the Einstein equation (1.1) becomes
$r_{i}=cx_{i},\qquad i=1,\ldots,s.$ (3.28)
3.6.1 Structure constants
We now introduce super-analogues of the structure constants; see, e.g., [WZ86].
Definition 3.22.
For each triple $i,j,k\in\{1,\ldots,s\}$, let
$[ijk]:=-\sum_{\alpha\in I^{i},\,\beta\in I^{j},\,\gamma\in I^{k}}Q\big(e_{\gamma}^{k},[e_{\alpha}^{i},e_{\beta}^{j}]_{\mathfrak{m}_{k}}\big)\,Q\big([\bar{e}_{\beta}^{j},\bar{e}_{\alpha}^{i}]_{\mathfrak{m}_{k}},\bar{e}_{\gamma}^{k}\big).$
The $\operatorname{ad}_{\mathfrak{g}}$-invariance of $Q$ implies that the structure constant $[ijk]$ is symmetric in all three indices.
Remark 3.23.
If $\dim(\mathfrak{m}_{\bar{1}})=0$ and $Q|_{\mathfrak{m}_{\bar{0}}}$ is positive definite, i.e., $M$ is a classical homogeneous Riemannian manifold, then $\bar{e}_{\alpha}^{i}=e_{\alpha}^{i}$ and the structure constant $[ijk]$ coincides with its classical counterpart.
Let $\operatorname{Str}_{\mathfrak{m}_{i}}(f)$ denote the supertrace of the restriction $f|_{\mathfrak{m}_{i}}$ for any $f\in\mathrm{End}(\mathfrak{g})$, and let $\operatorname{ad}_{\mathfrak{m}_{i}}(X):=[X,-]_{\mathfrak{m}_{i}}$ for any $X\in\mathfrak{g}$. We then have the following result.
Lemma 3.24.
For each triple $i,j,k\in\{1,\ldots,s\}$, we have
$[ijk]=-\sum_{\alpha\in I^{i}}(-1)^{[e_{\alpha}^{i}]}\operatorname{Str}_{\mathfrak{m}_{j}}\!\big(\!\operatorname{ad}_{\mathfrak{m}_{j}}(e_{\alpha}^{i})\operatorname{ad}_{\mathfrak{m}_{k}}(\bar{e}_{\alpha}^{i})\big).$
Moreover, $[ijk]$ is independent of the $Q$-normalised homogeneous bases of $\mathfrak{m}_{1},\ldots,\mathfrak{m}_{s}$.
Proof.
Using $\bar{e}^{j}_{\beta},\bar{e}_{\alpha}^{i}]_{\mathfrak{m}_{k}}=\sum_{\gamma\in I^{k}}Q([\bar{e}^{j}_{\beta},\bar{e}_{\alpha}^{i}]_{\mathfrak{m}},\bar{e}_{\gamma}^{k})e_{\gamma}^{k}$, we see that
$[ijk]=-\sum_{\alpha\in I^{i},\,\beta\in I^{j}}Q\big([\bar{e}^{j}_{\beta},\bar{e}_{\alpha}^{i}]_{\mathfrak{m}_{k}},[e_{\alpha}^{i},e_{\beta}^{j}]_{\mathfrak{m}_{k}}\big),$
so the desired expression for $[ijk]$ follows from
$\sum_{\beta\in I^{j}}Q\big([\bar{e}^{j}_{\beta},\bar{e}_{\alpha}^{i}]_{\mathfrak{m}_{k}},[e_{\alpha}^{i},e_{\beta}^{j}]_{\mathfrak{m}_{k}}\big)$ $=\sum_{\beta\in I^{j}}Q\big([[\bar{e}^{j}_{\beta},\bar{e}_{\alpha}^{i}]_{\mathfrak{m}_{k}},e_{\alpha}^{i}]_{\mathfrak{m}_{j}},e_{\beta}^{j}\big)=\sum_{\beta\in I^{j}}(-1)^{[e_{\alpha}^{i}]+[e_{\beta}^{j}]}Q\big(e_{\beta}^{j},[e_{\alpha}^{i},[\bar{e}_{\alpha}^{i},\bar{e}_{\beta}^{j}]_{\mathfrak{m}_{k}}]_{\mathfrak{m}_{j}}\big)$
$=(-1)^{[e_{\alpha}^{i}]}\operatorname{Str}_{\mathfrak{m}_{j}}\!\big(\!\operatorname{ad}_{\mathfrak{m}_{j}}(e_{\alpha}^{i})\operatorname{ad}_{\mathfrak{m}_{k}}(\bar{e}_{\alpha}^{i})\big).$ (3.29)
It remains to show that $[ijk]$ is independent of the choice of basis for $\mathfrak{m}_{i}$. Suppose there is another $Q$-normalised homogeneous basis $\{f_{\beta}^{i}\,|\,\beta\in I^{i}\}$ for $\mathfrak{m}_{i}$, with right dual basis $\{\bar{f}_{\beta}^{i}\,|\,\beta\in I^{i}\}$, and let $A^{i}$ and $B^{i}$ be the (even) transition matrices such that
$f_{\beta}^{i}=\sum_{\alpha\in I^{i}}A^{i}_{\beta\alpha}e_{\alpha}^{i},\qquad\bar{f}_{\beta}^{i}=\sum_{\alpha\in I^{i}}B^{i}_{\beta\alpha}\bar{e}_{\alpha}^{i}.$
Then, $B^{i}=((A^{i})^{-1})^{T}$ and
$\sum_{\beta\in I^{i}}(-1)^{[f_{\beta}^{i}]}\operatorname{Str}_{\mathfrak{m}_{j}}\!\big(\!\operatorname{ad}_{\mathfrak{m}_{j}}(f_{\beta}^{i})\operatorname{ad}_{\mathfrak{m}_{k}}(\bar{f}_{\beta}^{i})\big)$ $=\sum_{\alpha,\beta,\gamma\in I^{i}}(-1)^{[e_{\alpha}^{i}]}A^{i}_{\beta\alpha}B_{\beta\gamma}^{i}\operatorname{Str}_{\mathfrak{m}_{j}}\!\big(\!\operatorname{ad}_{\mathfrak{m}_{j}}(e_{\alpha}^{i})\operatorname{ad}_{\mathfrak{m}_{k}}(\bar{e}_{\gamma}^{i})\big)$
$=\sum_{\alpha\in I^{i}}(-1)^{[e_{\alpha}^{i}]}\operatorname{Str}_{\mathfrak{m}_{j}}\!\big(\!\operatorname{ad}_{\mathfrak{m}_{j}}(e_{\alpha}^{i})\operatorname{ad}_{\mathfrak{m}_{k}}(\bar{e}_{\alpha}^{i})\big).$
This proves the basis independence of $[ijk]$. ∎
Let $\{z_{1},\ldots,z_{\dim(\mathfrak{k})}\}$ be a homogeneous basis for $\mathfrak{k}$, with right $Q$-dual basis $\{\bar{z}_{1},\ldots,\bar{z}_{\dim(\mathfrak{k})}\}$. For each $i=1,\ldots,s$, define the Casimir operator on $\mathfrak{m}_{i}$ by
$C_{\mathfrak{m}_{i},Q|_{\mathfrak{k}}}:=-\sum_{\ell=1}^{\dim(\mathfrak{k})}(-1)^{[z_{\ell}]}\operatorname{ad}(z_{\ell})\operatorname{ad}(\bar{z}_{\ell}),$
where the negative sign is incorporated as a convention (cf. [WZ86]). Since $\mathfrak{m}_{i}$ is $\operatorname{ad}_{\mathfrak{k}}$-irreducible and $C_{\mathfrak{m}_{i},Q|_{\mathfrak{k}}}$ is an even operator, Schur’s lemma implies that there exists $c_{i}\in\mathbb{R}$ such that
$C_{\mathfrak{m}_{i},Q|_{\mathfrak{k}}}=c_{i}\mathrm{Id},$ (3.30)
where $\mathrm{Id}$ denotes the identity operator. Using this, the next result provides a super-analogue of [WZ86, Lemma 1.5]. For ease of notation, we set
$d_{i}:=\mathrm{sdim}(\mathfrak{m}_{i})=\dim((\mathfrak{m}_{i})_{\bar{0}})-\dim((\mathfrak{m}_{i})_{\bar{1}}),\qquad i=1,\ldots,s.$ (3.31)
Note that $d_{i}$, as well as $b_{i}$ and $[ijk]$, can be negative in the super-setting.
Proposition 3.25.
Let $M=G/K$ be a homogeneous superspace, with notation as in Theorem 3.13, and suppose $\mathfrak{m}$ admits a $Q$-orthogonal multiplicity-free decomposition of the form (3.24). Then,
$\sum_{j,k=1}^{s}[ijk]=d_{i}(b_{i}-2c_{i}),\qquad i=1,\dots,s,$
where $d_{i}$, $b_{i}$ and $c_{i}$ are given in (3.31), (3.25) and (3.30).
Proof.
Using Lemma 3.24, we have
$\sum_{j,k=1}^{s}[ijk]=-\sum_{j,k=1}^{s}\sum_{\alpha\in I^{i}}(-1)^{[e_{\alpha}^{i}]}\operatorname{Str}_{\mathfrak{m}_{j}}\!\big(\!\operatorname{ad}_{\mathfrak{m}_{j}}(e_{\alpha}^{i})\operatorname{ad}_{\mathfrak{m}_{k}}(\bar{e}_{\alpha}^{i})\big)=-\sum_{\alpha\in I^{i}}(-1)^{[e_{\alpha}^{i}]}\operatorname{Str}_{\mathfrak{m}}\!\big(\!\operatorname{ad}_{\mathfrak{m}}(e_{\alpha}^{i})\operatorname{ad}_{\mathfrak{m}}(\bar{e}_{\alpha}^{i})\big).$
Since $\mathfrak{g}$ admits the $Q$-orthogonal decomposition $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{m}$ and $\mathfrak{k},\mathfrak{m}]\subseteq\mathfrak{m}$, we have
$\operatorname{Str}_{\mathfrak{m}}\!\big(\!\operatorname{ad}_{\mathfrak{m}}(e_{\alpha}^{i})\operatorname{ad}_{\mathfrak{m}}(\bar{e}_{\alpha}^{i})\big)=\operatorname{Str}_{\mathfrak{g}}\!\big(\!\operatorname{ad}_{\mathfrak{g}}(e_{\alpha}^{i})\operatorname{ad}_{\mathfrak{g}}(\bar{e}_{\alpha}^{i})\big)-\operatorname{Str}_{\mathfrak{m}}\!\big(\!\operatorname{ad}_{\mathfrak{m}}(e_{\alpha}^{i})\operatorname{ad}_{\mathfrak{k}}(\bar{e}_{\alpha}^{i})\big)-\operatorname{Str}_{\mathfrak{k}}\!\big(\!\operatorname{ad}_{\mathfrak{k}}(e_{\alpha}^{i})\operatorname{ad}_{\mathfrak{m}}(\bar{e}_{\alpha}^{i})\big).$
Since $\operatorname{Str}_{V}(fg)=(-1)^{[f][g]}\operatorname{Str}_{W}(gf)$ for any two homogeneous linear maps $f:W\to V$ and $g:V\to W$ between vector superspaces $V$ and $W$, we see that $\operatorname{Str}_{\mathfrak{m}}(\operatorname{ad}_{\mathfrak{m}}(e_{\alpha}^{i})\operatorname{ad}_{\mathfrak{k}}(\bar{e}_{\alpha}^{i}))=(-1)^{[e_{\alpha}^{i}]}\operatorname{Str}_{\mathfrak{k}}(\operatorname{ad}_{\mathfrak{k}}(\bar{e}_{\alpha}^{i})\operatorname{ad}_{\mathfrak{m}}(e_{\alpha}^{i}))$. It follows that
$\sum_{j,k=1}^{s}[ijk]$ $=-\sum_{\alpha\in I^{i}}(-1)^{[e_{\alpha}^{i}]}B(e_{\alpha}^{i},\bar{e}_{\alpha}^{i})+\sum_{\alpha\in I^{i}}\operatorname{Str}_{\mathfrak{k}}\!\big(\!\operatorname{ad}_{\mathfrak{k}}(\bar{e}_{\alpha}^{i})\operatorname{ad}_{\mathfrak{m}}(e_{\alpha}^{i})\big)+\sum_{\alpha\in I^{i}}(-1)^{[e_{\alpha}^{i}]}\operatorname{Str}_{\mathfrak{k}}\!\big(\!\operatorname{ad}_{\mathfrak{k}}(e_{\alpha}^{i})\operatorname{ad}_{\mathfrak{m}}(\bar{e}_{\alpha}^{i})\big)$
$=\sum_{\alpha\in I^{i}}(-1)^{[e_{\alpha}^{i}]}b_{i}Q(e_{\alpha}^{i},\bar{e}_{\alpha}^{i})+2\sum_{\alpha\in I^{i}}(-1)^{[e_{\alpha}^{i}]}\operatorname{Str}_{\mathfrak{k}}\!\big(\!\operatorname{ad}_{\mathfrak{k}}(e_{\alpha}^{i})\operatorname{ad}_{\mathfrak{m}}(\bar{e}_{\alpha}^{i})\big)$
$=b_{i}d_{i}+2\sum_{\alpha\in I^{i}}(-1)^{[e_{\alpha}^{i}]}\sum_{\ell=1}^{\dim(\mathfrak{k})}(-1)^{[z_{\ell}]}Q\big(z_{\ell},[e_{\alpha}^{i},[\bar{e}_{\alpha}^{i},\bar{z}_{\ell}]_{\mathfrak{m}}]_{\mathfrak{k}}\big)$
$=b_{i}d_{i}+2\sum_{\alpha\in I^{i}}(-1)^{[e_{\alpha}^{i}]}\sum_{\ell=1}^{\dim(\mathfrak{k})}(-1)^{[z_{\ell}]}Q\big(e_{\alpha}^{i},[z_{\ell},[\bar{z}_{\ell},\bar{e}_{\alpha}^{i}]_{\mathfrak{m}}]_{\mathfrak{m}}\big)$
$=b_{i}d_{i}+2\sum_{\alpha\in I^{i}}(-1)^{[e_{\alpha}^{i}]}Q\big(e_{\alpha}^{i},-C_{\mathfrak{m}_{i},Q|_{\mathfrak{k}}}(\bar{e}_{\alpha}^{i})\big)$
$=b_{i}d_{i}-2d_{i}c_{i},$
where the fourth equality uses the $\operatorname{ad}_{\mathfrak{g}}$-invariance and supersymmetry of $Q$. ∎
3.6.2 Ricci curvature
For diagonal metrics, we now derive expressions for the Ricci coefficients $r_{i}$, $i=1,\ldots,s$, in (3.27). Although the superdimensions $d_{i}$ may be 0 (unlike the corresponding ordinary dimensions), the formula for the Ricci coefficients in Theorem 3.26 below is analogous to the corresponding classical result; cf. [PS97, Lemma 1.1] and [PR19, Lemma 3.3]. If $\dim(\mathfrak{m}_{\bar{1}})=0$, then Theorem 3.26 gives a Ricci curvature formula for a pseudo-Riemannian manifold.
Theorem 3.26.
Let $M=G/K$ be a homogeneous superspace, with notation as in Theorem 3.13, and suppose $\mathfrak{m}$ admits a $Q$-orthogonal multiplicity-free decomposition of the form (3.24). With the notation for $g$ as in (3.26), each of the corresponding Ricci coefficients $r_{i}$, $i=1,\ldots,s$, satisfies
$d_{i}r_{i}=\frac{b_{i}d_{i}}{2}+\sum_{j,k=1}^{s}\frac{[ijk]}{4}\Big(\frac{x_{i}^{2}}{x_{j}x_{k}}-\frac{2x_{j}}{x_{k}}\Big).$ (3.32)
Proof.
For each $i=1,\ldots,s$, we introduce $\epsilon_{i}=\pm 1$ such that $\tilde{x}_{i}=\epsilon_{i}x_{i}>0$, and write $\tilde{r}_{i}=\epsilon_{i}r_{i}$. Establishing the stated formula is then equivalent to showing that
$4d_{i}\tilde{r}_{i}=2b_{i}d_{i}\epsilon_{i}+\sum_{j,k=1}^{s}[ijk]\epsilon_{i}\epsilon_{j}\epsilon_{k}\Big(\frac{\tilde{x}_{i}^{2}}{\tilde{x}_{j}\tilde{x}_{k}}-\frac{2\tilde{x}_{j}}{\tilde{x}_{k}}\Big),\qquad i=1,\ldots,s.$
For each $i=1,\ldots,s$, let $\{X^{i}_{\alpha}\,|\,\alpha\in I^{i}\}$ be a $g$-normalised basis for $\mathfrak{m}_{i}$, with right $g$-dual basis $\{\widebar{X}_{\alpha}^{i}\,|\,\alpha\in I^{i}\}$, and let $\{X_{1},\dots,X_{d}\}$ be a corresponding basis for $\mathfrak{m}$ adapted to the decomposition (3.24). Then, $Q$-orthogonality implies that $\langle X^{i}_{\alpha},X^{j}_{\beta}\rangle=0$ for $i\neq j$ and any $\alpha,\beta$. For each $i=1,\ldots,s$, we also introduce the renormalised basis $\{e^{i}_{\alpha}=\sqrt{\tilde{x}_{i}}X_{\alpha}^{i}\,|\,\alpha\in I^{i}\}$ for $\mathfrak{m}_{i}$, whose right $Q$-dual basis is seen to be given by $\{\bar{e}_{\alpha}^{i}=\epsilon_{i}\sqrt{\tilde{x}_{i}}\widebar{X}^{i}_{\alpha}\,|\,\alpha\in I^{i}\}$.
We now fix $i\in\{1,\ldots,s\}$ and $\alpha\in I^{i}$, and set $X=e^{i}_{\alpha}$ and $\widebar{X}=\epsilon_{i}\bar{e}_{\alpha}^{i}$. Then,
$Q(X,\widebar{X})=\epsilon_{i},\qquad B(X,\widebar{X})=-b_{i}\epsilon_{i},\qquad\operatorname{Ric}(X,\widebar{X})=\tilde{r}_{i}.$
By Theorem 3.13, we have
$4\tilde{r}_{i}$ $=2b_{i}\epsilon_{i}-2\sum_{j=1}^{d}(-1)^{[X_{\alpha}^{i}][X_{j}]+[X_{j}]}\,\tilde{x}_{i}\,\langle[X_{\alpha}^{i},X_{j}]_{\mathfrak{m}},[\widebar{X}_{\alpha}^{i},\widebar{X}_{j}]_{\mathfrak{m}}\rangle$
$\quad+\sum_{j,k=1}^{d}(-1)^{[X_{k}][X_{j}]+[X_{\alpha}^{i}]}\,\tilde{x}_{i}\,\langle X_{\alpha}^{i},[X_{k},X_{j}]_{\mathfrak{m}}\rangle\,\langle[\widebar{X}_{k},\widebar{X}_{j}]_{\mathfrak{m}},\widebar{X}_{\alpha}^{i}\rangle$
$=2b_{i}\epsilon_{i}-2\sum_{j,k=1}^{d}(-1)^{[X_{\alpha}^{i}][X_{j}]+[X_{\alpha}^{i}]}\,\tilde{x}_{i}\,\langle X_{k},[X_{\alpha}^{i},X_{j}]_{\mathfrak{m}}\rangle\,\langle[\widebar{X}_{\alpha}^{i},\widebar{X}_{j}]_{\mathfrak{m}},\widebar{X}_{k}\rangle$
$\quad+\sum_{j,k=1}^{d}(-1)^{[X_{k}][X_{j}]+[X_{\alpha}^{i}]}\,\tilde{x}_{i}\,\langle X_{\alpha}^{i},[X_{k},X_{j}]_{\mathfrak{m}}\rangle\,\langle[\widebar{X}_{k},\widebar{X}_{j}]_{\mathfrak{m}},\widebar{X}_{\alpha}^{i}\rangle,$
$=2b_{i}\epsilon_{i}-2\sum_{j,k=1}^{s}\sum_{\beta\in I^{j},\,\gamma\in I^{k}}(-1)^{[X_{\alpha}^{i}][X_{\beta}^{j}]+[X_{\alpha}^{i}]}\,\tilde{x}_{i}\,\langle X_{\gamma}^{k},[X_{\alpha}^{i},X_{\beta}^{j}]_{\mathfrak{m}}\rangle\,\langle[\widebar{X}_{\alpha}^{i},\widebar{X}_{\beta}^{j}]_{\mathfrak{m}},\widebar{X}_{\gamma}^{k}\rangle$
$\quad+\sum_{j,k=1}^{s}\sum_{\beta\in I^{j},\,\gamma\in I^{k}}(-1)^{[X_{\gamma}^{k}][X_{\beta}^{j}]+[X_{\alpha}^{i}]}\,\tilde{x}_{i}\,\langle X_{\alpha}^{i},[X_{\gamma}^{k},X_{\beta}^{j}]_{\mathfrak{m}}\rangle\,\langle[\widebar{X}_{\gamma}^{k},\widebar{X}_{\beta}^{j}]_{\mathfrak{m}},\widebar{X}_{\alpha}^{i}\rangle$
$=2b_{i}\epsilon_{i}+2\sum_{j,k=1}^{s}\sum_{\beta\in I^{j},\,\gamma\in I^{k}}(-1)^{[e_{\alpha}^{i}]}\,\frac{\tilde{x}_{k}}{\tilde{x}_{j}}\epsilon_{i}\epsilon_{j}\epsilon_{k}Q\big(e_{\gamma}^{k},[e_{\alpha}^{i},e_{\beta}^{j}]_{\mathfrak{m}_{k}}\big)\,Q\big([\bar{e}_{\beta}^{j},\bar{e}^{i}_{\alpha}]_{\mathfrak{m}_{k}},\bar{e}_{\gamma}^{k}\big)$
$\quad-\sum_{j,k=1}^{s}\sum_{\beta\in I^{j},\,\gamma\in I^{k}}(-1)^{[e_{\alpha}^{i}]}\,\frac{\tilde{x}_{i}^{2}}{\tilde{x}_{j}\tilde{x}_{k}}\epsilon_{i}\epsilon_{j}\epsilon_{k}Q\big(e_{\alpha}^{i},[e_{\gamma}^{k},e_{\beta}^{j}]_{\mathfrak{m}_{i}}\big)\,Q\big([\bar{e}_{\beta}^{j},\bar{e}_{\gamma}^{k}]_{\mathfrak{m}_{i}},\bar{e}^{i}_{\alpha}\big)$
$=2b_{i}\epsilon_{i}-\sum_{j,k=1}^{s}\sum_{\beta\in I^{j},\,\gamma\in I^{k}}(-1)^{[e_{\alpha}^{i}]}\epsilon_{i}\epsilon_{j}\epsilon_{k}Q\big(e_{\gamma}^{k},[e_{\alpha}^{i},e_{\beta}^{j}]_{\mathfrak{m}_{k}}\big)\,Q\big([\bar{e}_{\beta}^{j},\bar{e}^{i}_{\alpha}]_{\mathfrak{m}_{k}},\bar{e}_{\gamma}^{k}\big)\Big(\frac{\tilde{x}_{i}^{2}}{\tilde{x}_{j}\tilde{x}_{k}}-\frac{2\tilde{x}_{k}}{\tilde{x}_{j}}\Big)$
$=2b_{i}\epsilon_{i}-\sum_{j,k=1}^{s}\sum_{\beta\in I^{j}}(-1)^{[e_{\alpha}^{i}]}\epsilon_{i}Q\big([\bar{e}_{\beta}^{j},\bar{e}^{i}_{\alpha}]_{\mathfrak{m}_{k}},[e_{\alpha}^{i},e_{\beta}^{j}]_{\mathfrak{m}_{k}}\big)\Big(\frac{x_{i}^{2}}{x_{j}x_{k}}-\frac{2x_{k}}{x_{j}}\Big)$
$=2b_{i}\epsilon_{i}-\sum_{j,k=1}^{s}\operatorname{Str}_{\mathfrak{m}_{j}}\!\big(\!\operatorname{ad}_{\mathfrak{m}_{j}}(e_{\alpha}^{i})\operatorname{ad}_{\mathfrak{m}_{k}}(\bar{e}_{\alpha}^{i})\big)\epsilon_{i}\Big(\frac{x_{i}^{2}}{x_{j}x_{k}}-\frac{2x_{k}}{x_{j}}\Big),$ (3.33)
where the second, second-to-last and last equalities follow from $X_{\alpha}^{i},X_{j}]_{\mathfrak{m}}=\sum_{k=1}^{d}\,\langle X_{k},[X_{\alpha}^{i},X_{j}]_{\mathfrak{m}}\rangle\,\widebar{X}_{k}$, $\bar{e}_{\beta}^{j},\bar{e}^{i}_{\alpha}]_{\mathfrak{m}_{k}}=\sum_{\gamma\in I^{k}}Q\big([\bar{e}_{\beta}^{j},\bar{e}^{i}_{\alpha}]_{\mathfrak{m}_{k}},\bar{e}_{\gamma}^{k}\big)e_{\gamma}^{k}$ and (3.6.1), respectively. Taking the graded sum over the $Q$-normalised homogeneous basis for $\mathfrak{m}_{i}$, we obtain
$4d_{i}r_{i}=4\sum_{\alpha\in I^{i}}(-1)^{[e_{\alpha}^{i}]}\,r_{i}=2b_{i}d_{i}-\sum_{\alpha\in I^{i}}\sum_{j,k=1}^{s}(-1)^{[e_{\alpha}^{i}]}\operatorname{Str}_{\mathfrak{m}_{j}}\!\big(\!\operatorname{ad}_{\mathfrak{m}_{j}}(e_{\alpha}^{i})\operatorname{ad}_{\mathfrak{m}_{k}}(\bar{e}_{\alpha}^{i})\big)\Big(\frac{x_{i}^{2}}{x_{j}x_{k}}-\frac{2x_{k}}{x_{j}}\Big).$
The desired formula now follows by application of Lemma 3.24 and the symmetry of $[ijk]$. ∎
Proposition 3.27.
With the notation as in Theorem 3.26, the scalar curvature $S$ is given by
$S=\frac{1}{2}\sum_{i=1}^{s}\frac{b_{i}d_{i}}{x_{i}}-\frac{1}{4}\sum_{i,j,k=1}^{s}[ijk]\frac{x_{k}}{x_{i}x_{j}}.$
Proof.
It follows from (3.27) that
$S=\sum_{i=1}^{d}(-1)^{[X_{i}]}\operatorname{Ric}(X_{i},\widebar{X}_{i})=\sum_{i=1}^{s}\sum_{\alpha\in I^{i}}(-1)^{[X_{\alpha}^{i}]}r_{i}Q(X_{\alpha}^{i},\widebar{X}_{\alpha}^{i})=\sum_{i=1}^{s}\frac{d_{i}r_{i}}{x_{i}}.$
Using Theorem 3.26 and the symmetry of $[ijk]$, we obtain the formula for $S$. ∎
If $d_{i}\neq 0$, then Theorem 3.26 says that
$r_{i}=\frac{b_{i}}{2}+\sum_{j,k=1}^{s}\frac{[ijk]}{4d_{i}}\Big(\frac{x_{i}^{2}}{x_{j}x_{k}}-\frac{2x_{j}}{x_{k}}\Big).$ (3.34)
If $d_{i}=0$, on the other hand, we cannot isolate $r_{i}$ from (3.32) but will now derive a formula for $r_{i}$ in case there exists an $i$-selected pair $(j,k)$. Here, $(j,k)$ is $i$-selected if $j,k\in\{1,\ldots,s\}$ are such that $j\leq k$ and
$[\mathfrak{m}_{i},\mathfrak{m}_{u}]_{\mathfrak{m}_{v}}={0}$ (3.35)
for all $1\leq u\leq v\leq s$ such that $(u,v)\neq(j,k)$. In Section 4.4 and Section 4.5, we give concrete examples where (3.35) is satisfied; see Proposition 4.4 and Proposition 4.17.
Proposition 3.28.
Let the notation be as in Theorem 3.26, and suppose $d_{i}=0$ for some $i\in\{1,\ldots,s\}$, and that there exists $i$-selected $(j,k)$. Then,
$r_{i}=\frac{b_{i}}{2}+\frac{b_{i}-2c_{i}}{4}\Big(\frac{x_{i}^{2}}{x_{j}x_{k}}-\frac{x_{j}}{x_{k}}-\frac{x_{k}}{x_{j}}\Big).$
Proof.
In the homogeneous $\mathfrak{m}_{i}$-basis $\{e^{i}_{\alpha}\,|\,\alpha\in I^{i}\}$ introduced in the proof of Theorem 3.26, there exists at least one even vector, because $d_{i}=0$. Let us select one such vector and denote it by $e_{\alpha}^{i}$. Since (3.35) holds, we have
$\operatorname{Str}_{\mathfrak{m}}\!\big(\!\operatorname{ad}_{\mathfrak{m}}(e_{\alpha}^{i})\operatorname{ad}_{\mathfrak{m}}(\bar{e}_{\alpha}^{i})\big)=\sum_{u,v=1}^{s}\operatorname{Str}_{\mathfrak{m}_{u}}\!\big(\!\operatorname{ad}_{\mathfrak{m}_{u}}(e_{\alpha}^{i})\operatorname{ad}_{\mathfrak{m}_{v}}(\bar{e}_{\alpha}^{i})\big)=(2-\delta_{jk})\operatorname{Str}_{\mathfrak{m}_{j}}\!\big(\!\operatorname{ad}_{\mathfrak{m}_{j}}(e_{\alpha}^{i})\operatorname{ad}_{\mathfrak{m}_{k}}(\bar{e}_{\alpha}^{i})\big).$
We also have
$\operatorname{Str}_{\mathfrak{m}}\!\big(\!\operatorname{ad}_{\mathfrak{m}}(e_{\alpha}^{i})\operatorname{ad}_{\mathfrak{m}}(\bar{e}_{\alpha}^{i})\big)$ $=\operatorname{Str}_{\mathfrak{g}}\!\big(\!\operatorname{ad}_{\mathfrak{g}}(e_{\alpha}^{i})\operatorname{ad}_{\mathfrak{g}}(\bar{e}_{\alpha}^{i})\big)-2\operatorname{Str}_{\mathfrak{k}}\!\big(\!\operatorname{ad}_{\mathfrak{k}}(e_{\alpha}^{i})\operatorname{ad}_{\mathfrak{m}}(\bar{e}_{\alpha}^{i})\big)$
$=B(e_{\alpha}^{i},\bar{e}_{\alpha}^{i})-2\sum_{\ell=1}^{\dim(\mathfrak{k})}(-1)^{[z_{\ell}]}Q\big(z_{\ell},[e_{\alpha}^{i},[\bar{e}_{\alpha}^{i},\bar{z}_{\ell}]_{\mathfrak{m}}]_{\mathfrak{k}}\big)$
$=-b_{i}Q(e_{\alpha}^{i},\bar{e}_{\alpha}^{i})-2Q\big(e_{\alpha}^{i},-C_{\mathfrak{m}_{i},Q|_{\mathfrak{k}}}(\bar{e}_{\alpha}^{i})\big)$
$=-b_{i}+2c_{i}.$
We now use (3.6.2) and obtain
$r_{i}$ $=\frac{b_{i}}{2}-\frac{1}{4}\sum_{u,v=1}^{s}\operatorname{Str}_{\mathfrak{m}_{u}}\!\big(\!\operatorname{ad}_{\mathfrak{m}_{u}}(e_{\alpha}^{i})\operatorname{ad}_{\mathfrak{m}_{v}}(\bar{e}_{\alpha}^{i})\big)\Big(\frac{x_{i}^{2}}{x_{u}x_{v}}-\frac{2x_{v}}{x_{u}}\Big)$
$=\frac{b_{i}}{2}-\frac{2-\delta_{jk}}{4}\operatorname{Str}_{\mathfrak{m}_{j}}\!\big(\!\operatorname{ad}_{\mathfrak{m}_{j}}(e_{\alpha}^{i})\operatorname{ad}_{\mathfrak{m}_{k}}(\bar{e}_{\alpha}^{i})\big)\Big(\frac{x_{i}^{2}}{x_{j}x_{k}}-\frac{x_{j}}{x_{k}}-\frac{x_{k}}{x_{j}}\Big),$
from which the expression for $r_{i}$ follows. ∎
4 Flag supermanifolds
In this section, we introduce a class of homogeneous superspaces called flag supermanifolds. In analogy with the construction of generalised flag manifolds in terms of so-called painted Dynkin diagrams in the non-super case [Arv03], we introduce what we call circled Dynkin diagrams to describe the flag supermanifolds in Section 4.1. In Section 4.2, we identify the compact real forms of the Lie superalgebras of types $A$ and $C$. In Section 4.3, Section 4.4 and Section 4.5, we then classify Einstein metrics on flag supermanifolds constructed from circled Dynkin diagrams of types $A$ and $C$. We refer to [Kac77] for Dynkin diagrams of complex Lie superalgebras, to [Ser83, Par90, Chu13] for classifications of real forms of Lie superalgebras, and to [Kac78] for the representation theory of complex Lie superalgebras.
4.1 Circled Dynkin diagrams
Let $\mathfrak{g}^{\mathbb{C}}=\mathfrak{g}_{\bar{0}}^{\mathbb{C}}\oplus\mathfrak{g}_{\bar{1}}^{\mathbb{C}}$ be a basic classical Lie superalgebra, with $Q$ a non-degenerate supersymmetric even bilinear form. Let $\mathfrak{h}^{\mathbb{C}}\subset\mathfrak{g}_{\bar{0}}^{\mathbb{C}}$ be a Cartan subalgebra of $\mathfrak{g}^{\mathbb{C}}$, and set $(\mathfrak{h}^{\mathbb{C}})^{\vee}:=\mathrm{Hom}_{\mathbb{C}}(\mathfrak{h}^{\mathbb{C}},\mathbb{C})$. For each $\alpha\in(\mathfrak{h}^{\mathbb{C}})^{\vee}$, define
$\mathfrak{g}_{\alpha}^{\mathbb{C}}:=\big\{X\in\mathfrak{g}^{\mathbb{C}}\,|\,[H,X]=\alpha(H)X,\ \text{for all}\ H\in\mathfrak{h}^{\mathbb{C}}\big\}.$
If $\mathfrak{g}_{\alpha}^{\mathbb{C}}\neq\{0\}$ for $\alpha\neq 0$, we call $\alpha$ a root and $\mathfrak{g}^{\mathbb{C}}_{\alpha}$ the corresponding root space. In a basic Lie superalgebra, every root $\alpha$ is either even (if $\mathfrak{g}^{\mathbb{C}}_{\alpha}\subseteq\mathfrak{g}_{\bar{0}}^{\mathbb{C}}$) or odd (if $\mathfrak{g}^{\mathbb{C}}_{\alpha}\subseteq\mathfrak{g}_{\bar{1}}^{\mathbb{C}}$), and $\dim(\mathfrak{g}^{\mathbb{C}}_{\alpha})=1$. The set of all roots, $\Delta$, called the root system, thus partitions as $\Delta=\Delta_{\bar{0}}\cup\Delta_{\bar{1}}$, where $\Delta_{\bar{0}}$ is the set of even roots (in fact, the root system of $\mathfrak{g}_{\bar{0}}^{\mathbb{C}}$) and $\Delta_{\bar{1}}$ the set of odd roots. As in the non-super case, we have a root space decomposition:
$\mathfrak{g}^{\mathbb{C}}=\mathfrak{h}^{\mathbb{C}}\oplus\bigoplus_{\alpha\in\Delta}\mathfrak{g}_{\alpha}^{\mathbb{C}}.$
We choose a distinguished Borel subsuperalgebra $\mathfrak{b}^{\mathbb{C}}$ such that the corresponding set $\Pi=\{\alpha_{1},\dots,\alpha_{\ell}\}$ of simple roots contains a single odd simple root. The Weyl vector $\rho=\rho_{\bar{0}}-\rho_{\bar{1}}$ of $\mathfrak{g}^{\mathbb{C}}$ is defined as half the graded sum of the positive roots, that is,
$\rho_{\bar{0}}=\tfrac{1}{2}\sum_{\alpha\in\Delta^{+}_{\bar{0}}}\alpha,\qquad\rho_{\bar{1}}=\tfrac{1}{2}\sum_{\alpha\in\Delta^{+}_{\bar{1}}}\alpha,$ (4.1)
where $\Delta^{+}=\Delta^{+}_{\bar{0}}\cup\Delta^{+}_{\bar{1}}$ denotes the set of positive roots. The bilinear form $Q|_{\mathfrak{h}^{\mathbb{C}}}$ is non-degenerate, so we may identify $\mathfrak{h}^{\mathbb{C}}$ with $(\mathfrak{h}^{\mathbb{C}})^{\vee}$. A root $\alpha$ is called isotropic if $Q(\alpha,\alpha)=0$. Note that isotropic roots are odd.
We now choose a subset $\Pi_{K}$ of $\Pi$ and denote by $\Delta_{K}\subseteq\Delta$ the system generated from $\Pi_{K}$. Then,
$\mathfrak{k}^{\mathbb{C}}:=\mathfrak{h}^{\mathbb{C}}\oplus\bigoplus_{\alpha\in\Delta_{K}}\mathfrak{g}_{\alpha}^{\mathbb{C}}$
is a Lie subsuperalgebra of $\mathfrak{g}^{\mathbb{C}}$. Note that $\mathfrak{k}^{\mathbb{C}}$ decomposes into its centre $Z(\mathfrak{k}^{\mathbb{C}})$ and the semisimple part $\mathfrak{k}_{\mathrm{ss}}^{\mathbb{C}}:=[\mathfrak{k}^{\mathbb{C}},\mathfrak{k}^{\mathbb{C}}]$. Let $\mathfrak{g}$ be the compact real form of $\mathfrak{g}^{\mathbb{C}}$ induced from a star operation $\ast$ on $\mathfrak{g}^{\mathbb{C}}$. Then, we can form the Lie subsupergroup $K=(K_{0},\mathfrak{k})$ of $G$, where $\mathfrak{k}=\mathfrak{k}^{\mathbb{C}}\cap\mathfrak{g}$ is a Lie subsuperalgebra of $\mathfrak{g}$, and $K_{0}$ is the Lie subgroup of $G_{0}$ generated by $\mathfrak{k}_{\bar{0}}$. In this way, we obtain a homogeneous superspace $M=G/K$ called a flag supermanifold.
The tangent space $T_{K}(M)$ of $M$ at $K$ can be described as follows. Let $\Pi_{M}=\Pi-\Pi_{K}$ and $\Delta_{M}=\Delta-\Delta_{K}$. We have the $Q$-orthogonal vector-space decomposition $\mathfrak{g}^{\mathbb{C}}=\mathfrak{k}^{\mathbb{C}}\oplus\mathfrak{m}^{\mathbb{C}}$, and we identify the complexified tangent space $T_{K}(M)^{\mathbb{C}}$ with
$\mathfrak{m}^{\mathbb{C}}=\bigoplus_{\alpha\in\Delta_{M}}\mathfrak{g}_{\alpha}^{\mathbb{C}}.$ (4.2)
Similarly to (2.4), we then have
$\mathfrak{m}=\operatorname{span}_{\mathbb{R}}\big\{X\in\mathfrak{m}_{\bar{0}}^{\mathbb{C}}\,|\,X^{*}=-X\big\}\oplus\operatorname{span}_{\mathbb{R}}\big\{\sqrt{\imath}X\in\mathfrak{m}_{\bar{1}}^{\mathbb{C}}\,|\,X^{*}=-X\big\},$ (4.3)
where $\ast$ is the star operation on $\mathfrak{g}^{\mathbb{C}}$. For each $\alpha\in\Delta^{+}_{M}$, let $E_{\pm\alpha}$ span $\mathfrak{g}^{\mathbb{C}}_{\pm\alpha}$ and define
$A_{\alpha}:=E_{\alpha}-E_{-\alpha},\qquad B_{\alpha}:=\imath(E_{\alpha}+E_{-\alpha}),$ (4.4)
noting that $(A_{\alpha})^{*}=-A_{\alpha}$ and $(B_{\alpha})^{*}=-B_{\alpha}$ under the star operation (4.5). It follows that $\mathfrak{m}$ has $\mathbb{R}$-basis
$\{A_{\alpha},B_{\alpha}\,|\,\alpha\in\Delta_{M}\cap\Delta_{\bar{0}}^{+}\}\cup\{\sqrt{\imath}A_{\alpha},\sqrt{\imath}B_{\alpha}\,|\,\alpha\in\Delta_{M}\cap\Delta_{\bar{1}}^{+}\}.$
In the non-super case, the above construction is encoded in a so-called painted Dynkin diagram, where the nodes of $\Pi_{M}$ are painted black while the other nodes of $\Pi$ remain white. In the super-setting, however, black nodes typically represent non-isotropic odd simple roots, so to avoid confusion, we shall circle the nodes of $\Pi_{M}$, thereby obtaining a circled Dynkin diagram of $G$. In addition, a node marked with a cross, $\otimes$, denotes an isotropic odd simple root, in accordance with the standard Dynkin diagram conventions for basic classical Lie superalgebras. The uncircled nodes of $\Pi$ form a simple root system of $\mathfrak{k}_{\mathrm{ss}}^{\mathbb{C}}$.
Conversely, given a Dynkin diagram of $G$ with some nodes circled, we can construct the corresponding homogeneous superspace $M=G/K$ as follows. Let $\mathfrak{g}^{\mathbb{C}}$ be the complex simple Lie superalgebra characterised by the (uncircled) Dynkin diagram. The Lie superalgebra $\mathfrak{g}$ in $G=(G_{0},\mathfrak{g})$ is then the compact real form of $\mathfrak{g}^{\mathbb{C}}$ induced from a specified star operation, whereas the Lie group $G_{0}$ is the unique simply connected Lie group generated by $\mathfrak{g}_{\bar{0}}$. The semisimple part $K_{\mathrm{ss}}$ of $K$ is obtained from the Dynkin diagram by removing the nodes that are circled, while $K=K_{ss}\times\operatorname{U}(1)^{n}$, where $n$ is the number of circled nodes of the Dynkin diagram. This determines the homogeneous superspace $M=G/K$.
4.2 Compact real forms
To construct a flag supermanifold $M=(G_{0},\mathfrak{g})/(K_{0},\mathfrak{k})$ from a given circled Dynkin diagram, we first specify the compact real form $\mathfrak{g}$. In the following, we recall Dynkin diagrams, star operations and compact real forms for the complex Lie superalgebras of types $A$ and $C$. It would be interesting to study (e.g. using Vogan diagrams [Chu13]) how the construction changes for non-compact real forms.
4.2.1 $\operatorname{SU}(m|n)$
The Lie superalgebra $\mathfrak{gl}(m|n)^{\mathbb{C}}$ has a basis consisting of matrix units $E_{ij}$, $i,j\in\{1,\ldots,m+n\}$, satisfying
$[E_{ij},E_{k\ell}]=\delta_{jk}E_{i\ell}-(-1)^{[E_{ij}][E_{k\ell}]}\delta_{i\ell}E_{kj},$
where $E_{ij}]=0$ if $i,j\leq m$ or $i,j>m$, and $E_{ij}]=1$ otherwise. Let $\mathfrak{h}^{\mathbb{C}}$ be the Cartan subalgebra spanned by all diagonal matrices, and denote by $\{\varepsilon_{1},\ldots,\varepsilon_{m+n}\}$ the $(\mathfrak{h}^{\mathbb{C}})^{\vee}$-basis where $\varepsilon_{i}(E_{jj})=\delta_{ij}$ for all $i,j$. For convenience, we introduce $\delta_{\mu}:=\varepsilon_{m+\mu}$ for each $\mu=1,\ldots,n$.
Let $\mathfrak{sl}(m|n)^{\mathbb{C}}:=\{X\in\mathfrak{gl}(m|n)^{\mathbb{C}}\,|\,\operatorname{Str}(X)=0\}$, and let $\mathfrak{b}^{\mathbb{C}}$ be the Borel subsuperalgebra spanned by all upper-triangular matrices. Then, the set of positive roots partitions as $\Delta^{+}=\Delta_{\bar{0}}^{+}\cup\Delta_{\bar{1}}^{+}$, where
$\Delta_{\bar{0}}^{+}=\{\varepsilon_{i}-\varepsilon_{j},\delta_{\mu}-\delta_{\nu}\,|\,1\leq i<j\leq m;1\leq\mu<\nu\leq n\},\qquad\Delta_{\bar{1}}^{+}=\{\varepsilon_{i}-\delta_{\mu}\,|\,1\leq i\leq m;1\leq\mu\leq n\},$
with simple root system
$\Pi=\{\alpha_{i}=\varepsilon_{i}-\varepsilon_{i+1},\alpha_{m}=\varepsilon_{m}-\delta_{1},\alpha_{m+\mu}=\delta_{\mu}-\delta_{\mu+1}\,|\,1\leq i\leq m-1;1\leq\mu\leq n-1\}$
and corresponding Dynkin diagram
$\alpha_{1}$$\alpha_{m-1}$$\alpha_{m}$$\alpha_{m+1}$$\alpha_{m+n-1}$$11111$
The numbers in the Dynkin diagram indicate the multiplicities in the decomposition of the highest root: $\varepsilon_{1}-\delta_{n}=\alpha_{1}+\cdots+\alpha_{m+n-1}$. The Weyl vector follows from
$2\rho=\sum_{i=1}^{m}(m-n-2i+1)\varepsilon_{i}+\sum_{\mu=1}^{n}(m+n-2\mu+1)\delta_{\mu}.$
Recall from [SNR77, GZ90b] that the so-called type (1) star operation on $\mathfrak{gl}(m|n)^{\mathbb{C}}$ is given by
$(E_{ij})^{*}=E_{ji},\qquad i,j\in\{1,\ldots,m+n\}.$ (4.5)
Using (2.4), we then obtain a compact real form of $\mathfrak{gl}(m|n)^{\mathbb{C}}$, denoted by $\mathfrak{u}(m|n)$, with $\mathbb{R}$-basis
$\{\imath E_{ii}\,|\,i=1,\ldots,m+n\}\cup\{A_{ij},B_{ij}\,|\,1\leq i<j\leq m\}\cup\{A_{ij},B_{ij}\,|\,m+1\leq i<j\leq m+n\}$
$\cup\,\{\sqrt{\imath}A_{ij},\sqrt{\imath}B_{ij}\,|\,1\leq i\leq m;\,m+1\leq j\leq m+n\},$
where $A_{ij}=E_{ij}-E_{ji}$ and $B_{ij}=\imath(E_{ij}+E_{ji})$; cf. (4.4).
The compact real form $\mathfrak{su}(m|n)$ of $\mathfrak{sl}(m|n)^{\mathbb{C}}$ consists of all matrices of $\mathfrak{u}(m|n)$ with supertrace zero, and we assume that $m\geq 3$ if $m=n$ (since for $m=n=2$, each odd root space has dimension $2$). The corresponding Lie supergroup is given by the Harish-Chandra pair
$\operatorname{SU}(m|n)=(\operatorname{SU}(m)\times\operatorname{SU}(n)\times\operatorname{U}(1),\mathfrak{su}(m|n)).$ (4.6)
The Killing form on $\mathfrak{sl}(m|n)^{\mathbb{C}}$ is given by
$B(X,Y)=2(m-n)\operatorname{Str}(XY),\qquad X,Y\in\mathfrak{sl}(m|n)^{\mathbb{C}}.$ (4.7)
In what follows, the non-degenerate supersymmetric even bilinear form $Q$ on $\mathfrak{sl}(m|n)^{\mathbb{C}}$ is defined by
$Q(X,Y)=-\operatorname{Str}(XY),\qquad\text{if $m\neq n$}.$ (4.8)
For $m=n$, $B$ is identically zero, and the Lie superalgebra $\mathfrak{sl}(n|n)^{\mathbb{C}}$ has a one-dimensional centre spanned by the identity matrix $I_{2n}$. The corresponding quotient Lie superalgebra $\mathfrak{sl}(n|n)^{\mathbb{C}}/\mathbb{C}I_{2n}$ is simple, and we define $Q$ on $\mathfrak{sl}(n|n)^{\mathbb{C}}=\mathbb{C}I_{2n}\oplus\mathfrak{sl}(n|n)^{\mathbb{C}}/\mathbb{C}I_{2n}$ by
$Q(I_{2n},I_{2n})=1,\qquad Q(I_{2n},X)=0,\qquad Q(X,Y)=-\operatorname{Str}(XY),\qquad X,Y\in\mathfrak{sl}(n|n)/\mathbb{C}I_{2n}.$ (4.9)
Example 4.1.
Consider $G=\operatorname{SU}(3|2)$ and the circled Dynkin diagram
$\varepsilon_{1}-\varepsilon_{2}$$\varepsilon_{2}-\varepsilon_{3}$$\varepsilon_{3}-\delta_{1}$$\delta_{1}-\delta_{2}$
We now construct the corresponding flag supermanifold $M=G/K$.
The Dynkin diagram corresponds to the complex Lie superalgebra $\mathfrak{g}^{\mathbb{C}}=\mathfrak{sl}(3|2)^{\mathbb{C}}$, while the Dynkin diagram obtained by removing the circled node corresponds to the Lie superalgebra $\mathfrak{k}^{\mathbb{C}}=\mathfrak{gl}(2)^{\mathbb{C}}\oplus\mathfrak{sl}(1|2)^{\mathbb{C}}$. The star operation gives rise to the real forms $\mathfrak{g}=\mathfrak{su}(3|2)$ and $\mathfrak{k}=\mathfrak{su}(2)\oplus\mathfrak{su}(1|2)\oplus\mathfrak{u}(1)\cong\mathfrak{u}(2)\oplus\mathfrak{su}(1|2)$. Thus, $K=\operatorname{SU}(2)\times\operatorname{SU}(1|2)\times\operatorname{U}(1)\cong\operatorname{U}(2)\times\operatorname{SU}(1|2)$, and the root systems of $G$ and $K$ are given by
$\Delta$ $=\big\{\!\pm\!(\varepsilon_{1}-\varepsilon_{2}),\pm(\varepsilon_{2}-\varepsilon_{3}),\pm(\varepsilon_{1}-\varepsilon_{3}),\pm(\delta_{1}-\delta_{2})\big\}\cup\big\{\!\pm\!(\varepsilon_{i}-\delta_{1}),\pm(\varepsilon_{i}-\delta_{2})\,|\,i=1,2,3\big\},$
$\Delta_{K}$ $=\big\{\!\pm\!(\varepsilon_{1}-\varepsilon_{2}),\pm(\delta_{1}-\delta_{2})\big\}\cup\big\{\!\pm\!(\varepsilon_{3}-\delta_{1}),\pm(\varepsilon_{3}-\delta_{2})\big\},$
respectively. Setting
$\Delta_{M}=\Delta-\Delta_{K}=\big\{\!\pm\!(\varepsilon_{i}-\varepsilon_{3}),\pm(\varepsilon_{i}-\delta_{1}),\pm(\varepsilon_{i}-\delta_{2})\,|\,i=1,2\big\},$
we let $\Delta_{M}^{+}=\Delta^{+}\cap\Delta_{M}$ (respectively $\Delta_{M}^{-}=\Delta^{-}\cap\Delta_{M}$) denote the set of positive (respectively negative) roots of $\Delta_{M}$. It follows from (4.2) that $\mathfrak{m}^{\mathbb{C}}=\mathfrak{m}_{1}^{\mathbb{C}}\oplus\mathfrak{m}_{-1}^{\mathbb{C}}$, where
$\mathfrak{m}_{\pm 1}^{\mathbb{C}}=\bigoplus_{\alpha\in\Delta_{M}^{\pm}}\mathfrak{g}_{\alpha}^{\mathbb{C}}.$
Note that $\mathfrak{m}_{1}^{\mathbb{C}}$ and $\mathfrak{m}_{-1}^{\mathbb{C}}$ are irreducible $\operatorname{ad}_{\mathfrak{k}^{\mathbb{C}}}$-modules, and that they are dual to each other. The highest-weight vector of $\mathfrak{m}_{1}^{\mathbb{C}}$ is $E_{15}$, with highest weight $\varepsilon_{1}-\delta_{2}$. It follows from (4.3) and (4.4) that
$\mathfrak{m}=\sum_{\alpha\in(\Delta_{M}^{+})_{\bar{0}}}\mathbb{R}A_{\alpha}\ \oplus\!\sum_{\alpha\in(\Delta_{M}^{+})_{\bar{0}}}\mathbb{R}B_{\alpha}\ \oplus\!\sum_{\alpha\in(\Delta_{M}^{+})_{\bar{1}}}\mathbb{R}\sqrt{\imath}A_{\alpha}\ \oplus\!\sum_{\alpha\in(\Delta_{M}^{+})_{\bar{1}}}\mathbb{R}\sqrt{\imath}B_{\alpha}.$
Since $\mathfrak{m}_{1}^{\mathbb{C}}$ is an irreducible $\operatorname{ad}_{\mathfrak{k}^{\mathbb{C}}}$-module, the real tangent space $\mathfrak{m}$ is irreducible as an $\operatorname{ad}_{\mathfrak{k}}$-module.
We extend the analysis above to general $\operatorname{SU}(m|n)$ with a single circled node in the Dynkin diagram in Section 4.3, and with two circled nodes in Section $4.$ A similar analysis is carried out for $\operatorname{SOSp}(2|2n)$ in Section 4.5, although for a single circled node only.
4.2.2 $\mathrm{SOSp}(2|2n)$
Let $V$ be a complex vector superspace with $\dim(V_{\bar{0}})=2$ and $\dim(V_{\bar{1}})=2n$, where $n\geq 2$, and let $\omega:V\times V\to\mathbb{C}$ be a non-degenerate supersymmetric even bilinear form whose matrix representation relative to a suitable basis for $V$ is given by
$\begin{pmatrix}0&1&0&0\\ 1&0&0&0\\ 0&0&0&I_{n}\\ 0&0&-I_{n}&0\end{pmatrix}$.
The corresponding orthosymplectic superalgebra is defined by
$\mathfrak{osp}(2|2n)^{\mathbb{C}}:=\big\{X\in\mathfrak{gl}(2|2n)^{\mathbb{C}}\,|\,\omega(Xu,v)+(-1)^{[X][u]}\omega(u,Xv)=0\text{ for all }u,v\in V\big\}.$
In the remainder of this subsection, we set $\mathfrak{g}^{\mathbb{C}}=\mathfrak{osp}(2|2n)^{\mathbb{C}}$.
Let $\mathfrak{h}^{\mathbb{C}}$ be the Cartan subalgebra spanned by the diagonal matrices of $\mathfrak{g}^{\mathbb{C}}$:
$\mathfrak{h}^{\mathbb{C}}=\big\{\mathrm{diag}(a,-a,f_{1},\dots,f_{n},-f_{1},\dots,-f_{n})\,|\,a,f_{1},\ldots,f_{n}\in\mathbb{C}\big\}.$ (4.10)
Relative to an arbitrary but fixed nonzero $H\in\mathfrak{h}^{\mathbb{C}}$, define the dual basis $\{\varepsilon,\delta_{1},\dots,\delta_{n}\}\subset(\mathfrak{h}^{\mathbb{C}})^{\vee}$ by
$\varepsilon(H)=a,\qquad\delta_{j}(H)=f_{j},\qquad j=1,\ldots,n.$
The root system partitions as $\Delta=\Delta^{+}\cup\Delta^{-}$, with the positive root system partitioning as $\Delta^{+}=\Delta^{+}_{\bar{0}}\cup\Delta^{+}_{\bar{1}}$, where
$\Delta_{\bar{0}}^{+}=\{\delta_{i}\pm\delta_{j}\,|\,1\leq i<j\leq n\}\cup\{2\delta_{j}\,|\,1\leq j\leq n\},\qquad\Delta_{\bar{1}}^{+}=\{\varepsilon\pm\delta_{j}\,|\,1\leq j\leq n\}.$
The corresponding simple root system is given by
$\Pi=\{\alpha_{1}=\varepsilon-\delta_{1},\alpha_{2}=\delta_{1}-\delta_{2},\dots,\alpha_{n}=\delta_{n-1}-\delta_{n},\alpha_{n+1}=2\delta_{n}\},$
with Dynkin diagram
[…]
The numbers in the Dynkin diagram indicate the multiplicities in the decomposition of the highest root: $\varepsilon+\delta_{1}=\alpha_{1}+2\alpha_{2}+\cdots+2\alpha_{n}+\alpha_{n+1}$. The Weyl vector is given by
$\rho=-n\varepsilon+\sum_{j=1}^{n}(n-j+1)\delta_{j}.$
For any root $\alpha\in\Delta$, let $E_{\alpha}\in\mathfrak{gl}(2|2n)^{\mathbb{C}}$ span the (one-dimensional) root space $\mathfrak{g}_{\alpha}^{\mathbb{C}}$, and let
$(E_{\alpha})^{*}=E_{-\alpha}\quad(\alpha\in\Delta^{+}),\qquad(H)^{*}=H\quad(H\in\mathfrak{h}^{\mathbb{C}}),$
define a type (1) star operation [SNR77, GZ90a]. In light of (2.4), the compact real form $\mathfrak{osp}(2|2n)$ of $\mathfrak{g}^{\mathbb{C}}$ satisfies
$\mathfrak{osp}(2|2n)_{\bar{0}}=\mathfrak{so}(2,\mathbb{R})\oplus\mathfrak{sp}(2n)\cong\mathfrak{so}(2,\mathbb{R})\oplus\mathfrak{u}(n,\mathbb{H}),\qquad\mathfrak{osp}(2|2n)_{\bar{1}}\cong\mathbb{H}^{n}.$
The connected orthosymplectic supergroup $\operatorname{SOSp}(2|2n)$ is given by the Harish-Chandra pair
$\operatorname{SOSp}(2|2n)=(\operatorname{SO}(2,\mathbb{R})\times\operatorname{Sp}(2n),\mathfrak{osp}(2|2n)),$ (4.11)
where $\operatorname{Sp}(2n)=\operatorname{Sp}(2n,\mathbb{C})\cap\operatorname{U}(2n)\cong\operatorname{U}(n,\mathbb{H})$ is our notation for the compact symplectic group.
As in the case of $\operatorname{SU}(m|n)$, we choose the non-degenerate $\operatorname{ad}_{\mathfrak{g}^{\mathbb{C}}}$-invariant supersymmetric even bilinear form to be
$Q(X,Y)=-\operatorname{Str}(XY),\qquad X,Y\in\mathfrak{osp}(2|2n)^{\mathbb{C}},$ (4.12)
and note that it is related to the Killing form $B$ as
$B(X,Y)=2n\,Q(X,Y).$ (4.13)
4.3 $\operatorname{SU}(m|n)$: one isotropy summand
Let $G=\operatorname{SU}(m|n)$ be as in (4.6). We will construct a flag supermanifold $M=G/K$ whose isotropy representation is an irreducible $\operatorname{ad}_{\mathfrak{k}}$-module, here denoted by $\mathfrak{m}_{1}$ (the index is for ease of comparison with results presented in Section 4.4 and Section 4.5). Retaining the notation from Section 3.6 and Section 4.2.1, a $G$-invariant graded Riemannian metric on $M$ is then of the form
$g=x_{1}Q|_{\mathfrak{m}_{1}},\qquad x_{1}\in\mathbb{R}^{\times}.$
Using (4.7), (4.8) and (4.9), the coefficient in the expression (3.25) is $b_{1}=2(m-n)$.
Generalising Example 4.1, we construct $M$ using a Dynkin diagram of $G$ with the $p^{\text{th}}\!$ node circled, where
$1\leq p\leq m+n-1,$
and we divide our considerations into two cases, with $K$ given as follows:
$K=\begin{cases}\operatorname{U}(p|0)\times\operatorname{SU}(m-p|n),\qquad&p\leq m,\\[2.84544pt] \operatorname{SU}(m|p-m)\times\operatorname{U}(0|m+n-p),\qquad&m<p.\end{cases}$ (4.14)
Note that $\operatorname{SU}(k|0)\cong\operatorname{SU}(0|k)\cong\operatorname{SU}(k)$ is an ordinary Lie group for $k\geq 1$. However, writing $\operatorname{U}(p|0)$, for example, keeps track of the concrete embedding of $\operatorname{U}(p)$ inside $G$.
Recall that $\mathfrak{m}^{\mathbb{C}}$ is the $Q$-orthogonal complement of $\mathfrak{k}^{\mathbb{C}}$ in $\mathfrak{g}^{\mathbb{C}}=\mathfrak{sl}(m|n)^{\mathbb{C}}$, where, with $K$ given in (4.14),
$\mathfrak{k}^{\mathbb{C}}=\begin{cases}\mathfrak{gl}(p|0)^{\mathbb{C}}\oplus\mathfrak{sl}(m-p|n)^{\mathbb{C}},\qquad&p\leq m,\\[4.26773pt] \mathfrak{sl}(m|p-m)^{\mathbb{C}}\oplus\mathfrak{gl}(0|m+n-p)^{\mathbb{C}},\qquad&m<p.\end{cases}$
First, let us focus on $p\leq m$ and introduce $\hat{\mathfrak{g}}^{\mathbb{C}}=\mathfrak{gl}(m|n)^{\mathbb{C}}$ and $\hat{\mathfrak{k}}^{\mathbb{C}}=\mathfrak{gl}(p|0)^{\mathbb{C}}\oplus\mathfrak{gl}(m-p|n)^{\mathbb{C}}$. Then, $\hat{\mathfrak{g}}^{\mathbb{C}}=\hat{\mathfrak{k}}^{\mathbb{C}}\oplus\mathfrak{m}^{\mathbb{C}}$ is a $Q$-orthogonal decomposition, and the decomposition of $\mathfrak{m}^{\mathbb{C}}$ as an $\operatorname{ad}_{\hat{\mathfrak{k}}^{\mathbb{C}}}$-module is the same as the decomposition of $\mathfrak{m}^{\mathbb{C}}$ as an $\operatorname{ad}_{\mathfrak{k}^{\mathbb{C}}}$-module. So, we may (and will) continue our analysis using $\hat{\mathfrak{g}}^{\mathbb{C}}$ and $\hat{\mathfrak{k}}^{\mathbb{C}}$ instead of $\mathfrak{g}^{\mathbb{C}}$ and $\mathfrak{k}^{\mathbb{C}}$. For all applicable $r,s$, we use $\mu_{r|s}$ to denote the natural representation of $\mathfrak{gl}(r|s)^{\mathbb{C}}$, with dual representation $\bar{\mu}_{r|s}$. In accord with the regular embedding $\hat{\mathfrak{k}}^{\mathbb{C}}\subseteq\hat{\mathfrak{g}}^{\mathbb{C}}$, we denote by $\varepsilon_{1},\dots,\varepsilon_{p}$ the weights of $\mu_{p|0}$, and by $\varepsilon_{p+1},\dots,\varepsilon_{m},\delta_{1},\dots,\delta_{n}$ the weights of $\mu_{m-p|n}$. It then follows that $\mathfrak{m}^{\mathbb{C}}_{1}$ is indeed an irreducible $\operatorname{ad}_{\mathfrak{k}^{\mathbb{C}}}$-module, with
$\mathfrak{m}^{\mathbb{C}}=\mathfrak{m}^{\mathbb{C}}_{-1}\oplus\mathfrak{m}^{\mathbb{C}}_{1},\qquad\mathfrak{m}^{\mathbb{C}}_{1}=\begin{cases}\mu_{p|0}\otimes\bar{\mu}_{m-p|n},\qquad&p\leq m,\\[4.26773pt] \mu_{m|p-m}\otimes\bar{\mu}_{0|m+n-p},\qquad&p>m,\end{cases}$
where $\mathfrak{m}^{\mathbb{C}}_{-1}$ is the dual $\operatorname{ad}_{\mathfrak{k}^{\mathbb{C}}}$-module of $\mathfrak{m}^{\mathbb{C}}_{1}$, and where we have included the expression for $\mathfrak{m}^{\mathbb{C}}_{1}$ obtained similarly in the case $p>m$. For all $p$, the corresponding highest weight of $\mathfrak{m}^{\mathbb{C}}_{1}$ is $\Lambda_{1}=\varepsilon_{1}-\delta_{n}$. We also have
$d_{1}:=\operatorname{sdim}(\mathfrak{m}_{1})=2\operatorname{sdim}_{\mathbb{C}}(\mathfrak{m}_{1}^{\mathbb{C}})=\begin{cases}2p(m-p-n),\quad&p\leq m,\\[4.26773pt] -2(2m-p)(m+n-p),\quad&p>m.\end{cases}$
The eigenvalue of $C_{\mathfrak{m}_{1},Q|_{\mathfrak{k}}}$ is given by $c_{1}=-(\Lambda_{1}+2\rho_{\mathfrak{k}^{\mathbb{C}}},\Lambda_{1})$, where the Weyl vector $\rho_{\mathfrak{k}^{\mathbb{C}}}$ of $\mathfrak{k}^{\mathbb{C}}$ (see (4.1)) follows from
$2\rho_{\mathfrak{k}^{\mathbb{C}}}=\sum_{i=1}^{p}(p-2i+1)\varepsilon_{i}+\sum_{i=p+1}^{m}(m+p-n-2i+1)\varepsilon_{i}+\sum_{\mu=1}^{n}(m-p+n-2\mu+1)\delta_{\mu}$
for $p\leq m$, while for $p>m$, we have
$2\rho_{\mathfrak{k}^{\mathbb{C}}}$ $=\sum_{i=1}^{m}(2m-p-2i+1)\varepsilon_{i}+\sum_{\mu=1}^{p-m}(p-2\mu+1)\delta_{\mu}+\sum_{\mu=p-m+1}^{n}(-m+p+n-2\mu+1)\delta_{\mu}.$
To compute $c_{1}$, we use that
$(\lambda,\mu)=\sum_{i=1}^{m+n}\lambda(E_{ii})\mu(\bar{E}_{ii}),\qquad\lambda,\mu\in(\mathfrak{h}^{\mathbb{C}})^{\vee},$
where the right $Q$-dual matrices, $\bar{E}_{11},\ldots,\bar{E}_{m+n,m+n}$, to the diagonal matrix units, $E_{11},\ldots,E_{m+n,m+n}$, are given by $\bar{E}_{ii}=-E_{ii}$ if $i\leq m$ and $\bar{E}_{ii}=E_{ii}$ if $i>m$. We thus have $(\varepsilon_{1},\varepsilon_{1})=-1$ and $(\delta_{n},\delta_{n})=1$, in particular. It follows that, for $p\leq m$,
$c_{1}=-(\Lambda_{1}+2\rho_{\mathfrak{k}^{\mathbb{C}}},\Lambda_{1})=-(2\rho_{\mathfrak{k}^{\mathbb{C}}},\varepsilon_{1}-\delta_{n})=-(p-1)(\varepsilon_{1},\varepsilon_{1})+(m-p-n+1)(\delta_{n},\delta_{n})=m-n,$
and similarly for $p>m$,
$c_{1}=-(2m-p-1)(\varepsilon_{1},\varepsilon_{1})-(m-p+n-1)(\delta_{n},\delta_{n})=m-n.$
In both cases, we thus have
$b_{1}=2c_{1},$
so by Proposition 3.25, the single structure constant vanishes: $[111]=0$.
The Ricci coefficient $r_{1}$ follows from (3.34) if $d_{1}\neq 0$, and from Proposition 3.28 if $d_{1}=0$. In both situations, we have $r_{1}=c_{1}$, so the Einstein equation (3.28) simply reads
$c_{1}=cx_{1}.$
For $m\neq n$, there is a unique solution up to scaling, namely $(g,c)=((m-n)Q|_{\mathfrak{m}_{1}},1)$, and the metric is positive (see text following (3.26)) with positive $c$ if $m>n$, and positive with negative $c$ if $m<n$. For $m=n$, there is a continuous family of solutions, namely $(g,c)=(x_{1}Q|_{\mathfrak{m}_{1}},0)$, so the metric is Ricci-flat.
Circling a single node in the Dynkin diagram of $\operatorname{SU}(m|n)$ results in a flag supermanifold with one isotropy summand. In the next section, we will show that circling two nodes leads to three isotropy summands.
4.4 $\operatorname{SU}(m|n)$: three isotropy summands
Let $G=\operatorname{SU}(m|n)$ be as in (4.6). We will construct a flag supermanifold $M=G/K$ whose isotropy representation decomposes into three inequivalent nonzero irreducible $\operatorname{ad}_{\mathfrak{k}}$-modules:
$\mathfrak{m}=\mathfrak{m}_{1}\oplus\mathfrak{m}_{2}\oplus\mathfrak{m}_{3}.$
Retaining the notation from Section 3.6 and Section 4.2.1, a $G$-invariant graded Riemannian metric on $M$ is of the form
$g=x_{1}Q|_{\mathfrak{m}_{1}}+x_{2}Q|_{\mathfrak{m}_{2}}+x_{3}Q|_{\mathfrak{m}_{3}},\qquad x_{1},x_{2},x_{3}\in\mathbb{R}^{\times}.$ (4.15)
We will use $(x_{1},x_{2},x_{3})$ as a shorthand notation for the sum in (4.15). Using (4.7), (4.8) and (4.9), the coefficients in the expression (3.25) are $b_{1}=b_{2}=b_{3}=2(m-n)$. In the following, we will denote this common value simply by $b$:
$b=2(m-n).$
We construct $M$ using a Dynkin diagram of $G$ with two circled nodes, with $p$ and $q$,
$1\leq p<q\leq m+n-1,$
denoting the indices of the circled nodes, and we divide our considerations into three cases, as depicted in Figure 1, with $K$ given as follows:
$K=\begin{cases}\operatorname{U}(p|0)\times\operatorname{U}(q-p|0)\times\operatorname{SU}(m-q|n),\qquad&p<q\leq m,\qquad(\text{Case 1}),\\[2.84544pt] \operatorname{U}(p|0)\times\operatorname{SU}(m-p|q-m)\times\operatorname{U}(0|m+n-q),\qquad&p<m<q,\qquad(\text{Case 2}),\\[2.84544pt] \operatorname{SU}(m|p-m)\times\operatorname{U}(0|q-p)\times\operatorname{U}(0|m+n-q),\qquad&m\leq p<q,\qquad(\text{Case 3}).\end{cases}$ (4.16)
Note that the odd isotropic node is circled if either $q=m$ (Case 1) or $p=m$ (Case 3).
[…]
Figure 1: Dynkin diagrams of $\operatorname{SU}(m|n)$ with two circled nodes.
In Section 4.4.4, we use results obtained in Section 3.6 to classify the $G$-invariant Einstein metrics (4.15) on $M=G/K$ constructed from a circled Dynkin diagram of $G=\operatorname{SU}(m|n)$, with $K$ given in (4.16). In preparation for these classification results, we examine the structure of the isotropy representation of $M$ in Section 4.4.1, and confirm that there are precisely three inequivalent isotropy summands in $T_{K}M$ for $M$ constructed as above. We subsequently determine the structure constants of $M$ in Section 4.4.2.
4.4.1 Isotropy summands
To describe the $\operatorname{ad}_{\mathfrak{k}}$-module structure of $\mathfrak{m}$, we first characterise the $\operatorname{ad}_{\mathfrak{k}^{\mathbb{C}}}$-module structure of $\mathfrak{m}^{\mathbb{C}}$. Recall that $\mathfrak{m}^{\mathbb{C}}$ is the $Q$-orthogonal complement of $\mathfrak{k}^{\mathbb{C}}$ in $\mathfrak{g}^{\mathbb{C}}=\mathfrak{sl}(m|n)^{\mathbb{C}}$, where, with $K$ given in (4.16),
$\mathfrak{k}^{\mathbb{C}}=\begin{cases}\mathfrak{gl}(p|0)^{\mathbb{C}}\oplus\mathfrak{gl}(q-p|0)^{\mathbb{C}}\oplus\mathfrak{sl}(m-q|n)^{\mathbb{C}},\qquad&p<q\leq m,\\[4.26773pt] \mathfrak{gl}(p|0)^{\mathbb{C}}\oplus\mathfrak{sl}(m-p|q-m)^{\mathbb{C}}\oplus\mathfrak{gl}(0|m+n-q)^{\mathbb{C}},\qquad&p<m<q,\\[4.26773pt] \mathfrak{sl}(m|p-m)^{\mathbb{C}}\oplus\mathfrak{gl}(0|q-p)^{\mathbb{C}}\oplus\mathfrak{gl}(0|m+n-q)^{\mathbb{C}},\qquad&m\leq p<q.\end{cases}$ (4.17)
For the root-space decomposition (4.2) of $\mathfrak{m}^{\mathbb{C}}$, we introduce
$\mathfrak{m}_{k_{1},k_{2}}^{\mathbb{C}}:=\bigoplus_{\alpha\in\Delta_{k_{1},k_{2}}}\mathfrak{g}_{\alpha}^{\mathbb{C}},\qquad\Delta_{k_{1},k_{2}}:=\Big\{\sum_{i=1}^{m+n-1}c_{i}\alpha_{i}\in\Delta_{M}\,|\,c_{p}=k_{1},c_{q}=k_{2}\Big\},\qquad(k_{1},k_{2})\in I_{\mathfrak{m}^{\mathbb{C}}},$
where $I_{\mathfrak{m}^{\mathbb{C}}}=\{(-1,-1),(0,-1),(-1,0),(1,0),(0,1),(1,1)\}$. If $\Delta_{k_{1},k_{2}}$ is non-empty, then $\mathfrak{m}_{k_{1},k_{2}}^{\mathbb{C}}$ is an $\operatorname{ad}_{\mathfrak{k}^{\mathbb{C}}}$-invariant subspace of $\mathfrak{m}^{\mathbb{C}}$.
Proposition 4.2.
Let $\mathfrak{k}^{\mathbb{C}}$ be one of the Lie superalgebras in (4.17). As an $\operatorname{ad}_{\mathfrak{k}^{\mathbb{C}}}$-module, $\mathfrak{m}^{\mathbb{C}}$ admits the $Q$-orthogonal decomposition
$\mathfrak{m}^{\mathbb{C}}=\mathfrak{m}_{-1,-1}^{\mathbb{C}}\oplus\mathfrak{m}_{0,-1}^{\mathbb{C}}\oplus\mathfrak{m}_{-1,0}^{\mathbb{C}}\oplus\mathfrak{m}_{1,0}^{\mathbb{C}}\oplus\mathfrak{m}_{0,1}^{\mathbb{C}}\oplus\mathfrak{m}_{1,1}^{\mathbb{C}}.$
Moreover, $\mathfrak{m}_{k_{1},k_{2}}^{\mathbb{C}}$ and $\mathfrak{m}_{-k_{1},-k_{2}}^{\mathbb{C}}$ are irreducible and dual for each $(k_{1},k_{2})\in\{(1,0),(0,1),(1,1)\}$.
Proof.
We present a proof in Case 1 ($p<q\leq m$); the other two cases can be treated similarly. Let $\hat{\mathfrak{g}}^{\mathbb{C}}=\mathfrak{gl}(m|n)^{\mathbb{C}}$ and $\hat{\mathfrak{k}}^{\mathbb{C}}=\mathfrak{gl}(p|0)^{\mathbb{C}}\oplus\mathfrak{gl}(q-p|0)^{\mathbb{C}}\oplus\mathfrak{gl}(m-q|n)^{\mathbb{C}}$. Then, $\hat{\mathfrak{g}}^{\mathbb{C}}=\hat{\mathfrak{k}}^{\mathbb{C}}\oplus\mathfrak{m}^{\mathbb{C}}$ is a $Q$-orthogonal decomposition, and the decomposition of $\mathfrak{m}^{\mathbb{C}}$ as an $\operatorname{ad}_{\hat{\mathfrak{k}}^{\mathbb{C}}}$-module is the same as the decomposition of $\mathfrak{m}^{\mathbb{C}}$ as an $\operatorname{ad}_{\mathfrak{k}^{\mathbb{C}}}$-module. So, we may (and will) continue our analysis using $\hat{\mathfrak{g}}^{\mathbb{C}}$ and $\hat{\mathfrak{k}}^{\mathbb{C}}$ instead of $\mathfrak{g}^{\mathbb{C}}$ and $\mathfrak{k}^{\mathbb{C}}$.
In accord with the regular embedding $\hat{\mathfrak{k}}^{\mathbb{C}}\subseteq\hat{\mathfrak{g}}^{\mathbb{C}}$, we denote by $\varepsilon_{1},\dots,\varepsilon_{p}$ the weights of the natural representation $\mu_{p|0}$, by $\varepsilon_{p+1},\dots,\varepsilon_{q}$ the weights of $\mu_{q-p|0}$, and by $\varepsilon_{q+1},\dots,\varepsilon_{m},\delta_{1},\dots,\delta_{n}$ the weights of $\mu_{m-q|n}$. We then have
$\hat{\mathfrak{k}}^{\mathbb{C}}\cong(\mu_{p|0}\otimes\bar{\mu}_{p|0})\oplus(\mu_{q-p|0}\otimes\bar{\mu}_{q-p|0})\oplus(\mu_{m-q|n}\otimes\bar{\mu}_{m-q|n}).$
Restricting to $\hat{\mathfrak{k}}^{\mathbb{C}}$, we have $\mu_{m|n}|_{\hat{\mathfrak{k}}^{\mathbb{C}}}\cong\mu_{p|0}\oplus\mu_{q-p|0}\oplus\mu_{m-q|n}$ and hence the $\hat{\mathfrak{k}}^{\mathbb{C}}$-module isomorphisms
$\hat{\mathfrak{g}}^{\mathbb{C}}|_{\operatorname{ad}_{\hat{\mathfrak{k}}^{\mathbb{C}}}}\cong\mu_{m|n}|_{\hat{\mathfrak{k}}^{\mathbb{C}}}\otimes\bar{\mu}_{m|n}|_{\hat{\mathfrak{k}}^{\mathbb{C}}}\cong(\mu_{p|0}\oplus\mu_{q-p|0}\oplus\mu_{m-q|n})\otimes(\bar{\mu}_{p|0}\oplus\bar{\mu}_{q-p|0}\oplus\bar{\mu}_{m-q|n}).$
Expanding this out, $\hat{\mathfrak{k}}^{\mathbb{C}}$ arises as direct summand, with remainder
$\mathfrak{m}^{\mathbb{C}}=\mathfrak{m}_{-1,-1}^{{}^{\prime}\mathbb{C}}\oplus\mathfrak{m}_{0,-1}^{{}^{\prime}\mathbb{C}}\oplus\mathfrak{m}_{-1,0}^{{}^{\prime}\mathbb{C}}\oplus\mathfrak{m}_{1,0}^{{}^{\prime}\mathbb{C}}\oplus\mathfrak{m}_{0,1}^{{}^{\prime}\mathbb{C}}\oplus\mathfrak{m}_{1,1}^{{}^{\prime}\mathbb{C}},$
where
$\mathfrak{m}^{{}^{\prime}\mathbb{C}}_{-1,-1}\cong\bar{\mu}_{p|0}\otimes\mu_{m-q|n},$ $\mathfrak{m}^{{}^{\prime}\mathbb{C}}_{0,-1}\cong\bar{\mu}_{q-p|0}\otimes\mu_{m-q|n},$ $\mathfrak{m}^{{}^{\prime}\mathbb{C}}_{-1,0}\cong\bar{\mu}_{p|0}\otimes\mu_{q-p|0},$
$\mathfrak{m}^{{}^{\prime}\mathbb{C}}_{1,0}\cong\mu_{p|0}\otimes\bar{\mu}_{q-p|0},$ $\mathfrak{m}^{{}^{\prime}\mathbb{C}}_{0,1}\cong\mu_{q-p|0}\otimes\bar{\mu}_{m-q|n},$ $\mathfrak{m}^{{}^{\prime}\mathbb{C}}_{1,1}\cong\mu_{p|0}\otimes\bar{\mu}_{m-q|n}.$
It follows that $\mathfrak{m}_{k_{1},k_{2}}^{{}^{\prime}\mathbb{C}}$ and $\mathfrak{m}_{-k_{1},-k_{2}}^{{}^{\prime}\mathbb{C}}$ are irreducible and dual to each other.
We now show that $\mathfrak{m}^{\mathbb{C}}_{k_{1},k_{2}}=\mathfrak{m}^{{}^{\prime}\mathbb{C}}_{k_{1},k_{2}}$ for all relevant $k_{1},k_{2}$, which then concludes the proof. Note that the weights of $\mathfrak{m}^{{}^{\prime}\mathbb{C}}_{1,0}$ are
$\varepsilon_{i}-\varepsilon_{j}=\alpha_{i}+\cdots+\alpha_{j-1},\qquad 1\leq i\leq p<j\leq q,$
so the coefficient of $\alpha_{p}$ is $1$, while the coefficient of $\alpha_{q}$ is 0. This implies that $\mathfrak{m}^{{}^{\prime}\mathbb{C}}_{1,0}\subseteq\mathfrak{m}^{\mathbb{C}}_{1,0}$. Similarly, it follows that $\mathfrak{m}^{{}^{\prime}\mathbb{C}}_{k_{1},k_{2}}\subseteq\mathfrak{m}^{\mathbb{C}}_{k_{1},k_{2}}$ for all $(k_{1},k_{2})\in I_{\mathfrak{m}^{\mathbb{C}}}$, so
$\bigoplus_{k_{1},k_{2}}\mathfrak{m}_{k_{1},k_{2}}^{\mathbb{C}}\subseteq\mathfrak{m}^{\mathbb{C}}=\bigoplus_{k_{1},k_{2}}\mathfrak{m}_{k_{1},k_{2}}^{{}^{\prime}\mathbb{C}}\subseteq\bigoplus_{k_{1},k_{2}}\mathfrak{m}_{k_{1},k_{2}}^{\mathbb{C}},$
where the direct sums are over $(k_{1},k_{2})\in I_{\mathfrak{m}^{\mathbb{C}}}$. It follows that $\mathfrak{m}^{\mathbb{C}}_{k_{1},k_{2}}=\mathfrak{m}^{{}^{\prime}\mathbb{C}}_{k_{1},k_{2}}$ for all $(k_{1},k_{2})\in I_{\mathfrak{m}^{\mathbb{C}}}$. ∎
Corollary 4.3.
Let the notation be as in Proposition 4.2. Then, the following holds.
  • In Case 1 $(p<q\leq m)$,
    $\mathfrak{m}_{1,0}^{\mathbb{C}}\cong\mu_{p|0}\otimes\bar{\mu}_{q-p|0},\qquad\mathfrak{m}_{0,1}^{\mathbb{C}}\cong\mu_{q-p|0}\otimes\bar{\mu}_{m-q|n},\qquad\mathfrak{m}_{1,1}^{\mathbb{C}}\cong\mu_{p|0}\otimes\bar{\mu}_{m-q|n},$
    as $\operatorname{ad}_{\mathfrak{k}^{\mathbb{C}}}$-modules, with corresponding highest weights
    $\Lambda_{1,0}=\varepsilon_{1}-\varepsilon_{q},\qquad\Lambda_{0,1}=\varepsilon_{p+1}-\delta_{n},\qquad\Lambda_{1,1}=\varepsilon_{1}-\delta_{n}.$
  • In Case 2 $(p<m<q)$,
    $\mathfrak{m}_{1,0}^{\mathbb{C}}\cong\mu_{p|0}\otimes\bar{\mu}_{m-p|q-m},\qquad\mathfrak{m}_{0,1}^{\mathbb{C}}\cong\mu_{m-p|q-m}\otimes\bar{\mu}_{0|m+n-q},\qquad\mathfrak{m}_{1,1}^{\mathbb{C}}\cong\mu_{p|0}\otimes\bar{\mu}_{0|m+n-q},$
    as $\operatorname{ad}_{\mathfrak{k}^{\mathbb{C}}}$-modules, with corresponding highest weights
    $\Lambda_{1,0}=\varepsilon_{1}-\delta_{q-m},\qquad\Lambda_{0,1}=\varepsilon_{p+1}-\delta_{n},\qquad\Lambda_{1,1}=\varepsilon_{1}-\delta_{n}.$
  • In Case 3 $(m\leq p<q)$,
    $\mathfrak{m}_{1,0}^{\mathbb{C}}\cong\mu_{m|p-m}\otimes\bar{\mu}_{0|q-p},\qquad\mathfrak{m}_{0,1}^{\mathbb{C}}\cong\mu_{0|q-p}\otimes\bar{\mu}_{0|m+n-q},\qquad\mathfrak{m}_{1,1}^{\mathbb{C}}\cong\mu_{m|p-m}\otimes\bar{\mu}_{0|m+n-q},$
    as $\operatorname{ad}_{\mathfrak{k}^{\mathbb{C}}}$-modules, with corresponding highest weights
    $\Lambda_{1,0}=\varepsilon_{1}-\delta_{q-m},\qquad\Lambda_{0,1}=\delta_{p-m+1}-\delta_{n},\qquad\Lambda_{1,1}=\varepsilon_{1}-\delta_{n}.$
Proposition 4.4.
In each case in (4.16), the $\operatorname{ad}_{\mathfrak{k}}$-representation $\mathfrak{m}$ admits a $Q$-orthogonal decomposition of the form
$\mathfrak{m}\cong\mathfrak{m}_{1,0}\oplus\mathfrak{m}_{0,1}\oplus\mathfrak{m}_{1,1},$
where the summands are irreducible $\operatorname{ad}_{\mathfrak{k}}$-representations satisfying
$[\mathfrak{m}_{1,0},\mathfrak{m}_{1,0}]\subseteq\mathfrak{k},\qquad[\mathfrak{m}_{0,1},\mathfrak{m}_{0,1}]\subseteq\mathfrak{k},\qquad[\mathfrak{m}_{1,1},\mathfrak{m}_{1,1}]\subseteq\mathfrak{k},$
$[\mathfrak{m}_{1,0},\mathfrak{m}_{1,1}]\subseteq\mathfrak{k}\oplus\mathfrak{m}_{0,1},\qquad[\mathfrak{m}_{0,1},\mathfrak{m}_{1,1}]\subseteq\mathfrak{k}\oplus\mathfrak{m}_{1,0},\qquad[\mathfrak{m}_{1,0},\mathfrak{m}_{0,1}]\subseteq\mathfrak{k}\oplus\mathfrak{m}_{1,1}.$
Proof.
As in (2.4), $\mathfrak{m}_{\bar{0}}$ is spanned by (1)(-1)-eigenvectors of the star operation restricted to $\mathfrak{m}_{\bar{0}}^{\mathbb{C}}$. Using Proposition 4.2, the star operation on $\mathfrak{m}_{\bar{0}}^{\mathbb{C}}$ splits into restrictions on $(\mathfrak{m}_{-1,0}^{\mathbb{C}}\oplus\mathfrak{m}_{1,0}^{\mathbb{C}})_{\bar{0}}$, $(\mathfrak{m}_{0,-1}^{\mathbb{C}}\oplus\mathfrak{m}_{0,1}^{\mathbb{C}})_{\bar{0}}$ and $(\mathfrak{m}_{-1,-1}^{\mathbb{C}}\oplus\mathfrak{m}_{1,1}^{\mathbb{C}})_{\bar{0}}$, respectively. For each pair $(k_{1},k_{2})\in\{(1,0),(0,1),(1,1)\}$, the real (1)(-1)-eigenspace of the star operation on $(\mathfrak{m}_{-k_{1},-k_{2}}^{\mathbb{C}}\oplus\mathfrak{m}_{k_{1},k_{2}}^{\mathbb{C}})_{\bar{0}}$ is spanned by $A_{\alpha}$ and $B_{\alpha}$, $\alpha\in(\Delta_{k_{1},k_{2}}^{+})_{\bar{0}}$, so
$\mathfrak{m}_{\bar{0}}=(\mathfrak{m}_{1,0})_{\bar{0}}\oplus(\mathfrak{m}_{0,1})_{\bar{0}}\oplus(\mathfrak{m}_{1,1})_{\bar{0}}.$
Similarly, $\mathfrak{m}_{\bar{1}}=(\mathfrak{m}_{1,0})_{\bar{1}}\oplus(\mathfrak{m}_{0,1})_{\bar{1}}\oplus(\mathfrak{m}_{1,1})_{\bar{1}}$, so the desired decomposition follows. Moreover, $\mathfrak{m}_{k_{1},k_{2}}$ is an irreducible $\operatorname{ad}_{\mathfrak{k}}$-representation since $\mathfrak{m}_{k_{1},k_{2}}^{\mathbb{C}}$ is an irreducible $\operatorname{ad}_{\mathfrak{k}^{\mathbb{C}}}$-representation. Finally, the commutation relations follow from those among the root vectors $E_{\alpha}$. ∎
By Corollary 4.3, the $\operatorname{ad}_{\mathfrak{k}^{\mathbb{C}}}$-modules $\mathfrak{m}_{1,0}^{\mathbb{C}},\mathfrak{m}_{0,1}^{\mathbb{C}}$ and $\mathfrak{m}_{1,1}^{\mathbb{C}}$ have distinct highest weights and are therefore mutually non-isomorphic. It follows from Proposition 4.4 that the flag supermanifold $G/K$ has three inequivalent irreducible isotropy summands.
It follows from the proof of Proposition 4.4 that, for each $(k_{1},k_{2})\in\{(1,0),(0,1),(1,1)\}$,
$\{A_{\alpha},B_{\alpha}\,|\,\alpha\in\Delta_{k_{1},k_{2}}\cap\Delta_{\bar{0}}^{+}\}\cup\{\sqrt{\imath}A_{\alpha},\sqrt{\imath}B_{\alpha}\,|\,\alpha\in\Delta_{k_{1},k_{2}}\cap\Delta_{\bar{1}}^{+}\}$
is an $\mathbb{R}$-basis for the $\operatorname{ad}_{\mathfrak{k}}$-module $\mathfrak{m}_{k_{1},k_{2}}$, with
$\operatorname{sdim}(\mathfrak{m}_{k_{1},k_{2}})=2\operatorname{sdim}_{\mathbb{C}}(\mathfrak{m}_{k_{1},k_{2}}^{\mathbb{C}}).$ (4.18)
For convenience, we introduce the notation
$\mathfrak{m}_{1}=\mathfrak{m}_{0,1},\qquad\mathfrak{m}_{2}=\mathfrak{m}_{1,1},\qquad\mathfrak{m}_{3}=\mathfrak{m}_{1,0},$ (4.19)
and write
$d_{i}=\operatorname{sdim}(\mathfrak{m}_{i}),\qquad i=1,2,3,$ (4.20)
as in (3.31). Explicit expressions for $d_{1},d_{2},d_{3}$ are readily obtained and are given in Lemma 4.7 below.
4.4.2 Structure constants and Ricci curvature
$c_{1}$ $c_{2}$ $c_{3}$
Case 1: $p<q\leq m$ mnpm-n-p mn+pqm-n+p-q $q$
Case 2: $p<m<q$ mnpm-n-p mn+p+q-m-n+p+q 2mq2m-q
Case 3: $m\leq p<q$ mn+p-m-n+p mnp+qm-n-p+q 2mq2m-q
Table 1: Casimir eigenvalues.
Proposition 4.5.
Let $K$ be as in (4.16) and $\mathfrak{m}$ as in Proposition 4.4 and (4.19). Then, the Casimir eigenvalues (3.30) are given as in Table 1.
Proof.
In each of the three cases, the eigenvalues of $C_{\mathfrak{m}_{i},Q|_{\mathfrak{k}}}$ are given by $c_{i}=-(\Lambda_{i}+2\rho_{\mathfrak{k}^{\mathbb{C}}},\Lambda_{i})$, $i=1,2,3$, where $\rho_{\mathfrak{k}^{\mathbb{C}}}$ is the Weyl vector of $\mathfrak{k}^{\mathbb{C}}$ (see (4.1)), and where $\Lambda_{3},\Lambda_{1},\Lambda_{2}$ are the highest weights given in Corollary 4.3. To compute these eigenvalues, we use that
$(\lambda,\mu)=\sum_{i=1}^{m+n}\lambda(E_{ii})\mu(\bar{E}_{ii}),\qquad\lambda,\mu\in(\mathfrak{h}^{\mathbb{C}})^{\vee},$
where the right $Q$-dual matrices, $\bar{E}_{11},\ldots,\bar{E}_{m+n,m+n}$, to the diagonal matrix units, $E_{11},\ldots,E_{m+n,m+n}$, are given by $\bar{E}_{ii}=-E_{ii}$ if $i\leq m$ and $\bar{E}_{ii}=E_{ii}$ if $i>m$. In Case 1, for example, we have
$2\rho_{\mathfrak{k}^{\mathbb{C}}}$ $=\sum_{i=1}^{p}(p-2i+1)\varepsilon_{i}+\sum_{i=p+1}^{q}\!\big(q-p-2(i-p)+1\big)\varepsilon_{i}$
$\quad+\sum_{i=q+1}^{m}\!\big(m-q-n-2(i-q)+1\big)\varepsilon_{i}+\sum_{\mu=1}^{n}(m-q+n-2\mu+1)\delta_{\mu},$
while the highest weight of $\mathfrak{m}_{2}^{\mathbb{C}}$ is $\Lambda_{2}=\varepsilon_{1}-\delta_{n}$, so
$c_{2}=-(\Lambda_{2}+2\rho_{\mathfrak{k}^{\mathbb{C}}},\Lambda_{2})=-(2\rho_{\mathfrak{k}^{\mathbb{C}}},\varepsilon_{1}-\delta_{n})=-(p-1)(\varepsilon_{1},\varepsilon_{1})+(m-q-n+1)(\delta_{n},\delta_{n})=m-n+p-q.$
The eigenvalues $c_{1}$ and $c_{3}$ are computed similarly, as are the eigenvalues in Cases 2 and 3. ∎
Corollary 4.6.
Let the notation be as in Proposition 4.5. Then,
$b=c_{1}+c_{2}+c_{3}.$
Lemma 4.7.
Let the notation be as in Proposition 4.4, (4.19) and (4.20). Then,
$d_{1}=\tfrac{1}{2}(b-2c_{2})(b-2c_{3}),\qquad d_{2}=\tfrac{1}{2}(b-2c_{1})(b-2c_{3}),\qquad d_{3}=\tfrac{1}{2}(b-2c_{1})(b-2c_{2}).$
Proof.
The result follows from Corollary 4.3, (2.1) and (4.18). ∎
Proposition 4.8.
Let $K$ be as in (4.16) and $\mathfrak{m}$ as in Proposition 4.4 and (4.19). Then, up to permutation, $[123]$ is the only possibly nonzero structure constant, and it is given as
$[123]=\tfrac{1}{4}(b-2c_{1})(b-2c_{2})(b-2c_{3}).$
Proof.
It follows from Proposition 4.4 that $[123]$ is the only possibly nonzero structure constant up to permutation. Using Proposition 3.25 and the symmetry of $[ijk]$, we have $2[123]=[312]+[321]=d_{3}(b_{3}-2c_{3})$. The result now follows using $b_{3}=b$, Proposition 4.5 and Lemma 4.7. ∎
Remark 4.9.
In the super-setting, structure constants can be negative, zero or positive, in contrast to the classical setting where they are always non-negative.
Proposition 4.10.
Let $g=(x_{1},x_{2},x_{3})$ be a $G$-invariant graded Riemannian metric on $M=G/K$, where $G=\operatorname{SU}(m|n)$ and $K$ is given in (4.16). Set $x^{2}=x_{1}^{2}+x_{2}^{2}+x_{3}^{2}$. Then, the following holds.
  • The Ricci coefficients in (3.27) are given by
    $r_{i}=\frac{b}{2}+\frac{(b-2c_{i})x_{i}(2x_{i}^{2}-x^{2})}{4x_{1}x_{2}x_{3}},\qquad i=1,2,3.$
  • The scalar curvature is given by
    $S=\frac{b}{2}\Big(\frac{d_{1}}{x_{1}}+\frac{d_{2}}{x_{2}}+\frac{d_{3}}{x_{3}}\Big)-\frac{[123]\,x^{2}}{2x_{1}x_{2}x_{3}},$
    with $d_{1},d_{2},d_{3}$ given in Lemma 4.7 and $[123]$ in Proposition 4.8.
Proof.
If $123]\neq 0$, then $d_{1},d_{2},d_{3}$ are nonzero and the desired expression for $r_{i}$ follows by applying $b_{1}=b_{2}=b_{3}=b$, Lemma 4.7, Proposition 4.8 and the symmetry of $[ijk]$ to the expression in (3.34). In Case 1, $b-2c_{i}=0$ is only possible if $i=3$, in which case $d_{1}=d_{2}=0$ while $d_{3}\neq 0$; in Case 2, only $b-2c_{2}$ can be zero, in which case $d_{1}=d_{3}=0$ while $d_{2}\neq 0$; and in Case 3, only $b-2c_{1}$ can be zero, in which case $d_{2}=d_{3}=0$ while $d_{1}\neq 0$. By Proposition 4.4, $(2,3)$ is $1$-selected, $(1,3)$ is $2$-selected, and $(1,2)$ is $3$-selected; see (3.35). We can thus use Proposition 3.28 to compute $r_{1}$ and $r_{2}$ in Case 1 if $2c_{3}=b$; $r_{1}$ and $r_{3}$ in Case 2 if $2c_{2}=b$; and $r_{2}$ and $r_{3}$ in Case 3 if $2c_{1}=b$. In every case, the desired expression for $r_{i}$ follows. Finally, the expression for $S$ follows by applying Lemma 4.7, Proposition 4.8 and the symmetry of $[ijk]$ to the expression in Proposition 3.27. ∎
4.4.3 Einstein metrics
Using Proposition 4.10, we can write the Einstein equations (3.28) as
$c=\frac{b}{2x_{i}}+\frac{(b-2c_{i})(2x_{i}^{2}-x^{2})}{4x_{1}x_{2}x_{3}},\qquad i=1,2,3.$ (4.21)
As always, if $(g,c)$ provides a solution to (4.21), then so does $(\lambda g,\frac{c}{\lambda})$ for any $\lambda\in\mathbb{R}^{\times}$. If $b=0$, then simultaneously changing the sign on $c$ and on any one of $x_{1},x_{2},x_{3}$ will also yield a solution. As in the non-super setting, an Einstein metric is Ricci-flat if $c=0$.
Eliminating $c$ by combining the equations (4.21) for $i=2,3$ and $i=1,3$, respectively, and using Corollary 4.6, yields
$(c_{1}x_{23}^{+}-b)x_{23}^{-}$ $=c_{3}-c_{2},$ (4.22)
$(b-c_{1})(x_{23}^{-})^{2}+\big(b-(b-2c_{3})x_{23}^{+}\big)x_{23}^{-}$ $=c_{1}(x_{23}^{+})^{2}-bx_{23}^{+}+2c_{2},$ (4.23)
where $x_{23}^{\pm}:=\frac{x_{2}\pm x_{3}}{x_{1}}$. We now multiply (4.23) by $(c_{1}x_{23}^{+}-b)^{2}$ and then apply (4.22) to replace factors of $(c_{1}x_{23}^{+}-b)x_{23}^{-}$ by $c_{3}-c_{2}$. This turns (4.23) into the quartic relation
$\big[x_{23}^{+}-1\big]\big[c_{1}x_{23}^{+}-(c_{1}+2c_{2})\big]\big[c_{1}x_{23}^{+}-(c_{1}+2c_{3})\big]\big[c_{1}x_{23}^{+}-(c_{2}+c_{3})\big]=0,$
which $x_{23}^{+}$ must satisfy to provide an Einstein metric. For $c_{1},c_{2},c_{3}$ generic, this quartic relation combined with (4.22) yields the following four solutions to the Einstein equations:
$(\mathrm{S1}):\quad(x_{1},x_{2},x_{3})=(c_{2}+c_{3},c_{2},c_{3}),\qquad c=1,$
$(\mathrm{S2}):\quad(x_{1},x_{2},x_{3})=(c_{1},c_{1}+c_{3},c_{3}),\qquad c=1,$
$(\mathrm{S3}):\quad(x_{1},x_{2},x_{3})=(c_{1},c_{2},c_{1}+c_{2}),\qquad c=1,$
$(\mathrm{S4}):\quad(x_{1},x_{2},x_{3})=(c_{1},c_{2},c_{3}),\qquad\qquad c=1+\frac{(b-2c_{1})(b-2c_{2})(b-2c_{3})}{4c_{1}c_{2}c_{3}}.$
In these solutions, we respectively have
$x_{1}=x_{2}+x_{3},\qquad x_{2}=x_{1}+x_{3},\qquad x_{3}=x_{1}+x_{2},\qquad x_{1}+x_{2}+x_{3}=b,$ (4.24)
and we note that the expression for $c$ in (S4) is well-defined as long as $x_{i}\neq 0$, $i=1,2,3$.
Since we must have $x_{1},x_{2},x_{3}\in\mathbb{R}^{\times}$, the Casimir eigenvalues $c_{1},c_{2},c_{3}$ could be related in a way that invalidates one or more of the four solutions. Indeed, if $c_{1}=0$, $c_{2}=0$ or $c_{3}=0$, then the only viable solution is (S1), (S2) or (S3), respectively. Similarly, if $c_{2}+c_{3}=0$, $c_{1}+c_{3}=0$ or $c_{1}+c_{2}=0$, then (S1), (S2), respectively (S3), is inviable. However, in these situations, other solutions may be present. To see this, let us analyse the cases $c_{1}=0$ and $c_{2}+c_{3}=0$, noting that some of the scenarios discussed below cannot arise for the Casimir eigenvalues $c_{1},c_{2},c_{3}$ given in Table 1.
First suppose $c_{1}=0$. If $c_{2}+c_{3}=0$, then, in fact, $c_{2}=c_{3}=0$, so $c=0$ while $g$ is unconstrained. If $c_{2}+c_{3}\neq 0$ and $c_{2}c_{3}=0$, then $c=0$ and $g\in\{(x_{1},x_{2},x_{2}\pm x_{1})\,|\,x_{2}\neq\mp x_{1};\,x_{1}x_{2}\neq 0\}$. If $c_{2}+c_{3}\neq 0$ and $c_{2}c_{3}\neq 0$, then (S1) holds, that is, $c=\frac{1}{\lambda}$ and $g=\lambda(c_{2}+c_{3},c_{2},c_{3})$, $\lambda\in\mathbb{R}^{\times}$.
Now suppose $c_{2}+c_{3}=0$. Since $c_{1}=0$ was covered above, we also suppose $c_{1}\neq 0$. If $c_{3}=0$, then $c=0$ and $g\in\{(x_{1},x_{2},x_{1}-x_{2})\,|\,x_{2}\neq x_{1};\,x_{1}x_{2}\neq 0\}$. If $c_{3}\neq 0$, then (S4) holds, that is, $c=\frac{c_{1}^{2}}{4\lambda c_{3}^{2}}$ and $g=\lambda(c_{1},c_{2},c_{3})$, $\lambda\in\mathbb{R}^{\times}$, provides a solution. However, this is not the only solution if $c_{3}\notin\{0,c_{1}\}$, as then (S3) also holds, that is, $c=\frac{1}{\lambda}$ and $g=\lambda(c_{1},c_{2},c_{1}+c_{2})$, $\lambda\in\mathbb{R}^{\times}$, provides a solution. Further, if $c_{3}\notin\{0,c_{1},-c_{1}\}$, then (S2) also holds, that is, $c=\frac{1}{\lambda}$ and $g=\lambda(c_{1},c_{1}+c_{3},c_{3})$, $\lambda\in\mathbb{R}^{\times}$, provides a solution.
4.4.4 Classification results
Following the analysis in Section 4.4.3, we now summarise the solutions to (4.21) in the specific cases of our interest, the flag supermanifolds $M=G/K$ built from $G=\operatorname{SU}(m|n)$ with $K$ given in (4.16). By invoking permutation symmetry, we have covered all possible scenarios in Section 4.4.3. As established in Section 4.4.1, the corresponding isotropy representation decomposes into a sum of three inequivalent nonzero irreducible $\operatorname{ad}_{\mathfrak{k}}$-modules, $\mathfrak{m}=\mathfrak{m}_{1}\oplus\mathfrak{m}_{2}\oplus\mathfrak{m}_{3}$, with the labelling for the summands given in (4.19) and explained in Section 4.4.1. For all $x_{1},x_{2},x_{3},c\in\mathbb{R}$, we introduce
$[\,x_{1}:x_{2}:x_{3}\mid c\,]=\begin{cases}\big\{(\lambda g,\tfrac{c}{\lambda})\,|\,g=(x_{1},x_{2},x_{3});\,\lambda\in\mathbb{R}^{\times}\big\},\ &\text{if $x_{1}x_{2}x_{3}\neq 0$,}\\[4.26773pt] \emptyset,\ &\text{if $x_{1}x_{2}x_{3}=0$,}\end{cases}$
and
$\mathcal{F}_{1}(u,v)$ $=\begin{cases}\big\{\big((x_{2}+x_{3},x_{2},x_{3}),0\big)\,|\,x_{2}+x_{3}\neq 0;\,x_{2}x_{3}\neq 0\big\},\ &\text{if $u=p\ \&\ v=q$,}\\[4.26773pt] \emptyset,\ &\text{otherwise,}\end{cases}$
$\mathcal{F}_{3}(u,v)$ $=\begin{cases}\big\{\big((x_{1},x_{2},x_{1}+x_{2}),0\big)\,|\,x_{1}+x_{2}\neq 0;\,x_{1}x_{2}\neq 0\big\},\ &\text{if $u=p\ \&\ v=q$,}\\[4.26773pt] \emptyset,\ &\text{otherwise.}\end{cases}$
Theorem 4.11.
Let $M=G/K$ be a flag supermanifold, where $G=\operatorname{SU}(m|n)$ and $K$ is given in (4.16), with $m,n\geq 1$ such that $m+n\geq 4$ and, in the case $m=n$, such that $n\geq 3$. Then, the $G$-invariant Einstein metrics $g$ on $M$, with corresponding Einstein constant $c$, are of the form (4.15) and classified as follows:
$\text{Case 1 }(p<q\leq m):$ $\qquad(g,c)\in\mathcal{C}_{1,1}\cup\mathcal{C}_{1,2}\cup\mathcal{C}_{1,3}\cup\mathcal{C}_{1,4}\cup\mathcal{F}_{3}(m-n,2(m-n)),$
$\text{Case 2 }(p<m<q):$ $\qquad(g,c)\in\mathcal{C}_{2,1}\cup\mathcal{C}_{2,2}\cup\mathcal{C}_{2,3}\cup\mathcal{C}_{2,4}\cup\mathcal{F}_{1}(n-m,2m)\cup\mathcal{F}_{3}(m-n,2n),$
$\text{Case 3 }(m\leq p<q):$ $\qquad(g,c)\in\mathcal{C}_{3,1}\cup\mathcal{C}_{3,2}\cup\mathcal{C}_{3,3}\cup\mathcal{C}_{3,4}\cup\mathcal{F}_{1}(3m-n,2m),$
where
$\mathcal{C}_{1,1}$ $=[\,m-n+p:m-n+p-q:q\mid 1\,]\,$,
$\mathcal{C}_{1,2}$ $=[\,m-n-p:m-n-p+q:q\mid 1\,]\,$,
$\mathcal{C}_{1,3}$ $=[\,m-n-p:m-n+p-q:2(m-n)-q\mid 1\,]\,$,
$\mathcal{C}_{1,4}$ $=\Big[m-n-p:m-n+p-q:q\mid 1+\frac{2p(q-p)(m-n-q)}{q(m-n-p)(m-n+p-q)}\Big],$
$\mathcal{C}_{2,1}$ $=[\,m-n+p:-m-n+p+q:2m-q\mid 1\,]\,$,
$\mathcal{C}_{2,2}$ $=[\,m-n-p:3m-n-p-q:2m-q\mid 1\,]\,$,
$\mathcal{C}_{2,3}$ $=[\,m-n-p:-m-n+p+q:-2n+q\mid 1\,]\,$,
$\mathcal{C}_{2,4}$ $=\Big[m-n-p:-m-n+p+q:2m-q\mid 1+\frac{2p(m+n-q)(2m-p-q)}{(m-n-p)(m+n-p-q)(2m-q)}\Big],$
$\mathcal{C}_{3,1}$ =[ $=[\,3m-n-p:m-n-p+q:2m-q\mid 1\,]\,$,
$\mathcal{C}_{3,2}$ $=[\,-m-n+p:m-n+p-q:2m-q\mid 1\,]\,$,
$\mathcal{C}_{3,3}$ $=[\,-m-n+p:m-n-p+q:-2n+q\mid 1\,]\,$,
$\mathcal{C}_{3,4}$ $=\Big[-m-n+p:m-n-p+q:2m-q\mid 1+\frac{2(p-q)(m+n-q)(2m-p)}{(m+n-p)(m-n-p+q)(2m-q)}\Big].$
In accordance with symmetry relations between circled $\operatorname{SU}(m|n)$ and $\operatorname{SU}(n|m)$ Dynkin diagrams, the family of solution spaces $\mathcal{C}_{k,l}$ and $\mathcal{F}_{k}(u,v)$ associated with $\operatorname{SU}(m|n)$ is closed under simultaneously mapping the parameters and arguments as
$m\mapsto n,\quad n\mapsto m,\quad p\mapsto m+n-q,\quad q\mapsto m+n-p;\qquad x_{1}\mapsto-x_{3},\quad x_{2}\mapsto-x_{2},\quad x_{3}\mapsto-x_{1}.$
Concretely, under these simultaneous maps,
$\mathcal{C}_{k,1}\mapsto\mathcal{C}_{4-k,3},\qquad\mathcal{C}_{k,2}\mapsto\mathcal{C}_{4-k,2},\qquad\mathcal{C}_{k,3}\mapsto\mathcal{C}_{4-k,1},\qquad\mathcal{C}_{k,4}\mapsto\mathcal{C}_{4-k,4},\qquad k=1,2,3,$
and
$\mathcal{F}_{1}\mapsto\mathcal{F}_{3},\qquad\mathcal{F}_{3}\mapsto\mathcal{F}_{1}.$
Every $\mathcal{C}_{k,l}$ in Theorem 4.11 is of the form $x_{1}:x_{2}:x_{3}\mid c\,]$, so $\mathcal{C}_{k,l}=\emptyset$ if and only if $x_{1}x_{2}x_{3}=0$. For q=2(mn)q=2(m-n), for example, $\mathcal{C}_{1,3}=[\,m-n-p:-m+n+p:0\mid 1\,]=\emptyset$. Corollary 4.12 below summarises the result of analysing all possible linear relations between the parameters $m,n,p,q$.
Corollary 4.12.
Let the notation be as in Theorem 4.11. Then, the numbers and types of $G$-invariant solutions to the Einstein equation are as follows.
Case 1 $(p<q\leq m)$ with $m>n$:
  • Four solutions up to scaling: $\mathcal{C}_{1,1},\mathcal{C}_{1,2},\mathcal{C}_{1,3},\mathcal{C}_{1,4},\quad$ if $p\neq m-n\ \&\ q-p\neq m-n\ \&\ q\neq 2(m-n)$.
  • Three solutions up to scaling: $\mathcal{C}_{1,1},\mathcal{C}_{1,2},\mathcal{C}_{1,4},\quad$ if $p\neq m-n\ \&\ q-p\neq m-n\ \&\ q=2(m-n)$.
  • One solution up to scaling: $\begin{cases}\mathcal{C}_{1,1},\quad\text{if }\ p=m-n\ \&\ q-p\neq m-n,\\[4.26773pt] \mathcal{C}_{1,2},\quad\text{if }\ p\neq m-n\ \&\ q-p=m-n.\end{cases}$
  • Continuous family of Ricci-flat solutions: $\mathcal{F}_{3}(m-n,2(m-n))$.
Case 1 $(p<q\leq m)$ with $m=n$:
  • Four solutions up to scaling: $\mathcal{C}_{1,1},\mathcal{C}_{1,2},\mathcal{C}_{1,3},\mathcal{C}_{1,4}$.
Case 1 $(p<q\leq m)$ with $m<n$:
  • Four solutions up to scaling: $\mathcal{C}_{1,1},\mathcal{C}_{1,2},\mathcal{C}_{1,3},\mathcal{C}_{1,4},\quad$ if $p\neq n-m\ \&\ q-p\neq n-m$.
  • Three solutions up to scaling: $\begin{cases}\mathcal{C}_{1,1},\mathcal{C}_{1,3},\mathcal{C}_{1,4},\quad\text{if }\ p\neq n-m\ \&\ q-p=n-m,\\[4.26773pt] \mathcal{C}_{1,2},\mathcal{C}_{1,3},\mathcal{C}_{1,4},\quad\text{if }\ p=n-m\ \&\ q-p\neq n-m.\end{cases}$
Case 2 $(p<m<q)$ with $m>n$:
  • Four solutions up to scaling: $\mathcal{C}_{2,1},\mathcal{C}_{2,2},\mathcal{C}_{2,3},\mathcal{C}_{2,4},\quad$ if $p\neq m-n\ \&\ p+q\neq 3m-n\ \&\ q\neq 2n$.
  • Three solutions up to scaling: $\begin{cases}\mathcal{C}_{2,1},\mathcal{C}_{2,2},\mathcal{C}_{2,4},\quad\text{if }\ p\neq m-n\ \&\ p\neq 3(m-n)\ \&\ q=2n,\\[4.26773pt] \mathcal{C}_{2,1},\mathcal{C}_{2,3},\mathcal{C}_{2,4},\quad\text{if }\ p\neq m-n\ \&\ p+q=3m-n\ \&\ q\neq 2n.\end{cases}$
  • Two solutions up to scaling: $\mathcal{C}_{2,1},\mathcal{C}_{2,4},\quad$ if p=3(mn)&q=2n\ p=3(m-n)\ \&\ q=2n.
  • One solution up to scaling: $\begin{cases}\mathcal{C}_{2,1},\quad\text{if }\ p=m-n\ \&\ q\neq 2n,\\[4.26773pt] \mathcal{C}_{2,2},\quad\text{if }\ p\neq m-n\ \&\ p+q=m+n\ \&\ q\neq 2n.\end{cases}$
  • Continuous family of Ricci-flat solutions: $\mathcal{F}_{3}(m-n,2n)$.
Case 2 $(p<m<q)$ with $m=n$:
  • Four solutions up to scaling: $\mathcal{C}_{2,1},\mathcal{C}_{2,2},\mathcal{C}_{2,3},\mathcal{C}_{2,4},\quad$ if $p+q\neq 2m$.
  • No solutions if $p+q=2m$.
Case 2 $(p<m<q)$ with $m<n$:
  • Four solutions up to scaling: $\mathcal{C}_{2,1},\mathcal{C}_{2,2},\mathcal{C}_{2,3},\mathcal{C}_{2,4},\quad$ if $p\neq n-m\ \&\ p+q\neq 3m-n\ \&\ q\neq 2m$.
  • Three solutions up to scaling: $\begin{cases}\mathcal{C}_{2,1},\mathcal{C}_{2,3},\mathcal{C}_{2,4},\quad\text{if }\ p\neq n-m\ \&\ p+q=3m-n\ \&\ q\neq 2m,\\[4.26773pt] \mathcal{C}_{2,2},\mathcal{C}_{2,3},\mathcal{C}_{2,4},\quad\text{if }\ p=n-m\ \&\ q\neq 4m-2n\ \&\ q\neq 2m.\end{cases}$
  • Two solutions up to scaling: $\mathcal{C}_{2,3},\mathcal{C}_{2,4},\quad$ if p=nm&q=4m2n\ p=n-m\ \&\ q=4m-2n.
  • One solution up to scaling: $\begin{cases}\mathcal{C}_{2,2},\quad\text{if }\ p\neq n-m\ \&\ p+q=m+n\ \&\ q\neq 2m,\\[4.26773pt] \mathcal{C}_{2,3},\quad\text{if }\ p\neq n-m\ \&\ q=2m.\end{cases}$
  • Continuous family of Ricci-flat solutions: $\mathcal{F}_{1}(n-m,2m)$.
Case 3 $(m\leq p<q)$ with $m>n$:
  • Four solutions up to scaling: $\mathcal{C}_{3,1},\mathcal{C}_{3,2},\mathcal{C}_{3,3},\mathcal{C}_{3,4},\quad$ if $q-p\neq m-n\ \&\ q\neq 2n$.
  • Three solutions up to scaling: $\begin{cases}\mathcal{C}_{3,1},\mathcal{C}_{3,2},\mathcal{C}_{3,4},\quad\text{if }\ q-p\neq m-n\ \&\ q=2n,\\[4.26773pt] \mathcal{C}_{3,1},\mathcal{C}_{3,3},\mathcal{C}_{3,4},\quad\text{if }\ q-p=m-n\ \&\ q\neq 2n.\end{cases}$
Case 3 $(m\leq p<q)$ with $m=n$:
  • Four solutions up to scaling: $\mathcal{C}_{3,1},\mathcal{C}_{3,2},\mathcal{C}_{3,3},\mathcal{C}_{3,4}$.
Case 3 $(m\leq p<q)$ with $m<n$:
  • Four solutions up to scaling: $\mathcal{C}_{3,1},\mathcal{C}_{3,2},\mathcal{C}_{3,3},\mathcal{C}_{3,4},\quad$ if $p\neq 3m-n\ \&\ q-p\neq n-m\ \&\ q\neq 2m$.
  • Three solutions up to scaling: $\mathcal{C}_{3,2},\mathcal{C}_{3,3},\mathcal{C}_{3,4},\quad$ if $p=3m-n\ \&\ q-p\neq n-m\ \&\ q\neq 2m$.
  • One solution up to scaling: $\begin{cases}\mathcal{C}_{3,2},\quad\text{if }\ q-p=n-m\ \&\ q\neq 2m,\\[4.26773pt] \mathcal{C}_{3,3},\quad\text{if }\ q-p\neq n-m\ \&\ q=2m.\end{cases}$
  • Continuous family of Ricci-flat solutions: $\mathcal{F}_{1}(3m-n,2m)$.
The Ricci-flat solutions $\mathcal{F}_{3}(m-n,2(m-n))$ are only present if $n<m\leq 2n$; $\mathcal{F}_{3}(m-n,2n)$ if $n<m<2n$; $\mathcal{F}_{1}(n-m,2m)$ if $m<n<2m$; and $\mathcal{F}_{1}(3m-n,2m)$ if $m<n\leq 2m$. Indeed, there are no Ricci-flat solutions for $m=n$. In fact, the Ricci-flat solutions arise exactly when $p,q,m,n$ are related so that $x_{1}=x_{2}=x_{3}=0$ in $\mathcal{C}_{k,l}=[\,x_{1}:x_{2}:x_{3}\mid c\,]$, which can only happen in $\mathcal{C}_{1,3}$, $\mathcal{C}_{2,1}$, $\mathcal{C}_{2,3}$ and $\mathcal{C}_{3,1}$.
Remark 4.13.
The isotropy representation of the homogeneous space $\operatorname{SU}(3)/\mathbb{T}^{2}$ splits into three inequivalent irreducible summands; see [Kim90]. One can identify an $\operatorname{SU}(3)$-invariant metric on this space with a triple $(x_{1},x_{2},x_{3})$, and it is well known [Kim90, PZ24] that such a metric is Kähler if and only if $x_{i}+x_{j}=x_{k}$ for some permutation $(i,j,k)$ of the set $\{1,2,3\}$. The similarity between this condition and the three first conditions in (4.24) suggests that the Ricci-flat metrics in Theorem 4.11 may be related to some Calabi–Yau-type structure on $M=G/K$.
By considering $n=0$ or $m=0$, our classification results reduce to the similar results in the non-super case. For $0=n<p<q<m$, we thus find that there are exactly four $G$-invariant Einstein metrics, namely $\mathcal{C}_{1,1},\mathcal{C}_{1,2},\mathcal{C}_{1,3},\mathcal{C}_{1,4}$, while for $0=m<p<q<n$, there are exactly four $G$-invariant Einstein metrics, namely $\mathcal{C}_{3,1},\mathcal{C}_{3,2},\mathcal{C}_{3,3},\mathcal{C}_{3,4}$. This is in agreement with the analysis in [Arv93].
4.4.5 Positive metrics
We recall that $g=(x_{1},x_{2},x_{3})$ is said to be positive if $x_{1},x_{2},x_{3}>0$. Following from Corollary 4.12, Proposition 4.14 below lists all the positive Einstein metrics found in Theorem 4.11. Corresponding to each $\mathcal{C}_{k,l}$, it is convenient to introduce
$\mathcal{C}_{k,l}^{\pm}=\big\{(\lambda g,\tfrac{c}{\lambda})\,|\,g=(x_{1}^{k,l},x_{2}^{k,l},x_{3}^{k,l});\,c=c^{k,l};\,\lambda\in\mathbb{R}^{\pm}\big\},$
where $x_{1}^{k,l},x_{2}^{k,l},x_{3}^{k,l},c^{k,l}$ are the parameters used in the characterisation of $\mathcal{C}_{k,l}=[\,x_{1}^{k,l}:x_{2}^{k,l}:x_{3}^{k,l}\mid c^{k,l}\,]$ in Theorem 4.11. For example, for $(m-n-p)(m-n+p-q)(2m-2n-q)\neq 0$, we have
$\mathcal{C}_{1,3}^{-}=\big\{\big(\lambda(m-n-p,m-n+p-q,2m-2n-q),\tfrac{1}{\lambda}\big)\,|\,\lambda<0\big\}.$
We also introduce
$\mathcal{F}_{1}^{+}=\big\{\big((x_{2}+x_{3},x_{2},x_{3}),0\big)\,|\,x_{2},x_{3}>0\big\},\qquad\mathcal{F}_{3}^{+}=\big\{\big((x_{1},x_{2},x_{1}+x_{2}),0\big)\,|\,x_{1},x_{2}>0\big\}.$
Proposition 4.14.
Let the notation be as in Theorem 4.11. Then, the numbers and types of positive $G$-invariant solutions to the Einstein equation are as follows.
Case 1 $(p<q\leq m)$:
  • Four solutions up to scaling: $\mathcal{C}_{1,1}^{+},\mathcal{C}_{1,2}^{+},\mathcal{C}_{1,3}^{+},\mathcal{C}_{1,4}^{+},\quad$ if $\max(p,q-p)<m-n$.
  • One solution up to scaling: $\begin{cases}\mathcal{C}_{1,1}^{+},\quad\text{if }\ q-p<m-n\leq p,\\[4.26773pt] \mathcal{C}_{1,2}^{+},\quad\text{if }\ p<m-n\leq q-p,\\[4.26773pt] \mathcal{C}_{1,3}^{-},\quad\text{if }\ m-n<\min(p,q-p).\end{cases}$
  • Continuous family of Ricci-flat solutions: $\mathcal{F}_{3}^{+},\quad$ if p=mn&q=2(mn)\ p=m-n\ \&\ q=2(m-n).
Case 2 $(p<m<q)$:
[…]
  • One solution up to scaling: $\begin{cases}\mathcal{C}_{2,1}^{+},\quad\text{if }\ m-n\leq p\ \&\ n-m<m+n-q<p,\\[4.26773pt] \mathcal{C}_{2,2}^{+},\quad\text{if }\ p\leq m+n-q\ \&\ p<m-n,\\[4.26773pt] \mathcal{C}_{2,2}^{-},\quad\text{if }\ m+n-q\leq p\ \&\ m+n-q<n-m,\\[4.26773pt] \mathcal{C}_{2,3}^{-},\quad\text{if }\ n-m\leq m+n-q\ \&\ m-n<p<m+n-q.\end{cases}$
  • Continuous family of Ricci-flat solutions: $\begin{cases}\mathcal{F}_{1}^{+},\quad\text{if }\ p=n-m\ \&\ q=2m,\\[4.26773pt] \mathcal{F}_{3}^{+},\quad\text{if }\ p=m-n\ \&\ q=2n.\end{cases}$
Case 3 $(m\leq p<q)$:
  • Four solutions up to scaling: $\mathcal{C}_{3,1}^{-},\mathcal{C}_{3,2}^{-},\mathcal{C}_{3,3}^{-},\mathcal{C}_{3,4}^{-},\quad$ if $\max(m+n-q,q-p)<n-m$.
  • One solution up to scaling: $\begin{cases}\mathcal{C}_{3,1}^{+},\quad\text{if }\ n-m<\min(m+n-q,q-p),\\[4.26773pt] \mathcal{C}_{3,2}^{-},\quad\text{if }\ m+n-q<n-m\leq q-p,\\[4.26773pt] \mathcal{C}_{3,3}^{-},\quad\text{if }\ q-p<n-m\leq m+n-q.\end{cases}$
  • Continuous family of Ricci-flat solutions: $\mathcal{F}_{1}^{+},\quad$ if p=3mn&q=2m\ p=3m-n\ \&\ q=2m.
For any $(g,c)\in\mathcal{C}_{k,l}^{\epsilon}$ where $l<4$ in Proposition 4.14, the Einstein constant is positive if $\epsilon=+$ and negative if $\epsilon=$-. If $l=4$, on the other hand, the sign of the Einstein constant depends intricately on the parameters $m,n,p,q$.
In the situation we are considering, when $M=G/K$ is constructed from $G=\operatorname{SU}(m|n)$ with $K$ given in (4.16), Corollary 4.3 and Proposition 4.4 imply that the isotropy summand $\mathfrak{m}_{i}$ is neither purely even nor purely odd for at least two $i\in\{1,2,3\}$. As a consequence, $Q$ has mixed signature on $\mathfrak{m}_{\bar{0}}=(\mathfrak{m}_{1})_{\bar{0}}\oplus(\mathfrak{m}_{2})_{\bar{0}}\oplus(\mathfrak{m}_{3})_{\bar{0}}$, and $Q|_{\mathfrak{m}_{\bar{0}}}$ is not positive-definite. This means that no $G$-invariant metric on $M$ induces a Riemannian metric on the reduced manifold $|M|$. For metrics on $M=G/K$, the notion of positivity introduced in Section 3.6 thus appears to be a natural super-geometric counterpart to positive definiteness.
The finiteness conjecture in classical geometry states that a compact homogeneous space $G_{0}/K_{0}$ with pairwise inequivalent isotropy summands can support only finitely many $G_{0}$-invariant Einstein metrics up to scaling; see [BWZ04]. The existence of continuous families of positive Ricci-flat metrics established in Proposition 4.14 demonstrates that this conjecture fails on supermanifolds. This is particularly intriguing since, as we showed in Section 3.6.2, the formulas for the Ricci curvature on $M$ are similar to those in the non-super setting. The main difference comes from the possible non-positivity of the dimension parameters and possible negativity of the structure constants.
In classical homogeneous geometry, Bochner’s vanishing theorem (see [Bes87, Theorem 1.84]) implies that a compact homogeneous space $G_{0}/K_{0}$ can support a $G_{0}$-invariant Riemannian metric with non-positive Ricci curvature only if the identity component of $G_{0}$ is a torus. The existence of positive $\mathrm{SU}(m|n)$-invariant Einstein metrics on $M$ with vanishing or negative Ricci curvature shows that this conclusion does not extend to homogeneous superspaces.
4.5 $\operatorname{SOSp}(2|2n)$: two isotropy summands
Let $G=\operatorname{SOSp}(2|2n)$ be as in (4.11). We will construct a flag supermanifold $M=G/K$ whose isotropy representation decomposes into two inequivalent nonzero irreducible $\operatorname{ad}_{\mathfrak{k}}$-modules:
$\mathfrak{m}=\mathfrak{m}_{1}\oplus\mathfrak{m}_{2}.$
Retaining the notation from Section 3.6 and Section 4.2.2, a $G$-invariant graded Riemannian metric on $M$ is of the form
$g=x_{1}Q|_{\mathfrak{m}_{1}}+x_{2}Q|_{\mathfrak{m}_{2}},\qquad x_{1},x_{2}\in\mathbb{R}^{\times}.$ (4.25)
We will use $(x_{1},x_{2})$ as a shorthand notation for the sum in (4.25). The coefficients in the expression (3.25) follow from (4.13) and are given by $b_{1}=b_{2}=-2n$. In the following, we will denote this common value simply by $b$:
$b=-2n.$
We construct $M$ using a Dynkin diagram of $G$ with the $p^{\text{th}}\!$ node circled, where
$2\leq p\leq n,$
as in Figure 2, and with
$K=\operatorname{U}(1|p-1)\times\operatorname{Sp}(2(n+1-p)).$ (4.26)
The compact real form associated with (4.26) is $\mathfrak{k}=\mathfrak{u}(1|p-1)\oplus\mathfrak{sp}(2(n+1-p))$, and the scalar superproduct $Q$ on $\mathfrak{m}$, where $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{m}$, is defined by the corresponding restriction of (4.12).
[…]
Figure 2: Dynkin diagram of $\operatorname{SOSp}(2|n)$ with one circled node.
4.5.1 Isotropy summands
Let $\mathfrak{m}^{\mathbb{C}}$ denote the $Q$-orthogonal complement of $\mathfrak{k}^{\mathbb{C}}=\mathfrak{gl}(1|p-1)^{\mathbb{C}}\oplus\mathfrak{sp}(2(n+1-p))^{\mathbb{C}}$ in $\mathfrak{g}^{\mathbb{C}}=\mathfrak{osp}(2|2n)^{\mathbb{C}}$, where $\mathfrak{k}^{\mathbb{C}}$ is the complexification of $\mathfrak{k}$. For the root-space decomposition (4.2) of $\mathfrak{m}^{\mathbb{C}}$, we introduce
$\mathfrak{m}_{k}^{\mathbb{C}}:=\bigoplus_{\alpha\in\Delta_{k}}\mathfrak{g}_{\alpha}^{\mathbb{C}},\qquad\Delta_{k}:=\big\{\sum_{i=1}^{n+1}c_{i}\alpha_{i}\in\Delta_{M}\,|\,c_{p}=k\big\},\qquad k\in I_{\mathfrak{m}^{\mathbb{C}}},$
where $I_{\mathfrak{m}^{\mathbb{C}}}=\{-2,-1,1,2\}$. If $\Delta_{k}$ is non-empty, then $\mathfrak{m}_{k}^{\mathbb{C}}$ is an $\operatorname{ad}_{\mathfrak{k}^{\mathbb{C}}}$-invariant subspace of $\mathfrak{m}^{\mathbb{C}}$.
[…]
Proposition 4.15.
As an $\operatorname{ad}_{\mathfrak{k}^{\mathbb{C}}}$-module, $\mathfrak{m}^{\mathbb{C}}$ admits the $Q$-orthogonal decomposition
$\mathfrak{m}^{\mathbb{C}}=\mathfrak{m}_{-2}^{\mathbb{C}}\oplus\mathfrak{m}_{-1}^{\mathbb{C}}\oplus\mathfrak{m}_{1}^{\mathbb{C}}\oplus\mathfrak{m}_{2}^{\mathbb{C}}.$
Moreover, $\mathfrak{m}_{i}^{\mathbb{C}}$ and $\mathfrak{m}_{-i}^{\mathbb{C}}$ are irreducible and dual for each $i=1,2$.
Proof.
Let $\pm\varepsilon,\pm\delta_{1},\dots,\pm\delta_{n}$ denote the weights of $\nu_{2|2n}$. In accord with the regular embedding $\mathfrak{k}^{\mathbb{C}}\subseteq\mathfrak{g}^{\mathbb{C}}$, we denote by $\varepsilon,\delta_{1},\dots,\delta_{p-1}$ the weights of $\mu_{1|p-1}$, by $\varepsilon,-\delta_{1},\dots,-\delta_{p-1}$ the weights of $\bar{\mu}_{1|p-1}$, and by $\pm\delta_{p},\dots,\pm\delta_{n}$ the weights of $\nu_{0|2(n+1-p)}$. We then have
[…]
Restricting to $\mathfrak{k}^{\mathbb{C}}$, we have $\nu_{2|2n}|_{\mathfrak{k}^{\mathbb{C}}}=\mu_{1|p-1}\oplus\bar{\mu}_{1|p-1}\oplus\nu_{0|2(n+1-p)}$ and hence the $\mathfrak{k}^{\mathbb{C}}$-module isomorphisms
[…]
$\quad\oplus(\mu_{1|p-1}\otimes\bar{\mu}_{1|p-1})\oplus(\mu_{1|p-1}\otimes\nu_{0|2(n+1-p)})\oplus(\bar{\mu}_{1|p-1}\otimes\nu_{0|2(n+1-p)})$
[…]
As $\mathfrak{k}^{\mathbb{C}}$ arises as direct summand, it follows that
$\mathfrak{m}^{\mathbb{C}}=\mathfrak{m}_{-2}^{{}^{\prime}\mathbb{C}}\oplus\mathfrak{m}_{-1}^{{}^{\prime}\mathbb{C}}\oplus\mathfrak{m}_{1}^{{}^{\prime}\mathbb{C}}\oplus\mathfrak{m}_{2}^{{}^{\prime}\mathbb{C}},$
where
[…]
noting that $\mathfrak{m}_{-i}^{{}^{\prime}\mathbb{C}}$ and $\mathfrak{m}_{i}^{{}^{\prime}\mathbb{C}}$ are irreducible and dual for each $i=1,2$.
We now show that $\mathfrak{m}_{k}^{\mathbb{C}}=\mathfrak{m}_{k}^{{}^{\prime}\mathbb{C}}$ for all $k\in I_{\mathfrak{m}^{\mathbb{C}}}$, which then concludes the proof. The weights of $\mathfrak{m}_{1}^{{}^{\prime}\mathbb{C}}$ are $\delta_{i}\pm\delta_{j}$ and $\varepsilon\pm\delta_{j}$ with $1\leq i<p\leq j\leq n$, and since they are all contained in $\Delta_{1}$, the coefficient of $\alpha_{p}=\delta_{p-1}-\delta_{p}$ must be $1$. This implies that $\mathfrak{m}_{1}^{{}^{\prime}\mathbb{C}}\subseteq\mathfrak{m}_{1}^{\mathbb{C}}$. Because
[…]
the weights of $\mathfrak{m}_{2}^{{}^{\prime}\mathbb{C}}$ are $\delta_{i}+\delta_{j}$ with $1\leq i\leq j\leq p-1$ and $\varepsilon+\delta_{i}$ with $1\leq i\leq p-1$. As all these weights are contained in $\Delta_{2}$, it follows that $\mathfrak{m}_{2}^{{}^{\prime}\mathbb{C}}\subseteq\mathfrak{m}_{2}^{\mathbb{C}}$. We similarly have $\mathfrak{m}_{-1}^{{}^{\prime}\mathbb{C}}\subseteq\mathfrak{m}_{-1}^{\mathbb{C}}$ and $\mathfrak{m}_{-2}^{{}^{\prime}\mathbb{C}}\subseteq\mathfrak{m}_{-2}^{\mathbb{C}}$, so
$\bigoplus_{k\in I_{\mathfrak{m}^{\mathbb{C}}}}\mathfrak{m}_{k}^{\mathbb{C}}\subseteq\mathfrak{m}=\bigoplus_{k\in I_{\mathfrak{m}^{\mathbb{C}}}}\mathfrak{m}_{k}^{{}^{\prime}\mathbb{C}}\subseteq\bigoplus_{k\in I_{\mathfrak{m}^{\mathbb{C}}}}\mathfrak{m}_{k}^{\mathbb{C}}.$
It follows that $\mathfrak{m}_{k}^{\mathbb{C}}=\mathfrak{m}_{k}^{{}^{\prime}\mathbb{C}}$ for all $k\in I_{\mathfrak{m}^{\mathbb{C}}}$. ∎
Corollary 4.16.
Let the notation be as in Proposition 4.15. Then,
[…]
with corresponding highest weights
$\Lambda_{1}=\varepsilon+\delta_{p},\qquad\Lambda_{2}=\varepsilon+\delta_{1}.$
Proposition 4.17.
The $\operatorname{ad}_{\mathfrak{k}}$-representation $\mathfrak{m}$ admits a $Q$-orthogonal decomposition of the form
$\mathfrak{m}\cong\mathfrak{m}_{1}\oplus\mathfrak{m}_{2},$
where the summands are irreducible $\operatorname{ad}_{\mathfrak{k}}$-representations satisfying
$[\mathfrak{m}_{1},\mathfrak{m}_{1}]\subseteq\mathfrak{k}\oplus\mathfrak{m}_{2},\qquad[\mathfrak{m}_{1},\mathfrak{m}_{2}]\subseteq\mathfrak{m}_{1},\qquad[\mathfrak{m}_{2},\mathfrak{m}_{2}]\subseteq\mathfrak{k}.$
Proof.
The proof is analogous to that of Proposition 4.4. ∎
By Corollary 4.16, the $\operatorname{ad}_{\mathfrak{k}^{\mathbb{C}}}$-modules $\mathfrak{m}_{1}^{\mathbb{C}}$ and $\mathfrak{m}_{2}^{\mathbb{C}}$ have distinct highest weights and are therefore non-isomorphic. It follows from Proposition 4.17 that the flag supermanifold $G/K$ has two inequivalent irreducible isotropy summands.
It follows from the proof of Proposition 4.17 that, for each $i=1,2$,
$\{A_{\alpha},B_{\alpha}\,|\,\alpha\in\Delta_{i}\cap\Delta_{\bar{0}}^{+}\}\cup\{\sqrt{\imath}A_{\alpha},\sqrt{\imath}B_{\alpha}\,|\,\alpha\in\Delta_{i}\cap\Delta_{\bar{1}}^{+}\}$
is an $\mathbb{R}$-basis for the $\operatorname{ad}_{\mathfrak{k}}$-module $\mathfrak{m}_{i}$, with
$d_{i}:=\operatorname{sdim}(\mathfrak{m}_{i})=2\operatorname{sdim}_{\mathbb{C}}(\mathfrak{m}_{i}^{\mathbb{C}}).$ (4.27)
Explicit expressions for $d_{1},d_{2}$ are readily obtained and are given in Lemma 4.20 below.
4.5.2 Structure constants and Ricci curvature
Proposition 4.18.
Let $K$ be as in (4.26) and $\mathfrak{m}$ as in Proposition 4.17. Then, the Casimir eigenvalues (3.30) are given by
$c_{1}=-n+\tfrac{1}{2}(p-1),\qquad c_{2}=-(p-1).$
Proof.
The eigenvalues of $C_{\mathfrak{m}_{i},Q|_{\mathfrak{k}}}$ are given by $c_{i}=-(\Lambda_{i}+2\rho_{\mathfrak{k}^{\mathbb{C}}},\Lambda_{i})$, $i=1,2$, where $\rho_{\mathfrak{k}^{\mathbb{C}}}$ is the Weyl vector of $\mathfrak{k}^{\mathbb{C}}$, and where $\Lambda_{1},\Lambda_{2}$ are the highest weights given in Corollary 4.16. To compute these eigenvalues, we use that
$(\lambda,\mu)=\sum_{i=1}^{n+1}\lambda(h_{i})\mu(\bar{h}_{i}),\qquad\lambda,\mu\in(\mathfrak{h}^{\mathbb{C}})^{\vee},$
where $\{h_{1}=E_{11}-E_{22},\,h_{i}=E_{i+1,i+1}-E_{n+1+i,n+1+i}\,|\,2\leq i\leq n+1\}$ is a basis for the Cartan subalgebra $\mathfrak{h}^{\mathbb{C}}$ (cf. (4.10)). The right $Q$-dual basis elements are given by
$\bar{h}_{1}=-\tfrac{1}{2}h_{1},\qquad\bar{h}_{i}=\tfrac{1}{2}h_{i},\qquad i=2,\ldots,n+1,$
so
$(\varepsilon,\varepsilon)=\sum_{i=1}^{n+1}\varepsilon(h_{i})\varepsilon(\bar{h}_{i})=-\frac{1}{2},\qquad(\delta_{j},\delta_{j})=\sum_{i=1}^{n+1}\delta_{j}(h_{i})\delta_{j}(\bar{h}_{i})=\frac{1}{2},\qquad j=1,\ldots,n.$
Since
$2\rho_{\mathfrak{k}^{\mathbb{C}}}=(1-p)\varepsilon+\sum_{j=1}^{p-1}(p-2j+1)\delta_{j}+\sum_{j=p}^{n}\!\big(2(n+1-j)\big)\delta_{j},$
and the highest weights are given by $\Lambda_{1}=\varepsilon+\delta_{p}$ and $\Lambda_{2}=\varepsilon+\delta_{1}$, we have
$c_{1}$ $=-(\Lambda_{1}+2\rho_{\mathfrak{k}^{\mathbb{C}}},\Lambda_{1})=(p-2)(\varepsilon,\varepsilon)-(2(n+1-p)+1)(\delta_{p},\delta_{p})=-\tfrac{1}{2}(2n-p+1),$
$c_{2}$ $=-(\Lambda_{2}+2\rho_{\mathfrak{k}^{\mathbb{C}}},\Lambda_{2})=(p-2)(\varepsilon,\varepsilon)-p(\delta_{1},\delta_{1})=1-p.$
Corollary 4.19.
Let the notation be as in Proposition 4.18. Then,
$b=2c_{1}+c_{2}.$
Lemma 4.20.
Let the notation be as in Proposition 4.17 and (4.27). Then,
$d_{1}=-2(p-2)(b-2c_{2}),\qquad d_{2}=-(p-2)(b-2c_{1}).$
Proof.
The result follows from Corollary 4.16, (2.1) and (4.27). ∎
Proposition 4.21.
Let $K$ be as in (4.26) and $\mathfrak{m}$ as in Proposition 4.17. Then, up to permutation, $[112]$ is the only possibly nonzero structure constant, and it is given as
$[112]=-(p-2)(b-2c_{1})(b-2c_{2}).$
Proof.
Proposition 4.17 implies that $[112]$ is the only possibly nonzero structure constant up to permutation. Using Proposition 3.25 and the symmetry of $[ijk]$, we have $112]=[211]=d_{2}(b_{2}-2c_{2})$. The result now follows using $b_{2}=b$, Proposition 4.18 and Lemma 4.20. ∎
Concretely, $b-2c_{1}=-(p-1)$ and $b-2c_{2}=-2(n+1-p)$, and since $2\leq p\leq n$, it follows that $(b-2c_{1})(b-2c_{2})>0$, $112]\leq 0$ and that $[112]=0$ if and only if $p=2$.
Proposition 4.22.
Let $g=(x_{1},x_{2})$ be a $G$-invariant graded Riemannian metric on $M=G/K$, where $G=\operatorname{SOSp}(2|2n)$ and $K$ is given in (4.26). Set $x^{2}=2x_{1}^{2}+x_{2}^{2}$. Then, the following holds.
  • The Ricci coefficients in (3.27) are given by
    $r_{i}=\frac{b}{2}+\frac{(b-2c_{i})x_{i}(2x_{i}^{2}-x^{2})}{4x_{1}x_{1}x_{2}},\qquad i=1,2.$
  • The scalar curvature $S$ is given by
    $S=\frac{b}{2}\Big(\frac{d_{1}}{x_{1}}+\frac{d_{2}}{x_{2}}\Big)-\frac{[112]\,x^{2}}{4x_{1}x_{1}x_{2}},$
    with $d_{1},d_{2}$ given in Lemma 4.20 and $[112]$ in Proposition 4.21.
Proof.
If $p\neq 2$, then, by Lemma 4.20, $d_{1},d_{2}$ are nonzero, in which case Theorem 3.26 implies that
$r_{1}=\frac{b_{1}}{2}-\frac{[112]}{2d_{1}}\frac{x_{2}}{x_{1}},\qquad r_{2}=\frac{b_{2}}{2}+\frac{[211]}{4d_{2}}\Big(\frac{x_{2}^{2}}{x_{1}^{2}}-2\Big).$
Using $b_{1}=b_{2}=b$ and Proposition 4.21, the desired expressions for the Ricci coefficients follow. If $p=2$, then $d_{1}=d_{2}=0$. By Proposition 4.17, $(2,1)$ is $1$-selected, and $(1,1)$ is $2$-selected. The desired expressions for the Ricci coefficients now follow from Proposition 3.28. Finally, the expression for $S$ follows from Proposition 3.27. ∎
4.5.3 Classification of Einstein metrics
Using Proposition 4.22, we can write the Einstein equations (3.28) as
$c=\frac{b}{2x_{i}}+\frac{(b-2c_{i})(2x_{i}^{2}-x^{2})}{4x_{1}x_{1}x_{2}},\qquad i=1,2.$ (4.28)
Eliminating $c$ by combining the two equations in (4.28) and using $b=2c_{1}+c_{2}$ and $x^{2}=2x_{1}^{2}+x_{2}^{2}$, yields
$(2x_{1}-x_{2})(c_{2}x_{1}-c_{1}x_{2})=0.$
It follows from Proposition 4.18 that $c_{1},c_{2}\neq 0$, so, up to scaling, there are exactly two solutions to the Einstein equations:
$(\mathrm{S1}):\quad(x_{1},x_{2})=(1,2),\qquad\ \ c=c_{1},$
$(\mathrm{S2}):\quad(x_{1},x_{2})=(c_{1},c_{2}),\qquad c=1+\frac{(2c_{1}-c_{2})c_{2}}{4c_{1}^{2}}.$
This completes our analysis of the Einstein equations and gives rise to the classification result in Theorem 4.23 below. Mimicking the presentation in Section 4.4.4, for all $x_{1},x_{2},c\in\mathbb{R}$, we introduce
$[\,x_{1}:x_{2}\mid c\,]:=\begin{cases}\big\{(\lambda g,\tfrac{c}{\lambda})\,|\,g=(x_{1},x_{2});\,\lambda\in\mathbb{R}^{\times}\big\},\ &\text{if $x_{1}x_{2}\neq 0$,}\\[4.26773pt] \emptyset,\ &\text{if $x_{1}x_{2}=0$.}\end{cases}$
Theorem 4.23.
Let $M=G/K$ be a flag supermanifold, where $G=\operatorname{SOSp}(2|2n)$ and $K$ is given in (4.26) with $2\leq p\leq n$. Then, the $G$-invariant Einstein metrics $g$ on $M$, with corresponding Einstein constant $c$, are of the form (4.25) and classified as follows:
$(g,c)\in[\,1:2\mid-n+\tfrac{1}{2}(p-1)\,]\cup\Big[2n+1-p:2(p-1)\mid-\frac{1}{2}-\frac{(p-1)(n+1-p)}{(2n+1-p)^{2}}\Big].$
We highlight the similarities between the solutions in Theorem 4.23 and the ones in Section 4.4.3 when the former are expressed as
$(g,c)\in[\,c_{1}:b-c_{2}\mid 1\,]\cup\Big[c_{1}:c_{2}\mid 1+\frac{(b-2c_{1})(b-2c_{1})(b-2c_{2})}{4c_{1}c_{1}c_{2}}\Big].$
Remark 4.24.
The various expressions above, including the ones in Theorem 4.23, shorten if $p$ is replaced by $q+1$. This would correspond to enumerating the nodes associated to $\alpha_{2},\ldots,\alpha_{n}$ in the $\operatorname{SOSp}(2|2n)$ Dynkin diagram in Section 4.2.2 from $1$ to n1n-1. For consistency with the enumeration in the $\operatorname{SU}(m|n)$ cases in Section 4.3 and Section 4.4, we have opted to use $p$ as indicated.
When restricting the scaling parameter $\lambda$ to $\mathbb{R}^{+}$, both solutions in Theorem 4.23 are positive with negative Einstein constant.
Acknowledgements
This work was supported by the Australian Research Council under the Discovery Project scheme, project numbers DP180102185, DP200102316 and DP220102530. The authors thank Andreas Arvanitoyeorgos, Ruibin Zhang and Gabriele Tartaglino-Mazzucchelli for discussions and comments.